A Complete Bibliography of Publications in the Journal of Theoretical Biology: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

27 July 2020 Version 1.01

Title word cross-reference

2 [KMT65, KT65, Tri64]. _{12} [VP65a]. _{max} [Fis62a]. α [Kno65, Par69, SS64]. n [Sav69b, TR65]. π [IKTN66]. r [Ber62c]. σ [IKTN66]. T_{m} [WGP66].

-chymotrypsin [Kno65]. -fibrous [Par69]. -fold [TR65]. -halogenoethylamines [KMT65, KT65, Tri64]. -limited [WGP66]. -pool [Sav69b]. -salicylic [BLT62a].

1-130 [Ano69s]. 1-185 [Ano69u]. 11 [Ano66f]. 13 [Ano69g].

2 [BS69]. 235 [Ano69g].

4 [Ber62c]. 4- [BLT62a].

86 [Ano66f]. ability [Her69]. Abnormal [Ess66]. absence [Pol65]. absolute [Wal69a]. absorption [ADB61, LMB68, WO67]. accelerated [Wan63]. account [WO67]. accounting [Doe66]. acety1cholinesterase [SW66]. acetylcholine [Kar67, KS68a]. Acid [Rus68b, Ata66, BLT62a, Boc67, Cam67, Dou62, Eck62, GD63, Haa69, Hav66, KS68b, Lin61, Mil66, RFCH68, RS66, Smi66, Syk66, ZES68]. Acid-base [Rus68b]. acids [CRJ66, KS67, Yca61]. across [EH69, ERH69b, HK66, Kim66, Løv63, PGH63, SKB69, VPT64, Vai65c, Vai66a, Vai69]. acting [Bla67a]. actinomycin [Sin65]. action [BBS63, CT67, Fel68, HOD63, Kla66, Koc67b, MR68, Mur66, NM68, PVP61, Per62a, Rig66, Sim65a, SW67, SA66, Tri64, VWM61]. actions [TP63]. activated [Sem67]. Activation [Lov69, Mad69, Sem67]. activators [Kat69]. Active [Acl66, Gro67, Rap68, Acl67, Acl69, Bre67, Bri69, Coh69, CD69, HK66, Jah62, Vai65a]. activity [CRJ66, Heb68, Kau67, Lau69b, LG68, Ros68, Sha61, Sne66]. adaptation [Cur68, Gau64, HCS66, LM65]. Adaptive [PK68, Lin63]. addition [Dia69]. adenyl [Mad69]. adenylate [RB68a]. adjusted [SB65]. Adjustment [QZ62]. adsorbed [HHT65a]. adsorption [Hed66]. adult [Zaj68]. advantages [Tom66b]. aerogenes [DH69]. affecting [RB68b]. after [Wev66]. Age [Noo68, Bur66]. age-specific [Bur66]. agencies [Why67]. agent [Alm69]. agents [Alm69, Sim65a, SA66]. aggregates [Bro64, Van67]. aggregating [Mag69a]. aging [Sam66]. aid [Ber63]. aids [Day65]. aldosterone [VWM61]. Algorithmically [Sta65]. algorithms [SG63]. alkaline [Mar64]. alkylating [BLT^{+}62b, BLT62a, HOD63, Heb68, Sim65a]. Alleles [Fis62d, Eck62, Lew62]. allied [Pet64]. Allocation [Noo66]. allometry [Ebe69]. allosteric [Kar67]. allows [AP65]. almost [EDT66]. almost-simultaneous [EDT66]. along [ZS65]. alternately [Ghe66]. alternately-oriented [Ghe66]. alternative [Par69, SP69]. amines [Sha61]. amino [Ata66, Eck62, GD63, Hav66, KS67, KS68b, Smi66, Yca61, ZES68]. aminotransferase [Haa69]. ammonium [Fel68, SW66, SW67]. amniotes [PK68]. amniotic [Cor65]. amoebae [BT62, Sha65]. amoeboid [Kav63, RB69]. among [BW63, PP68]. amphibian [HCS66]. amplification [Mor69]. amylase [Tho68]. Analog [Hei64, HE65a, JH69, KS68a, She67, YP64]. analogue [SF63]. analogues [Haa69, VP65a]. Analysis [BCB64a, Eng62, Fan69, Ged68, Lan63, Mag69a, RS62, RB68b, AKT62, AB61, BR66, BW63, CS65, CGTK67a, CGTK67b, CFY68, Cla67, FM69, Fri69b, GFR68, Haa69, HvR69, Hei64, Kel65, KS67, KS68b, Loc65, Mar61, Myh69, Orl68, PS68, Pri67a, Pri67b, Pri67c, Rap68, Reh68, RP68, Ric65, Sav69a, Sav69b, SS67b, SS67a, Ste68, Sug63, SF63, Sug66, Tak66, Tak68, TP67, Tri64, Van67, Zel68, SF63]. angular [Koc68b]. animal [AB61, Brl64, Geo69, Had67, Løv63, Mar61, O’N69, RD69, Wag68]. animals [BW63, Bro62]. anionic [SW66]. anions [Des65]. anisotropic [Bri69, Rap68]. Announcement [Ano63a, Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano69a].

Announcements [Ano66e, Ano69b]. anomalies [Boe63]. antagonistic [TP63]. antennal [Daw66]. anti [HOD63]. anti-tumour [HOD63]. antibodies [AP65, BA63]. antibody [AP65, Cam67, Lur69, PA63, PA68, TP63, Win63, WA65, vO63]. Antigen [vO63, AP65, PA63, PA68, TP63]. Antigen-antibody [vO63, AP65, PA63, PA68]. Antigenesis [Rin65]. Antigenic [Kil66a]. antigens [Sie69]. antimetabolites [BLT^{+}62b, BLT62a]. antisera [Kin64]. antitumor [Heb68]. Apparatus [Mel64a]. appearance [Fis62a, Yam66]. Applicability [MH69]. applicable [O’S62]. Application [Atl68, Cze64a, Cze64b, Cze67, Hen68, Ohk67, CS65, CFY68, DS67, Kar67, Kat67, Sch68, Sem67, Res69]. applied [Mar67, Pol63, Reh68]. approach [BO65, Ber61a, Eva62, FBG^{+}66a, HH69, Kor69, Lau65, Lau69a, Lau69b, Rub67, She62a, TL66]. approximation [BR66, Sav69b]. April [Ano61c, Ano66x, Ano67n, Ano68h, Ano69j]. aqueous [EH68, HT63, VG68]. area [HHT65c]. areas [Kom63]. arising [BBKL66, Par66, Par68]. Aromatic [Sol66a, NIT^{+}65]. Ascension [KCE69]. aspartate [Haa69]. aspect [ZZ69]. Aspects [Gre66a, VP65a, Alm69, Ber64b, CM62, Con69, Els62, FW65, IJ68, Jeh67, Lew67, Mur66, PVP61, Per62a, Pol61, Sie69, Sin67, Str62, Wal66b, Wal66a, ZZ67]. assemblies [Lin68c]. assembly [McM63]. associating [CFY68]. Association [PP68, BO65]. associations [KS67]. assumption [MH69, Wal67]. asymmetric [Yam66]. Asymptotic [Mac69, dLR67]. Atherosclerosis [Tab67]. Atmospheric [Ids68]. Atmospheric- [Ids68]. atomic [RM68]. ATP [BDF69]. attack [Mal63, Ozo69]. augmentation [Vid66]. August [Ano66w, Ano67q, Ano68j, Ano69l]. authors [ZZ69]. auto [Bur66]. auto-immunity [Bur66]. Autogenous [Wad65]. Automata [Arb67, Ros67]. Automatic [RP68, Cla63]. automaton [SR66b, Sta65, Sug63]. auxinic [CRJ66]. average [Big67]. axiomatic [HR69c].

b [Wad65, VP65a]. background [Mor66]. bacteria [AAB69, DeL67, Kal68, Koc68a, Koc68b, Noa68, Syk66]. Bacterial [Mar63, Sai68b, Sai68a, SAFT66, Ble69, De 69, Doe66, Jar64, War67b, Wu67, YP64]. bactericidal [LG68]. Bactérienne [BP62]. bacteriophage [Ber62c]. Bakteriengeißeln [Jar64]. Balance [JM66, LM66]. balances [Ell68, WW68a]. barriers [vO63]. Base [Gat64, BS69, Fre62a, GD63, Goe67a, Rus68b]. based [Ghe66, Jah62, MM68a, QZ62, TP63, War67a]. bases [RFCH68, WP63]. basis [Aid66, Brl64, Fon67a, Gin67, Han69, Lin67, Lin68c, Tak66, Tak68, Wad65]. be [Bra65, Coh67]. bee [Wen61]. beetle [Sta66]. behavior [Eml66, HR69c, HR69a, Nor68, RD69, Win67, Zaj68]. behaviour [Ham64a, Ham64b]. beim [Kli66]. benzoic [CRJ66]. berücksichtigung [WO67]. betrifft [Hel64]. betulinus [PP68]. between [BNW62, Big67, Boy62, CPC66, Cla67, CRJ66, Coh67, Dia69, GD63, Hav66, Hel68a, Hel68b, Kat69, Kau67, Koi66, NN66, Par67, Par64, PR66, Ren62, RPS61, Sha61, SZ65, Sim65b, Sim67, Sne66, Wag68]. beyond [Pla61]. bilayer [OF67, Ohk69]. bilayers [Ohk68, RH69]. bimolecular [EH69, ERH69b, HHT65b, HHT65c, HH66, HT63, MR68, Par67]. binary [WPY67]. Binding [Dan69, Dia69, Fis62a, Koc67c, SW66, WA65]. Biochemical [Psc63, Sav69a, Sav69b, Cze64a, Cze64b, Pri67a, Pri67b, Pri67c, RP68, Ros67, WPY67]. Biochemistry [DS66]. biochimique [WP63]. Bioelectric [BT62]. Bioelectrochemical [Low69]. biogenesis [McM63]. Biol [Ano66f, Ano69g]. Biological [UE65a, UE65b, Win67, BR66, Ber64b, Cin69, CC65, Cze67, DKGK66, Els62, Enr65, Est66, Eyl66, GH67, GdlB66, GM68, Koc62, Koc68a, Koc68b, Lin61, MH69, Mor66, Myh69, Pet64, Pie66a, Pie67, RTB67, RS62, Ros68, Sam66, Seg69, Sha61, Sim65b, Sne66, Ste64, Sti68, TR63, Tol63, Vai64, VP64, Vai66a, WW68a, Wal69a, Wal69b, Wil69, ZS66]. biologiques [VP65b]. biology [AC63, Ano65n, CM62, IJ68, Koc66, Koc69, SG61, ZZ67, ZZ69]. Biomathematics [Ano63b]. biomolecules [Yam66]. biopolymer [Man69b]. biosphere [RM68]. biosyntheses [Low61]. biosynthesis [RB66]. biosynthetic [Sol66a, Sol66b]. Black [Tie67, HHT65a]. blood [Eva62, EDT66, Per62b, PH66, Spe67, WZ69, Zaj68]. blue [Wil61b]. Board [Ano61a, Ano62a, Ano62b, Ano63e, Ano63f, Ano64b, Ano64c, Ano65c, Ano65d, Ano66i, Ano66j, Ano66k, Ano66l, Ano67d, Ano67e, Ano67f, Ano67g, Ano68a, Ano68b, Ano68c, Ano68d, Ano69c, Ano69d, Ano69e, Ano69f]. body [FM69, LM66, Per62b]. Boer [De 67]. Boltzmann [MM68a, Mac69, She67]. bond [RM68]. bonding [Loe68, PR66]. bones [Cur68]. borders [Pas67]. both [Nao68]. boundary [ERH69b, Mel64a, Mel64b, Mel64c]. Bounds [RKB67]. Boyle [Nob69]. braid [Shi66c]. brain [Ber64c, Deu63, Gar62]. branching [Lin68b]. break [Bla67b, Wol62a]. break-up [Bla67b]. bridge [Koi66, Koi67]. bridges [Con69]. brief [Ser64]. brush [Pas67]. building [Ber63]. Bulliard [PP68].

c [KS68b]. calcium [HR66, WW68a]. calculating [Car64, Lau65]. Calculation [JI69, SZ65, GFR68, Shi63]. Calculations [Shi66a, BS69]. Camin [Est68]. can [Bra65]. cancer [Ser64, TS69]. cancerous [Con69]. capacitance [EH68, HHT65a, HHT65c, Seg67]. capacity [Tom66a]. Capillary [Wil66]. Caplan [Ano66f]. capture [Bas63]. carbamates [Heb68]. Carolina [Ano63b]. carotenoids [PLB69]. Carrier [RW63, Bri66, EKH66, Vid66]. cases [Boe63, Sem67]. castaneum [Daw66]. catalysed [Ain68, DPW66a]. catalysis [Lov69, Tho68]. catalysts [Wil61a, Wil62]. catalyzed [Dia69, Koc67a, Mag69b]. Cation [Dun65]. Cation-permeability [Dun65]. Cations [Wya64]. caused [LMB68]. Cell [O’B67, Wei62, Bel61, Bel62, Ber62a, Ber62b, Ber65, Ble69, DG63, Dic64, GG67, Hel68a, Hel68b, HE65a, HE65b, Kau67, Kie68, KK64, LR69, Loc65, Luc64, Mel64a, Mel64b, Mel64c, MM68b, MR63, Rob68, Rot63, RS66, Sai68b, Sai68a, Sha65, SB65, Sim65c, SR66a, SR66b, Sim67, Sta65, Ste68, Tak66, Tak68, TM68, Tru67, TB68, Wat65, Wei63, Wei64, Wei68, Wil67, Wya64, YSB65, YP64, Zaj68]. cell-culture [HE65a]. cells [Bro69, Bro63, DSB68, Doe67, Doe69, Dun65, Ehr67, Geo69, Had67, Koi66, Løv63, Sag67, Ste68, VNM69, WZ69, Zaj68]. Cellular [Boy62, Wei64, Aid66, Bel61, Doe68, FBG^{+}66a, Gil68, Gin67, Gri68a, Gri68b, Gro69, Lin68a, Lin68b, Res69, SR66b, Sta67, Sug66, TS66, Tyl66, Wad65, Wol69]. centering [Beg63]. central [Cla64]. cerebral [Lau65, Lau69a, Lau69b]. certain [Boe63, Wal69a]. chain [Kli66, PLB69]. chainlength [SZ65]. chains [Ebe69, Wil61a, Wil62]. champ [WP63]. Change [Shi67a]. Changes [LMB68, BCB64a, Gin68, JW67, Tol62]. changing [GGW68, Sai68b, Sai68a]. character [Acl67]. characteristic [Ano66f, Cap66a, RM68]. Characteristics [TI67]. characterization [ZES68]. charge [FBG66b, FBG66c]. charged [Mac69, Vai65b, Vai65c, Vai66a]. check [KM65]. Chemical [Cze68a, Cze68b, Cze68c, Wan63, Ber64c, Bri69, Cze64a, Cze64b, Cze67, FBG66c, Gro69, Hig68, KSK65, Pla61, Ros68, She67, Sne66, SS67b, SS67a, Sug63, SF63, Sug66]. chemo [Dun63]. chemo- [Dun63]. chemostat [Han69, RS67]. chemotherapy [BBKL66]. chimaerism [RG69]. chlorinated [CRJ66]. chlorophyll [Bro64]. choice [Coh67, Pri67a]. chromatid [Hed69]. chromatids [Kau67]. chromosome [SP66, SP69, Shi66c, SH64, Wol62a]. chromosomes [Kil66b, Par64]. chronon [ET67]. chymotrypsin [Kno65]. cilia [RB68a, SS64]. Circadian [Swa69, ET67, PZO68, Pav69, Wev66]. circuit [CS65, GH67, Sug63, SF63, Sug66]. circuiting [Reh68]. circular [JT69]. circularity [RR68]. circulation [She62b, SU69, WO67]. cis [PLB69]. classical [BLT^{+}62b, BLT62a]. classification [Est66, Fis62c, FT68, Orl68, Rub67]. clearance [Per62b, PH66]. cleavage [Rap65]. clock [ET67]. closed [Wal65]. clostridial [Per62a]. clustering [Sar65]. CNDO [BS69]. CNDO/2 [BS69]. co [Gre66a, Ric65, Sch68]. co-factor [Gre66a]. co-ordinates [Sch68]. co-orientation [Ric65]. code [Her62, Mar67, QZ62, RR67]. coding [Blu62, Chi65, Jeh67]. coefficient [Dic64, Ged68, Shi66a, Shi66d, Shi67a, Shi67c]. coefficients [CC65, Lau65, Lau69a, TM68]. coexist [RS67]. coexistence [PP67]. Cohesion [Bro63]. coil [MM68a, Par69]. coiled [Par69]. coiled-coil [Par69]. coincidence [TR65]. Coleocheate [Kor69]. coli [Eng62, SH64]. colicin [Hed66]. colicinogenic [Fre63]. colicins [CT67]. collagen [Dav66]. Collection [Tru67]. collections [Pie66a, Pie67]. collective [Sha65]. College [Ano63b]. Collision [Man69a, Pol63]. colloids [De 67]. colloquium [Ano65n]. colonies [Wal65]. colony [Wen61]. colour [Wil61b]. columnar [Lau69b]. combined [SF63]. commaless [Her62]. comment [SW66, Tri69, vdS69]. Comments [Agi69, Fel68, Goe67a, Hug68, IJ68, ZZ69, Agi63, Agi64, Ros68, Vai66a]. common [Boy66, BBS63]. communication [Alt65, BW63, Bos68, Mar61]. community [Bla66]. comparative [Ain68]. comparison [Kos62]. compartment [Myh69]. compartmental [JM66]. compartmentalization [MLH66]. compatibility [Gof66]. compensatory [Sam66]. competent [AAB69]. competition [Sim67, Wue69]. Competitive [WA65]. complementarity [Els69]. Complements [VP65b]. complex [Cen68, GC68, KT65, Sch68, SS66, Ste68]. complexes [Cin69, PA68, Whi65]. complexity [Ter65]. component [Sav69a]. components [Kil66a, Mae65a, Shi66d, Shi69b]. composed [BCB64a]. Composition [O’B67, Ata66, Fre62a, GD63, Gat64, Goe67a, HHT65a, Hav66, PA68, RH69, RM68, RB68b, Smi66, Ata66]. Compositional [De 69]. compounds [Fel68, JI69, PLB69, SW67]. comprehensive [Mar64]. computational [BBKL66]. Computed [Gin68]. Computer [Day65, Gar62, ST68, WC69, Hei64, PCB68, Pen69, Pri67a, Pri67b, Pri67c, RP68, RD69, Sta67, Sti68, SF63, TS69, War67b, YP64, Zaj68]. Computing [Her69, Sug66]. Concentration [RB68a, Kat69, MM66b, NS61, Wag68]. concentrations [Ged68]. concept [ET67, Els61, Sam66, Ser64]. Concepts [MM68b, Ber61a]. concerned [YP64]. Concerning [Bra68, Bri69, Rig66, Rot63]. condition [Rus68b]. conditions [Coh67, Noa68, RS67, Tyl66]. Conductance [Vai69, HHT65c]. conduction [Jah62, Ste68]. configuration [CF69]. conformation [LMB68, War64]. conformations [Cen68, Cen69, TP67]. confusum [Sta66]. conjugation [Wu67]. connection [Hel68a, Hel68b]. considerations [DKGK66, Man69b]. considered [vO63]. consistant [WP63]. consistency [Lau69a]. constant [Fri69b, Ged68, MH69, Ohk68, Zel68]. constants [CFY68, FCR67, Mag69b]. constrained [RB68a]. constraints [KS67]. constructed [Kau69]. consumption [Fuj63]. contact [Doe68]. contacts [Wei64]. contain [Les69]. containing [CGTK67b]. content [BW69, Gat66, Gat68, Hav66]. Contents [Ano63c, Ano63d, Ano64a, Ano65a, Ano65b, Ano66g, Ano66h, Ano67b, Ano67c]. continuous [DH69, Has66a]. continuous-time [Has66a]. continuously [Sai68b, Sai68a]. contraction [Arn61, Cor65, Loe68, Ull67]. Contrast [Alm69]. Contribution [Smi67, Blu62, Har63]. contributions [Bri69]. Control [Res69, CHS67, DS68, Dun65, Gri68a, Gri68b, Gro69, Kau67, Koc67b, Mil62, RTB67, Rob68, RR66, Sha65, Sug66, Wya64]. Controlled [Dan64, Kie68, Lin63, Wal69a, Wal69b]. convenience [Pet64]. conversion [Bas63, Cal61, Cap66b, Cap66c]. converters [Ano66f, Cap66a]. cooling [Hen68]. cooperative [ZS66]. copolymeric [Goe67b, GM69]. cornea [YS67]. corphyrins [HPW62]. correcting [ADB61]. Correction [Fri69a]. Corrections [Agi64, Lea65]. correlates [Psc63]. correlating [Pul62]. correlation [Coh67, Hav66]. correspondences [HM68]. corrin [JI69]. cortex [Lau65, Lau69a, Lau69b]. count [Fri69a, LR69]. counter [Vid66]. coupled [KSK65, She67, Win67]. coupling [Cap66b, Cap66c, Lau65, Lau69a, Res69, RPS61, UE65b]. covalent [Loe68, RM68]. crassa [Ber64d]. Crick [PR68b]. criteria [AC63, Bre67, Vai65a]. Critical [Bla66, Rap68]. Cross [Shi66b, Shi64, Shi66c, Sie69]. Cross-over [Shi66b]. cross-overs [Shi64, Shi66c]. cross-reacting [Sie69]. crossing [Ghe66, RTB67]. crossing-over [Ghe66]. cryptogram [Eck62]. cube [TI67]. culture [Cin69, DH69, HE65a, HE65b]. cultured [DSB68, MM68b]. cultures [War67b]. Curie [Acl66]. current [GH67, Kim66]. curve [GK69, Pri67b]. Curve-fitting [Pri67b]. curves [Ber62a, Ber62b, Ber65, GFR68, HvR69, Spr67, TB68]. cut [DSB68]. cut-and-patch [DSB68]. cutaneous [Str69]. Cybernetics [AW65]. cyclase [Mad69]. cycle [DS66, Hor69, Tak66, Tak68]. cycles [Mor66, Swa69, Wil69]. cyclic [Cze68a, Cze68b, Cze68c]. Cycliques [LL62]. cyclohexene [PLB69]. cylinder [CGTK67a, CGTK67b]. cytochemistry [O’S62]. cytochromes [KS68b]. cytosine [BS69]. cytostatic [Sim65a].

D [Fis62d, BBS63, Sin65]. dairy [RKB67]. dark [Gau64]. darmdurchblutung [WO67]. Darwin [Lei69b]. Darwinian [Bla67a, Lur69]. data [BR64, Dar69, HE65b, Koc62, Koc68a, Myh69, Par67, QZ62, RTB67]. December [Ano66v, Ano66r, Ano67v, Ano68q, Ano69q]. deducible [PL67]. defense [Dix68, Ozo69]. definition [Ber61a, Des65]. deformability [Wei68]. degradation [Coh69]. degree [Cap66b, Cap66c]. degrees [PP67]. dehydrogenase [BLT^{+}62b, Per62a, RB68b]. dehydrogenases [Whi65]. dehydrogenating [Whi65]. denaturation [Bra68, Urb68]. density [Gar67b, Mel64b]. density-dependent [Gar67b]. deoxynucleotide [Rot63]. deoxyribonucleic [Lin61]. dependence [Koc68b, Wal66b]. dependency [NN66]. dependent [Gar67b]. depolarization [Dun65]. derepressed [Sim67]. derivation [RP68]. derived [Lew66, MM68b, Sar65]. descent [Shi66b, Shi67d]. described [Shi66c]. describing [Ain68, Gof66, HR69c]. description [Ros68, dLR67]. descriptions [Lin68c]. Design [Nel63, Alm69, BLT^{+}62b]. detailed [Gra64]. detection [Fis62b, PP67, Sat67]. Determination [BR64, CFY68, HE65b, Koi67, Ohk69, Vay65, Ber64a, BA63, CC69, Cla64, DPW66a, Day65, FCR67, Fri69a, Koi66, Mag69b, Per62b, RB63, Tie67]. determinations [RB68b]. determined [Bro64]. determining [JD69]. deterministic [Eng68]. developing [GC69]. Development [DP69, AW65, Arb67, Daw66, Kor69, Lin68a, Lin68b, VNM69]. developmental [Her69, IJ68, ZZ67, ZZ69]. diagrammatic [Hil66]. diagrams [Lov69]. dicarboxylic [Haa69]. Dielectric [Ohk68]. difference [Fis62b]. differences [Kau67, Nas69]. different [Pie66a, PP67, Pie67]. differential [BR66, Cal68, YP64]. Differentiation [Gil68, Kau67, Rin65, Sim65c, Wol69]. difficult [BNW62]. diffuses [O’S62]. diffusion [Dic64, EH69, ERH69b, FW65, Gol69, Pas67, SKB69, TM68, Vai65b]. Diffusional [GB62b]. digestive [WO67]. digital [Pri67a, Pri67b, Pri67c, RP68, RD69, SF63, War67b]. Dilemma [Lei69b]. dilution [EDT66]. dimensions [FBG66b]. dimeric [Ain68]. d’intérêt [WP63]. diphosphopyridine [Gra64]. diploid [Shi66a, Shi69a]. dipole [FBG66b, FBG66c]. dipoles [Ghe66]. Directive [Why67]. discontinuous [Nor68]. discrimination [Boy62]. discussion [Hel68b, Mae65a]. disease [Fre62b]. diseases [Gof66]. dissociable [CGTK67b]. distinction [Dia69]. Distortion [SP69, Par66, Par68]. distribution [BW69, De 69, GM69, HE65b, Koc66, Per62b, RM68, SP66, SZ62, SZ65]. distributions [FBG66c, Hol66b, LR69, Noo68, Koc69]. disulfide [Lev62]. diversity [Pie66a, Pie66b, Pie67]. dividing [Noo68]. Division [Wen61, Bel61, Psc63]. DNA [AAB69, CF69, De 69, DV66, Erh69a, Fon67a, Fon67b, Fre62a, GD63, Gat63, Gat64, Gat66, Gat68, GC68, Goe67a, Goe67b, GFR68, GM69, Kal68, Kau67, Les69, Lew67, MM66b, MM66a, Nao68, Pai62, Phi69, PR68a, Pol61, Sib64, Sto65, Zár69]. DNA-dye [GC68]. Do [DSB68]. documents [ZP65]. does [GH67, Les69]. dog [EDT66, HR66]. dominance [Cro63]. dosage [Ano69g, Har63, KT66, KT69, Nel62]. dose [GK69]. double [Fis62b, McM63, Phi69, Sch62]. double-helical [Phi69]. doubling [HE65b]. doubling-time [HE65b]. down [Pic66]. drastically [Kom63]. drop [GG67]. Drosophila [CB61, Car64]. drug [Ano69g, Hom67, KT66, KT69, Nel63, Wag68]. dual [RB68a]. Duplication [Ble69, SH64, WW68b]. duration [Doe68, Wev66]. during [Cla67, DD66, Erh69a, EH69, Hen68, Koc64, McM63, Nel62, Phi69]. Duval [Sta66]. dye [GC68]. Dynamic [KP67]. Dynamical [Fon67a]. dynamics [Ber61b, Ber61a, Ber64a, Ord69, Wil67].

E. [Ano69g, Eng62]. early [PK68]. earth [Mar64, Yam66]. ecological [Gar67a, Gar67b, KP67, Pie66b]. ecology [AB61, BKK66, Mae65a, Mae65b]. editor [Agi69, Fis62d]. Editorial [Dan67a, Dan67b, Ano61a, Ano62a, Ano62b, Ano63e, Ano63f, Ano64b, Ano64c, Ano65c, Ano65d, Ano66i, Ano66j, Ano66k, Ano66l, Ano67d, Ano67e, Ano67f, Ano67g, Ano68a, Ano68b, Ano68c, Ano68d, Ano69c, Ano69d, Ano69e, Ano69f]. Effect [Gar67a, GB62a, Shi67b, Sin65, Gar67b, KT69, Lad64, Myh69, NS61, Shi64, Shi66b, Tol62]. effective [Bro64, Wil69]. effectiveness [HR66]. Effects [Cen69, Ain68, Atl68, BDF69, DKGK66, Fru65, Ids68, PZO68, RTB67, SZ62, Wei68, Cen68]. efficiency [Cap66b, Cap66c]. eggs [Rap65]. einfaches [Hel64]. ejection [Zár69]. electret [Wob68]. electric [Vai66b]. Electrical [EH68, Seg67, Deu63, Fru65, HHT65b, HWJ69, PR68a, Ste68, Vay65]. electroencephalography [Dew64]. electrolyte [Lev66, Vai65b, VG68, Vai69]. electrolytes [Pas67, Vai64, VP64, Vai65c]. Electron [Mur66, Mal65, Pul62, UE65a, UE65b]. Electronic [BL69, PVP61, BS69, CRJ66, Jah62, JI69, Per62a]. electronique [VP65b, WP63, VP65a]. electrophoresis [GB62a, Mel64c]. Electrophoretic [WOB62a]. Electrostatic [VG68, GC68]. Electrostatics [FBG^{+}66a, FBG66b]. elements [SR66b]. elongation [Loc65]. elution [HvR69]. Emerson [Tol62]. emission [Bro64]. empathy [HR69b]. empirical [GFR68, Haa69, IKTN66]. empiricism [Sok62]. encoding [Par66, Par68]. endemicity [Bla66]. endings [KS68a]. endomessenger [Had67]. endothelium [SKB69]. energetics [PR68b]. energies [GFR68]. Energy [Acl69, Bas63, Ano66f, Atl68, BL69, Cap66a, Cap66b, Cap66c, Cla63, D’A63, FBG66c, Hel68a, Hel68b, LM66, Nas69, O’N69, Res69, TR63, Wol62b]. energy-rich [BL69]. enlargement [Koc64]. entrainment [Han69, Swa69, Wev66]. Entropieerzeugung [Hel64]. entropies [GFR68]. entropy [Wil66]. Enumeration [Fis62c, HM68]. Environment [Mae65b, Coh66, Coh67, Hel68a, Hel68b, Lev64, Mae65a, Sai68b, Sai68a]. environmental [Atl68]. environments [Bro62]. Enzymatic [Bla67b, Sav69a]. Enzyme [Kno65, Ain68, BLT^{+}62b, Cen68, Cen69, Cze68a, Cze68b, Cze68c, DPW66a, DPW66b, Dar68, Dia69, FCR67, Fis62a, Haa69, Hei64, HH69, Kat67, Kat69, Koc67b, Lov69, Mag69b, O’S62, PL67, RB68a, Res69, Sem67, Wal66c, Wal67]. enzyme-catalyzed [Dia69, Mag69b]. enzyme-substrate [Cen68]. enzymes [EMM68, Lin63, Sch62, Wya64]. enzymic [Cen68, Cen69, DS67, Kos62, RP68]. epidemics [RKB67]. epigenesis [Kau69]. epithelial [Pas67]. equation [Bar65, BBKL66, Fri69b, Ged68, JD69, Kat67, MM68a, Mac69, WA65, Zel68]. Equations [Bur66, Agi63, Agi64, Bri69, Cal68, Dar69, Lau65, Lau69a, Sem67, YP64]. equilibrium [CFY68, RM68]. equivalence [PEC69]. equivalent [Sug63, SF63, Sug66]. equivivant [Tru67]. Errat [Ano64d]. Errata [Ano64e, Ano64f, Ano67h, Ano67i, Ano68e, Ber62b]. Erratum [Ano63g, Ano65e]. error [Fri69a, Myh69]. errors [Noo66]. erythrocytes [LL69, RPS61]. Escherichia [SH64]. estimate [Bro64]. Estimation [Spr67, TM68, CB61, O’N69]. Etude [BP62, WP63]. evaluation [Koc62, Koc68a]. even [Zár69]. Événements [LL62]. events [Car64]. Evidence [Par67]. Evolution [Hol68, Lew61, Ata66, Cro63, DD66, Els63, Fre62a, Ham64a, Ham64b, Kil66b, Lei69b, RR68, RTB67, Spe64, WW68b, Why67]. evolutionary [Bla66, Cal68, Hol66b, ZP65]. exact [KT69]. exactly [Koc66]. examination [Hed66, Par67, PR68b, RPS61]. exchange [CGTK67a, CGTK67b, ERH69b, Løv63, SP69]. exchanger [Seg67]. exchanges [Hed69, SP66, Wol62a]. excitability [MR63]. excitable [Dun65, JH69]. excitation [Kno68]. Excitatory [Dun63]. excited [BS69]. excretion [Mil66, Nel62]. exist [Coh67]. existence [GM69]. exo [BLT^{+}62b, BLT62a]. exo-alkylating [BLT^{+}62b, BLT62a]. expanding [LP66]. experiment [Noo66]. Experimental [RPS61, Cha66, FM69, MR68, Cze68c]. experiments [She62b, SU69]. explanation [Boe63]. exponential [Van67]. exponentially [TB68]. expression [Els69, Fis62a]. extended [Lew66, RFCH68]. extension [Bel61, QZ62]. Eye [Beg63]. Eye-centering [Beg63].

Facilitated [Sne65, Gol69]. factor [Gre66a, Kau67, Ros67, Sin65]. factors [Atl68, BLT^{+}62b, Ber64a, Fre63, Kor67, RB68b]. false [Hed69]. families [Lei68]. fat [BBS63, PF67]. fauna [PP68]. feature [Gre66b]. February [Ano66t, Ano67p, Ano68m, Ano69o]. feedback [DS68, Gri68a, Gri68b, HR66, Kie68, Mel64a, TM68]. feedback-controlled [Kie68]. female [Hor69]. ferroporphyrines [VP65b]. fertilization [PK68]. fibers [Pic66]. fibrillar [Dav66]. fibrous [Par69]. Field [Dan64, Fri69b, Ged68, MH69, Zel68]. fields [Hol66a, Wad65]. filament [RB69]. filamentous [Her69]. filaments [Lin68b, Lin68a]. films [CC69, VG68, Vai69]. find [BNW62]. Fine [CHS67]. finite [BCB64a, FBG66b]. fishery [Pel69]. fitness [Bod63, CB61, Lev64]. fitting [Haa69, Pri67b]. fixation [Pra62]. fixed [CS65, FBG66c, FBG66b]. fixed-rotating [CS65]. fixing [DP69]. flagella [Doe66, Jar64]. Flavin [Gra64]. flocculation [WOB62a, WOB62b]. flour [Sta66]. flow [Bri69, Cap66b, Cap66c, EKH66, Eva62, EH69, FW65, Kli66, Mel64a, Mel64b, Mel64c, PGH63, Per62b, PH66, Sch68]. flows [EDT66]. fluctuating [Swa69]. fluctuations [Mor69]. fluid [YS67]. fluids [Cin69]. fluorescence [Cla67]. flux [Ess66, Vid66]. Fluxes [Bri66, Hil66, O’N69]. fold [TR65]. follicles [Cla64]. following [WZ69]. food [Ebe69, O’N69]. Force [Ell68, Ano66f, Cap66a]. Force-balances [Ell68]. forces [Doe69, Kor67]. foreign [Boy62, FM69]. form [Dar69]. Formal [Ano69g, KT66, KT69, KS67, KS68b, WO67]. formale [WO67]. Formalized [dLR67]. formation [FM69, Kor67, Lur69]. formulation [Hel68a]. Founding [Lei68]. four [Cze68b]. four-step [Cze68b]. fractionation [Mel64a, Mel64b, Mel64c, dD64]. Franck [Ros64]. free [Mel64a, Mel64b, Mel64c]. frequencies [BCB64b, Gat63, Wu67]. frictional [DG63, Ess66]. Fries [PP68]. frost [Lev62]. Function [O’B67, Ata66, Hig68, Lin61, Mar64, She67, SS67b, SS67a, Vai69, Wal66b, Wal66a]. Functional [Sug63, SF63, Sug66, DL69b, GD63, Gat64, Goe67a, Sim67, Tri69]. functions [BR64, Blu68, FBG66c, Tru67, Wil61a, Wil62]. fundamental [Ber61a, Gre66b]. Further [BLT^{+}62b, Tak68].

games [Lew61, War67a]. Ganghöhen [Jar64]. gap [Koi67]. gas [Ohk67]. Gatlin [Goe67a]. gene [CS66, Daw66, Gri68a, Gri68b, Kau67, Lin63, Ren62, Shi66a, WW68b]. gene-controlled [Lin63]. gene-set [Shi66a]. general [Est68, Hel68a, Hel68b, Wal66c, Wat65, Cze68a, Cze68b, RW63]. generalization [De 67]. Generalized [Shi67c, DL69b, Els69, Shi66d, Tri69]. generating [Koc66]. generation [Fri69a, Kal68, Lew62]. Generator [Pri67c]. genesis [Tab67]. Genetic [RR68, BCB64a, Fra62, FBM66, Gat63, HM68, Kau69, Nao68, RR67, RR66, Ser64, Ter65, Wu64]. genetical [Bar62, Ham64a, Ham64b]. genetics [Has66a, Hol66b, WGP66]. genotype [HM68]. genotype-phenotype [HM68]. genotypes [BO65]. genotypic [BCB64b]. Geometrical [Rap65]. geometrically [RB68a]. German [Jar64, Kli66, WO67]. gestation [Fei67]. gland [DS68]. glandular [DiS69]. Globular [Luc64]. glucose [Pol63]. glucosuria [WGP66]. glutamate [Gar62]. glycoproteins [Eyl66]. go [Pla61]. Goldman [Bar65, MH69]. gradient [Mel64b]. gradients [Doe67]. Grain [LR69, Fri69a]. grain-count [Fri69a]. graph [Bro69, Est66]. Graphic [Van67]. graphical [HvR69]. great [Pen69]. green [KCE69]. group [HHT65b, Shi63, Shi66c, Shi67c, UE65a]. groups [Pla61, Rub67, Sar65, Spr67]. growing [GB66, GGW68, Koc68a, TB68]. growth [Blu68, Bro63, Fuj63, HE65a, HE65b, Koc64, PCB68, Spr67, Sta66, SAFT66, TI67, Wil67, ZS65]. growth-rate [HE65b]. guanine [BS69]. guanine-cytosine [BS69]. guessing [War67a].

H [She67]. H-theorem [She67]. halogenoethylamines [KMT65, KT65, Tri64]. halving [Fri69a]. Hamaker [De 67]. hamsters [Cha66]. haploid [DD66]. hapten [WA65]. hapten-antibody [WA65]. harmonic [Mar67]. heart [Spe67]. Heat [Per62b, PH66]. Helical [CF69, Fon67a, Hav66, Phi69, TP67]. helix [MM68a, McM63]. helix-coil [MM68a]. hemoglobin [Sne65]. hemoglobins [Ata66]. hemolysis [WZ69]. herds [RKB67]. hereditary [RTB67]. heredity [Pat67]. hermaphroditism [Smi67, Tom66b]. heterogeneity [AP65, BA63]. heterogeneous [Kal68, Lev64]. heterozygote [Bod63]. heterozygotes [Fis62b]. hexagonally [Ell68]. hexagonally-packed [Ell68]. hibernating [Cha66]. hibernation [Hen68]. High [Mal65, Boe63, D’A63]. higher [Arn68]. Hill [Ano66f, Cap66a]. histograms [Bra65]. Histological [Lau65]. histology [She62a]. histone [Bae64]. history [Hol67, ZP65]. hit [Hed66]. Hodgkin [Agi63, Agi64, Lew66]. Hoff [Nob69]. home [JT69]. homeostasis [Sim67]. homologous [KS68b]. homology [De 69]. homozygosity [Shi66b, Shi67d, Wal65]. honey [Wen61]. hormone [DiS69, Mad69, Rig66]. host [BNW62, Gre66b]. hosts [BNW62]. Hückel [RFCH68]. Human [Wil61b, CHS67, CS68, Eml66, Fan69, HR69c, HR69a, Kom63, LL69, MM68b]. Huxley [Agi63, Agi64, Lew66]. hybrid [Sug66]. Hydration [YS67, Lew67]. hydrocarbon [HHT65a]. hydrodynamic [Mel64a]. hydrogen [PR66, Res69]. hydrophobicity [Big67]. hydroxylamine [NM68]. hydroxylation [Sol66a]. hyperbolic [Gat64, Goe67a]. hypochromicity [HL64]. hypophysial [DS68]. hypothalamo [DS68]. hypothalamo-hypophysial [DS68]. hypotheses [Hed66, Mil66]. hypothesis [BBS63, CC65, Dou62, Gra64, Had67, HKG68, Lev62, Lin63, PR68b, Rig66, Rot63, RPS61, SP69, Sim65a, Tab67, Wei62, Yam66]. hypotheson [Bar62].

icosahedron [Han67]. idea [AB61, Sug63]. idealized [FBG66c]. ideas [Atl68]. identification [DS68]. II [Alt65, Ber61b, Ber62c, Ber64d, Cen69, Cze64b, Cze68b, FBG66b, Gat68, Gri68b, Ham64b, Hel68b, Koc68a, Koc69, KT69, KS68b, Lau69a, Lin68b, Mel64b, Pri67b, QZ62, Sav69b, SU69, SR66a, Sug63, Tak68, UE65b, WOB62b, Wil62]. III [Cze67, Cze68c, HK66, Lan63, Lau69b, Lev64, Mel64c, Pri67c, SR66b, SF63]. illustrating [Tyl66]. imidazole [UE65b, UE65a]. immaturity [Len62]. immune [Sie69]. immunity [Bur66, For69]. immunoelectrophoresis [AKT62]. immunological [Fit64, Len62]. impact [Lei69b]. impermeable [vO63]. implanted [Kor67]. implication [Bla66, Cze68c]. implications [Mac69, WW68b]. imposing [Gar67a]. impulse [NN66, Pic66]. inbred [Wal65]. inbreeding [Shi66a, Shi66d, Shi67a, Shi67c]. inception [Sat67]. include [UE65b]. incompatibility [Lew62]. incorporation [PS68]. increase [Gar67b]. increasing [Kom63]. Increment [Gau64]. independence [Seg69]. independently [GM69]. Index [Ano62c, Ano63h, Ano63i, Ano64g, Ano64h, Ano65f, Ano65g, Ano66m, Ano66n, Ano67j, Ano67k, Ano69h, Ano69i, Ohk68, Shi69a]. Indicator [EDT66, PEC69]. indices [Shi69b]. indigenous [Boy62]. Indirect [O’N69]. individual [Shi69a, Wal66a]. individuals [BNW62]. Induced [MR63, Eng62, Hei64, Ids68, JW67, LL69, Mal65, SP66, SP69, Tol62]. induction [Len62, Lin63, PL67]. Inertial [Brl64]. infectious [Gof66]. infinite [Cal68]. inflection [Wal67]. Influence [EH69, Hed69, Kat69, RH69, HHT65a, Koc62, MM66a]. Information [Orl68, Atl68, BW69, Gat66, Gat68, Had67, Lei66, Lei69a, Nao68, Spe64, Wol69, AW65, Kli66]. information-transmission [Spe64]. Informationsfluß [Kli66]. infrequent [PP68]. Ingraham [Kat67]. inheritance [Boy66, Fis62c]. Inhibition [Vid66, Cen69, Haa69, Sem67]. inhibitor [BLT62a]. inhibitors [BLT^{+}62b, Kat69, SW66]. initial [HH69, JD69, Wei64]. initiation [Ble69, Bur66]. initiation-rates [Bur66]. injection [NS61]. injury [DSB68, Koi67, Lev62]. inlet [PEC69]. Inorganic [Cin69]. Inorganic-organic [Cin69]. input [PEC69]. inputs [JM66, Lin68a, Lin68b]. Insect [Daw66, Mil62, PP68]. insects [Mil66]. insoluble [PA68]. instability [Tyl66]. insular [Bla66]. insulin [Kla66]. intact [Wag68]. integrated [JD69]. integration [Pri67a]. integumental [Bue67]. integuments [Bue66]. interacting [CGTK67b, Pav69, PCB68]. Interaction [BNW62, Ain68, De 67, Ess66, Gin68, KMT65, Koc67c, OF67]. Interactions [Par64, CPC66, GC68, Gin67, Lew69, Lin68a, Lin68b, VP66a, VG68]. intercalation [GC68]. intercellular [Con69, Gin67, Koi66, Koi67]. interdependent [Wal66a]. interest [Cal68]. interfacial [Kor67]. Interference [Shi64, Boe63]. Interferometry [Tol63]. interferon [GK69]. interior [Dic64]. Interlayer [OF67]. Interlayer-interaction [OF67]. intermediary [Kla66]. intermediate [DPW66b, DS67]. intermediate-one [DS67]. intermediates [DPW66a, Mag69b]. Internal [PK68, Ged68]. interpretation [Ber62a, Ber62b, Boy66, Lad64, Tol62]. Interrelated [SR66b]. interval [Bra65]. intervals [NN66]. intra [Cor65]. intra-amniotic [Cor65]. intracellular [Coh69]. intragenic [Ber62c, Ber64d]. intravenous [NS61]. Introduction [NO66]. Inverse [BKK66]. inversion [BBKL66, FBM66]. invertebrate [Rap65]. Investigation [Myh69, Dan69, Wev66]. involved [Kor67, Pul62, Wer65]. involving [Cen68, Cen69, Vid66]. Iodoacetamido [BLT62a]. Ion [KK64, Kim66, Dan64, Ged68, Mil66, Rob68, Seg67, VP66a]. ion-exchanger [Seg67]. Ionic [Lew69, Fru65, Lin67, MM66a]. ionizable [Sar65]. ionization [Sha61]. ionizing [Atl68]. ions [Dan69, Jah62, Vai66a]. iron [Jan64]. Irreversible [DG63, Acl67, BLT62a, Ber64b, HK66, Hil66, Loc65]. Island [KCE69]. Isolated [Cze64a]. isomerism [PLB69]. Isotope [Ess66, CGTK67a, CGTK67b, Mal65]. isotropic [Rap68]. Iterative [RFCH68]. IV. [Hil66, PA63, RP68, Sim67, Sug66, War64]. IX [Sta66].

J [Ano66f, Ano69g]. James [Ros64]. January [Ano61b, Ano62e, Ano63j, Ano64i, Ano65i, Ano66p, Ano67l, Ano68i, Ano69k]. July [Ano61d, Ano62d, Ano63k, Ano64j, Ano65h, Ano66u, Ano67m, Ano68f, Ano69s]. June [Ano67u, Ano68n, Ano69r].

keeping [KM65]. keratin [SS64]. kettenmolekülen [Kli66]. kidney [FW65, PH66]. kinase [RB68a]. Kinetic [CGTK67a, CGTK67b, PS68, BR64, Ber64c, Cze68a, Cze68b, Dar69, Dia69, Haa69, Koc62, Koc68a, Pri67a, Pri67b, Pri67c, RP68, SS67b, Cen69]. Kinetics [Fuj63, GL65, Goe67b, Gol69, Koc67a, LL69, Nel62, Wag68, WZ69, Cen68, Cen69, DPW66b, Dar68, Gar67c, Hed66, KSK65, PL67, Pol63, Wal67, WO67, Wol62a, ZS65, ZS66, RS66, SR66a]. kinetik [WO67]. king [Shi67c]. Klebsiella [DH69]. Krüger [Ano69g].

labeled [Bro69, LR69, PEC69, TB68]. labeling [Bro69]. labelled [PS68, Tak68, Tak66]. labile [Koc62]. labor [Wen61]. lack [LM65]. lactic [BLT^{+}62b, RB68b]. Laminar [Mel64b]. lanosterol [Sol66b]. lanosterol-possible [Sol66b]. Laplace [BBKL66]. large [Pla61]. lattice [EKH66, Ohk67]. law [Els61, Hel68a, Hel68b, Hen68, O’S62, Sav69a, Sav69b]. laws [RP68]. layers [EH69, ERH69b]. Lea [Hug68]. leaf [Boy66]. leaflets [HHT65b, HT63, Par67]. learned [Bra65]. Learning [Deu63, Aid66, Mal63, RTB67]. Lebewesen [Hel64]. lecithin [HHT65b, Par67]. legumes [DP69]. Lethal [Sim65a]. Letter [Fis62d, Agi69]. leucocytes [WOB62a, WOB62b]. level [CT67, Gra64, dF62]. Lewis [Fis62d]. Liapunov [Hig68, She67]. liberator [Kel67]. life [CS66, Cin69, Coe68, Fit64, Fon68, Gil64, Hol67, Kom63, MM68b, Pat67]. life-span [Fit64]. lifetimes [Man69a]. light [DKGK66, FR64, Koc68b, LMB68, Mil66, Swa69, Tol62]. light-induced [Tol62]. limit [Hol66b]. Limitations [HvR69, Hul69]. limited [SP69, WGP66]. limiting [Ber61b, Sin65]. limits [Gar67a]. Line [GG67]. linear [Ano66f, Cap66a, Haa69, LP66, MM68a, Van67]. lines [BCB64a, Cha66, Wal65]. linkage [Ben63]. linkages [Lew69]. lipid [CC69, Dan66, EH68, EH69, ERH69b, Gre66b, HHT65a, HHT65c, HH66, HT63, Luc64, MR68, Ohk67, OF67, Ohk68, Ohk69, RH69, Tie67, VG68, Vai69]. lipoprotein [TB66]. lipoproteins [DL69a]. liver [ST68]. living [Pol65, RS66, SR66a, SR66b, Sim67]. Local [PH66, Per62b]. Localized [Kor67]. locomotion [Kav63]. locomotive [Wei64]. log [Koc66, Koc69]. log-normal [Koc66, Koc69]. logarithm [Koc66, Koc69]. logic [WPY67]. logical [Mar61, Sug63, SF63, Sug66]. logics [HR69c, HR69a, HR69b]. Logistic [Blu68]. London [CPC66]. long [Fit64, Hug68, KS67, Lea63, Lea65, Par67]. long-range [KS67, Par67]. Lotka [Gar67a]. low [Alm69]. lunar [Gil64]. lung [LG68]. lymphoblasts [MM68b]. lymphocytes [CS68, Cou69, Fan69, Fit64]. lysogenic [Noa68].

machine [RTB67]. machines [Cai61, Chi65, RTB67]. macromolecales [ZS66]. macromolecule [Man69a, ZS65]. macromolecules [Coh69, Dan69]. macrophages [Jan64]. made [Coh67]. Magnesium [WW68a]. Magnesium/calcium [WW68a]. magnitude [Fre62b]. Maintenance [BDF69]. makes [Mar64]. making [Gar67b]. Makower [Kat67]. mammalian [Cro68, DSB68, Low61, Tom66a]. man [Beg63, Wag68]. manufacture [Sha65]. mapping [Boe63, Wu64]. March [Ano62i, Ano63l, Ano64k, Ano65k, Ano67t, Ano68o, Ano69t]. markers [SZ65, Wu67]. Markov [DS67, Wen61]. mass [Sta66]. masses [PEC69, Ste68]. material [LM66]. Mathematical [Kel65, Lau65, Lau69a, Lau69b, Lin68a, Lin68b, Noo65, Wal66b, Wal66a, Els69, Est66, Gof66, KS69, PZO68, Sav69a, Sta66]. Mathematics [Gri68a, Gri68b, Har63]. mating [BCB64b, Ben63, BO65, Shi67a]. matter [Boy62, Per62b]. maximum [Car64]. May [Ano62g, Ano63n, Ano64m, Ano65m, Ano66q, Ano67r, Ano68k, Ano69m, Coh67]. Mean [Coe68, RKB67, Spr67, Wil69]. means [Kom63]. Measles [Bla66]. measure [GH67]. measured [Fre62b]. Measurement [JT69, Koi67, PH66, Pie66a, Pie67]. measurements [EDT66]. measuring [ADB61, LM66]. Mechanical [Sha65, Cro68, NIT^{+}65, Pol65, Vai65b]. mechanics [Arn61, Pat67]. Mechanism [Zár69, AAB69, Bel61, Ber62c, Ber64d, Cam67, Cen68, Cen69, Coh69, Con69, DSB68, Dan64, DS67, DiS69, Gin67, Gra64, Had67, KMT65, Kla66, KM65, Mor69, NM68, Nao68, PVP61, Per62a, PR68a, Sib64, Sim65a, Sol66a, Spe64, Sto65, Syk66, Tho68, TS66, Wal66b]. Mechanisms [Rus68a, Ain68, Bra65, Cen68, Cen69, Cze68a, Cze68b, Dun63, Res69, Tri64, WW68b, Koc66]. mechanistic [Els62]. mechano [Vai66b]. mechano-electric [Vai66b]. mechanoreceptors [Dun63]. media [Cin69, GB62a, VG68]. medullated [Pic66]. meiosis [Nic66, Ric65]. melanogaster [Car64]. melting [GFR68, GM69, MM66b, MM66a]. Membrane [Bar65, Gin67, JH69, Lin67, Lin68c, Sch68, TM68, Tom66a, CT67, DG63, Dic64, EH68, FBG^{+}66a, FBG66b, Ged68, GG67, Gin68, Hal61, HWJ69, HKG68, KK64, MR63, Ohk69, PGH63, PF67, RH69, Sem67, VPT64, Vai65b, Vai65c, VP66a, Vai66a, Vai66b, Wob68]. membranes [Bel62, CC65, Dan64, Dan66, EH69, ERH69b, Gro67, HHT65a, HHT65c, HH66, HK66, Jah62, Kim66, Luc64, MR68, O’B67, Ohk69, Seg67, Seg69, SS66, Tie67, Vai64, VP64, Wan63, Wat65]. memory [Bra68, Urb68, Wil64]. Menten [O’S62]. mesentery [GB62b]. messenger [Gar67c, Pol63, Sin69]. Metabolic [Kau69, Koc67b, BR66, Ber64a, Koc62, MHMQ64, MLH66]. metabolically [Coh69]. metabolism [Gar62, Gre66a, Gre66b, Kla66, Koc68a, Low69, Mar64, MP63, Nel63, Pol63, Rot63, Rus68b, Wil69, Wya64]. metal [Dan69]. Métastables [LL62]. method [BA63, Car64, Dan69, FCR67, Fri69a, Kin64, Pri67a, Tak66, Tak68, Wal67, KS68b]. méthode [WP63]. Methods [ADB61, Lau65, Per62b, Tie67, ZES68]. metric [Pet64]. micelles [Luc64]. Michaelis [O’S62]. micro [RS67]. micro-organisms [RS67]. microbial [Fuj63]. microscopy [Mal65]. microtabules [Sub68]. microtopography [Tol63]. mictic [Shi69a, Shi69b]. migration [KCE69, Mel64a, Mel64b, Mel64c, Mel64c]. milk [PF67]. minimum [Wil66]. misuse [Seg69]. mitochondrial [BDF69, SA66]. mitomycin [Mur66]. mitoses [TB68]. mitosing [Sag67]. mitosis [Bro69, Nic66, Ric65, Sub68, Tak66, Tak68]. Mitotic [Sim67, Eng68]. mixed [BCB64b, Shi67a, SZ65]. mixtures [BCB64a, Mel64a, Mel64b, Mel64c, Mor69]. Mobile [VP66a, Vid66]. mobility [GB62a]. mode [Bae64, BBS63, CT67, Mur66, VWM61]. Model [Ste68, Ber63, Ber64c, Bue66, CT67, Cro68, CD69, DiS69, DV66, DD66, Ess66, Est66, Est68, Ghe66, GC68, Gof66, GC69, GB66, GGW68, Has66a, Has66b, Hei64, Her69, HCS66, JH69, Kar67, Kie68, KS69, KS68a, Lau69a, Lau69b, Lei66, Lev66, LG68, Lew66, Mad69, Mar64, Mil66, Noa68, OF67, PZO68, Pel69, PCB68, Pen69, RB69, SS64, SKB69, SH64, Sim65c, Sin69, SS67a, Sta67, Sti68, Swa69, TP63, TS66, TS69, Tyl66, UE65a, UE65b, VPT64, WPY67, War67a, Wil67, Wob68, YSB65, Zaj68]. model-system [Hei64]. modeling [Wue69]. Models [TB66, DG63, ET67, EKH66, FBG66c, Haa69, KP67, Lei69a, Lin68a, Lin68b, Noo65, Phi69, Pri67a, Pri67b, Pri67c, RP68, RD69, Ros67, Rus68a, She62b, SU69, Vid66, War64, HK66]. modern [Mil66]. modification [Mil66]. modified [HKG68, O’S62]. modifiers [Bod63, Cen68, Cen69]. moleculaire [WP63]. Molecular [DL69a, ET67, SG63, VP66b, War64, dF62, NO66, Pul62, Rin65, Sug63, Tie67, WW68b, Wer65, Wol62b]. molecule [Erh69a, Fon67b, IKTN66, Mur66]. Molecules [ZP65, GFR68, Hel68a, Hel68b, Kli66, NIT^{+}65, Pla61, Sim65b, Wan63, LL62]. monkeys [Alt65]. monolayers [Ohk67]. monovalent [Bri66]. morphogenetic [Wad65]. morphologically [Ste68]. mosaic [War64]. most [RM68]. motion [RB69]. motor [KS68a]. moulding [Ham66]. movement [Bel61, BT62, Doe66, Wei62]. Mulluscan [WC69]. Multi [Sim65c, Hed66, PEC69]. multi-hit [Hed66]. multi-inlet [PEC69]. Multi-steady-state [Sim65c]. multibarrier [Sch68]. multicenter [ZS65]. multicompartmented [RS62]. multimolecular [Cin69]. multiple [Cap66b, Cap66c]. multiple-flow [Cap66b, Cap66c]. multiplication [TS66]. multistate [BW69]. multivariate [Haa69]. muscarinic [SW67]. muscle [Ano66f, Bar63, Cap66a, Cro68, Koi66, Koi67, Loe68, Mat68, Tom66a, Ull67, Vay65]. muscular [Arn61]. mutagenic [Lad64, NM68]. mutant [RS67]. Mutants [Her62]. mutation [Eng62, RTB67, Sim65a, Str62]. mutational [Car64]. mutations [KM65]. myoglobins [Ata66].

nachbau [Kli66]. narrow [Hug68, Lea63, Lea65]. native [Dav66, Lew67]. natriuresis [Wal66b]. Natural [Bod63, Eml66, Spe64, Bla67a, Ham66, Has66b, Mil62, War67a]. nature [CS66, Cro63, Fre63, Koc67c, Nor68]. navigation [Brl64, KCE69]. necessary [Coh69]. negative [Boe63, HR66, Gri68a]. neighbor [SZ62]. nephron [Kel65]. nerve [Dan64, KS68a, Par66, Par68, VP66b, Wob68]. nervous [Pic66, Psc63]. Net [EKH66]. nets [Kau69, Ros67]. network [MHMQ64]. networks [Sti68]. neural [Ros67, Wer65]. Neuroelectric [Lew66]. neuron [Enr67, dLR67]. neuronal [NN66]. neuronic [Lau65, Lau69a]. Neurospora [Ber64d]. Newton [Hen68]. Niche [Wue69]. ninefold [SS64]. Nitella [HWJ69]. nitrogen [DP69, Pra62]. nitrogen-fixing [DP69]. NMR [Dan69]. Non [BLT^{+}62b, BLT62a, Ber65, Cha66, CGTK67b, Els62, FBG^{+}66a, Haa69, JT69, Kos62, MM68a, Pas67, RM68, Tru67, VP64, Eck62]. non-circular [JT69]. Non-classical [BLT^{+}62b, BLT62a]. non-electrolytes [Pas67, VP64]. non-enzymic [Kos62]. non-equilibrium [RM68]. non-equivivant [Tru67]. non-hibernating [Cha66]. non-linear [Haa69, MM68a]. non-mechanistic [Els62]. non-permeable [CGTK67b]. Non-random [Eck62]. non-stationary [Ber65]. non-statistical [FBG^{+}66a]. nonelectrolytes [Bue66, CC65]. Nonlinear [Dew64]. nonreabsorbed [Kel65]. noradrenaline [KMT65, KT65]. normal [DiS69, Koc66, Koc69, MM68b]. Normalization [Mat68]. North [Ano63b]. Note [Els63, LP66, Nic66, Wu64]. November [Ano62h, Ano63o, Ano64n, Ano65l, Ano66s, Ano67s, Ano68l, Ano69n]. nuclear [Mal65]. nucleic [Cam67, Dou62, RFCH68]. nucleophilicity [EMM68]. nucleotide [De 69, Gra64]. nucleotides [Kla66]. number [Car64, DPW66a, HM68, Mag69b, SP69]. numeric [RTB67]. Numerical [FT68, BBKL66]. numero [SB65].

objects [Koc68b]. obtained [Shi67c]. occupation [Hom67]. occurrence [Eck62, PP68]. October [Ano67o, Ano68g, Ano69u]. Octopus [Mal63]. ocular [Beg63]. odor [Sin67]. odours [Dav65]. oil [CC65, GG67]. olfactory [BW63, Bos68]. oncogenesis [Gil68]. one [Cen68, Cen69, DPW66b, DS67, Gri68a, Gri68b, HH69, Lin68a]. one-sided [Lin68a]. one-substrate [Cen68, Cen69]. Onsager [SS66]. Ontogeny [RG69]. Open [Han67, KCE69]. open-sea [KCE69]. Operational [GM68]. operon [Geo69, Koc67c, SR66b]. operons [Sim67]. Optical [CC69, Tie67]. optima [Sch62]. Optimal [Rub67, SAFT66]. Optimizing [Coh66, Coh67]. orbitals [Pul62]. order [BR64, Dia69, Lan63]. orders [Est68, SZ62]. ordinates [Sch68]. organ [AC63]. organic [Cin69, Des65, Why67]. organismic [Els61, Els63, Els64]. organisms [DD66, Her69, RS67, Spr67]. organization [Atl68, Geo69, GC69, Lau69b, Str69, Tie67, Wol62b]. orientation [HHT65b, Nic66, Ric65]. oriented [Ghe66]. origin [CS66, Cin69, Pat67, RR67, Sag67]. origins [Eng68]. oscillations [DH69, Dew64]. oscillator [Mar67]. oscillators [Pav69, Win67]. oscillatory [SS67a]. osmolality [Alm69]. Osmotic [JW67, EH69, Wal66b]. other [Atl68, Whi65]. outcome [Coh67]. Outline [Cai61]. overs [Shi64, Shi66c]. Oxford [Pen69]. oxidase [PVP61]. oxidations [Low61]. oxidative [Gra64, Rus68a]. oxygen [Gol69, KS69, LM65, Sne65].

packed [Ell68]. Packing [Fon67b]. Pages [Ano61c, Ano61b, Ano61d, Ano62e, Ano62d, Ano62i, Ano62g, Ano62f, Ano62h, Ano63j, Ano63k, Ano63l, Ano63n, Ano63o, Ano63m, Ano64i, Ano64j, Ano64k, Ano64m, Ano64n, Ano64l, Ano65i, Ano65h, Ano65k, Ano65m, Ano65l, Ano65j, Ano66x, Ano66w, Ano66v, Ano66r, Ano66t, Ano66p, Ano66u, Ano66q, Ano66s, Ano66o, Ano67n, Ano67q, Ano67v, Ano67p, Ano67l, Ano67u, Ano67m, Ano67t, Ano67r, Ano67s, Ano67o, Ano67w, Ano68h, Ano68j, Ano68q, Ano68m, Ano68i, Ano68n, Ano68f, Ano68o, Ano68k, Ano68l, Ano68g, Ano68p, Ano69j, Ano69l, Ano69q, Ano69o, Ano69k, Ano69r, Ano69s, Ano69t, Ano69m, Ano69n, Ano69u, Ano69p]. Pair [TP67, BS69, KS67, Lan63]. pair-order [Lan63]. pairing [GFR68]. paper [Goe67a, Hug68, IJ68, ZZ69]. Parallel [She62b, Vid66]. parameter [SP66]. Parameters [PL67, HWJ69, Vay65]. parasite [TL66]. parasites [BNW62]. Paratene [Ehr67]. parathyroid [Rig66]. Part [Lev64, Kel65, Arb67]. parthenogenesis [Tom66b]. partial [Est68]. partially [PEC69]. particle [Bar62, LMB68]. particular [BR64, DG63]. particularly [BLT^{+}62b]. Partition [Orl68, CC65, SW67]. Passive [MHMQ64, Bri66]. patch [DSB68]. pathway [Cou69, Lei69a, Sol66b]. pathways [D’A63, Han67]. pattern [Pie66b, Sat67, WC69, Wol69]. pattern-diversity [Pie66b]. patterning [Bos68]. patterns [Cla64, Lan63, RTB67, Rot63]. pedigree [Shi66d, Shi69b]. penetration [Bue66, Bue67, CC65, GG67]. Peptide [IKTN66]. peptides [Sne66]. Percentage [TB68]. perception [Wil61b]. periodically [SB65]. periodicities [Gil68, Wad65]. peritoneum [GB62b]. permeability [Dic64, Dun65, Ged68, HH66, RH69, Sch68, Vai64, VP64, Wat65]. permeable [CGTK67b, JW67]. permease [Koc67a]. Permeation [Hal61, Lea63, Lea65, Hug68]. Permissive [Cam67]. persistence [Sta66]. Petal [Sat67]. pH [Sch62]. phage [Boe63, Pol63]. pharmacogenetic [Nel63]. pharmacologic [Wag68]. Pharmacological [Wat65, DS66]. phase [Dar68, GC69, HH69, Koc64, VP66a, Wev66]. phase-shift [GC69]. phases [EH68]. Phenogram [HM68]. phenomena [FBG66c, Hol67, Wat65]. Phenomenological [Fon68, Lin68c]. phenotype [HM68, Ren62]. philosophy [Fei67]. phosphate [BDF69, D’A63]. phosphates [BL69]. phospholipids [DL69b, Tri69]. phosphorylation [Gra64, Rus68a]. photochemistry [Cla67]. photodimerization [NIT^{+}65]. Photoreceptor [Wol62b]. photosynthesis [Bas63, Cal61, Cla63, Cla67, FR64, Ids68, Tol62]. photosynthetic [Kno68]. phyletic [Psc63]. phylogeny [Est68]. Physical [Els62, Mor66, Pol61, Gin67, Kor67, Tab67]. physical-thermodynamic [Tab67]. physico [Ros68]. physico-chemical [Ros68]. physics [Ste64]. physiology [TR65]. phytosociology [Orl68]. pigmentation [WC69]. pipiens [Mat68]. pitches [Jar64]. placental [KS69]. planar [FBG66b]. plant [FR64, Loc65]. plants [Arn68, Bro62, Ids68, Lev62]. plaques [Koc64]. plasma [HR66, PF67]. plastics [Kor67]. plateau [KT69]. plausible [Kar67, Nao68]. point [Hol66a, Wal67]. Poisson [MM68a, Mac69]. Polar [HHT65b]. polarized [DKGK66]. polyelectrolyte [Ell68]. polyene [PLB69]. polymer [McM63]. polymerase [Sib64]. polymorphism [FBM66]. polynomial [CFY68]. polynucleotide [McM63]. polynucleotides [HL64, SZ62, SZ65]. polypeptides [Bla67b, Chi65]. Polyporus [PP68]. Polyribosome [Ord69]. Polyribosomes [Sin69, GL65]. Polysaccharide [Bel62]. polysomic [Fis62c]. polyvalent [Bri66]. pool [Ber64a, Sav69b]. pool-sizes [Ber64a]. pools [Koc62]. Population [Hol67, Gar67b, Has66a, Mil62, PCB68, Pen69, RS67, Shi63]. Populations [Pav69, BCB64a, Bla66, Bla67a, CB61, Coe68, Gar67a, Han69, Kie68, LR69, Noa68, Noo68, PCB68, Sai68b, Sai68a, Shi64, Shi66a, Shi67a, Shi67b, Shi67d, Sta66, TL66, Tru67, TB68, Win67]. pores [Hug68, Lea63, Lea65]. position [Wil64]. Positional [Wol69]. Positive [Gri68b]. possibilities [Pul62]. possibility [Gil64]. Possible [Lad64, Wil61a, Wil62, AAB69, Bae64, Cal68, Cin69, GK69, Gin67, Hed66, Kau67, Kin64, Kor67, Mar64, Par66, Par68, RB68a, Sol66b, Tho68, Tyl66, WW68b]. postulate [Ber63]. postulating [Cen68]. potassium [RPS61]. potential [Bar65, Con69, FBG66c, Gin67, Gin68, Kim66, Kom63, Pul62]. potentialities [O’S62]. potentials [BT62, Jah62, Koi67, Lew66, MR68, VP66b]. power [DPW66b, Mar64, Sav69b]. power-law [Sav69b]. practical [Kat67]. practicality [Wal67]. prebiological [Hul69]. precipitates [vO63]. precursor [BR64]. precursors [PS68]. Prediction [Boc67, Dar69, De 69, Orl68]. Predictive [Acl67]. Preface [Dan61]. Preliminary [Atl68]. preparation [Kat69]. presence [Kel65]. present [Kat69]. Pressure [Doe67, Cor65, Wei64]. pressures [CC65]. prevalence [Bur66]. previous [Cze68c]. primary [Cla64, Tol62, WW68b]. primitive [CS66]. principle [Seg69, Acl66]. Principles [dD64, Mel64a, Mel64c, vdS69]. Prinzip [Hel64]. pristine [Gil64]. Probabilistic [dLR67]. probabilities [Shi67b]. probability [Shi66a]. probable [RM68]. Problem [ERH69b, Blu62, Jeh67, KCE69, Mar67, RB63, Rub67, Sat67, TR65]. Problems [Boy66, BKK66, FT68, Lei69a, Sta65]. procedures [Pri67b]. process [BR66, Bla67b, SB65, SAFT66, Tol62, Wen61]. processes [BCB64b, Cai61, DG63, DS67, Gri68a, Gri68b, Koc62, Koc68a, Res69, Sam66, ST68, SR66a, SR66b, Sim67, Van67, Wal66a, Wer65]. processing [Lei66, Lei69a, RTB67]. producing [BDF69, Kin64]. product [Cen69, DS67, HH69]. production [Gro69, Hel68a, Hel68b, Wil66].

Professor [Fis62d, Ros64]. profiles [Bar65, RB68a]. program [Gat63, Sti68]. programming [Pri67c]. proliferating [Bro69]. proliferation [Con69, Rot63]. Propagation [Bar63, Doe69, Pic66]. Properties [Ohk69, Pla61, Ber61b, HHT65b, HT63, Sav69a]. Proposed [Wag68]. protein [Big67, CFY68, Day65, Eck62, Gar67c, GD63, GL65, Jar64, KS67, KS68b, Lew69, MP63, Pol61, QZ62, RB66, Rob65, SR66a, Sin65, Sin69, TP67, War64]. proteins [Ata66, Big67, HHT65a, Hav66, Kar67, KS68b, Par69, Smi66, Whi65, Yca61, ZES68]. Proteinschrauben [Jar64]. Proton [BS69, Rus68b]. protoplasts [Mar63]. proximal [FW65, Kel65, Wu67]. Psycho [HR69c, HR69a, HR69b]. Psycho-logics [HR69c, HR69a, HR69b]. pulse [Par66, Par68]. pump [Fru65, RPS61, UE65a, UE65b]. pumps [Rob68]. pure [BCB64a]. purines [CPC66, PR66]. puriques [WP63]. purse [Pel69]. pyrimidines [CPC66, PR66]. Pyrimidiques [WP63].

quality [Dav65]. Quanta [Els61]. Quantitative [O’S62, AP65, Jar64, PA63, Rig66]. quantitativen [Jar64]. Quantum [Cal61, NIT^{+}65, Pat67, Ste64, CM62]. Quantum-mechanical [NIT^{+}65]. Quasi [Wal66c]. Quasi-steady [Wal66c]. quaternary [Dav66, Fel68, SW66, SW67, Whi65].

R [Ano66f]. racemic [Mor69]. radiation [Atl68, DSB68, Eng62, Lad64, SP66, SP69]. radiation-induced [Eng62, SP69]. radioactive [Bro69]. Radiosensitivity [Ter65]. rana [Mat68]. Random [Ben63, BCB64b, BO65, Eck62, She62b, SU69, Shi67a]. randomly [Coh66, Coh67, Kau69]. range [JT69, KS67, Par67]. Rapidly [Sin65]. rat [Gar62]. Rate [Sem67, Blu68, Dar69, DH69, FCR67, Gar67b, HE65b, JD69, Kat69, Løv63, Mag69b, NS61, O’S62, Par66, Par68, RP68, Rot63, Sav69a, Sch68]. rates [Bur66, CC65, Koc62, Koc68a, LM66, PEC69]. ratio [GGW68, RR66]. ratios [AC63, Ess66, Ged68, Hed69]. re [Wev66]. re-entrainment [Wev66]. Reabsorption [Wal66a, Kel65, Wal66b, WGP66]. reacting [Sie69]. reaction [Ain68, AP65, HH69, Kat67, Kat69, Kos62, PA63, Tri64, Wan63]. reactions [AP65, Bri69, Cze64a, Cze64b, Cze67, Cze68a, Cze68b, Cze68c, DPW66a, DPW66b, Dar68, Dia69, FCR67, Fis62a, Hig68, KT65, KSK65, Mag69b, Mal65, PA63, Res69, Sav69a, She67]. read [Nao68]. read-out [Nao68]. reality [Noo65]. Rearrangement [Kat67]. recentor [KMT65]. Receptive [Hol66a]. receptor [CS68, DS66, Fan69, Fel68, Hom67, Kar67, KT65, Lin63, SW67]. receptor-hypothesis [Lin63]. receptor-storage [KT65]. receptors [HCS66, Wat65]. reciprocal [SS66]. Recognition [Cou69, Sim65b, Man69a]. Recombination [Wu67, Ber62c, Ber64d, Fis62b, Shi67b]. reconstituted [MR63]. reconstruction [Kli66]. recovery [DS66]. recurrence [RKB67]. red [Ber65, WZ69, Zaj68, Zaj68]. redox [Jah62, Pul62]. Reductions [Low61]. reference [Cla64, DG63, Lei66, Ser64]. reflex [Beg63]. reflexive [Koc67b]. refractive [Ohk68]. regarding [Aid66]. regenerative [ST68]. regimens [Ano69g, KT66, KT69, Nel62]. region [Ber62c]. regional [EDT66, PEC69]. regression [CFY68, Haa69]. regular [HM68]. regulated [Ano66f, Cap66a]. Regulating [Rob68, HR66]. Regulation [Kla66, RB66, RS66, SR66a, SR66b, Sim67, CT67, CHS67, Geo69, Lev66, Pai62, Syk66, TM68, YSB65]. regulatory [Noa68, TS66]. related [FBG66c, Gro69, JI69, PLB69]. relating [CC65]. relation [Ano66f, Ata66, Big67, BT62, Cap66a, Cap66b, Cap66c, Gin67, Hed66, Nob69, Rot63, Sha61, Wei64]. Relations [CRJ66, Sne66, Cla67, Kat69, Rap65, SS66]. relationship [DS66, GD63, Goe67a, Gre66b, Ren62, Spe67]. relationships [DS68, Gat64, Wag68]. relative [AC63, CB61, Wal65]. relaxation [Cze64a, Cze64b, Cze67, Cze68a, Cze68b, Cze68c, KSK65]. release [Gro69, KS68a]. remark [TR63]. remarks [Hig68, HL64]. renal [Des65, Lev66, Wal66b, Wal66a, WGP66]. renal-electrolyte [Lev66]. renewal [SB65]. repair [DSB68]. Replacement [Yca61, QZ62]. Replication [Kal68, Dou62, DV66, Erh69a, Goe67b, Kau67, Phi69, Sto65, ZS66]. Reply [Fis62d]. Representation [Shi63, Hil66]. repression [PL67]. repressor [Koc67c]. repressor-operon [Koc67c]. reproducing [RTB67]. reproduction [Coh66, Coh67, DD66, Sta67]. repulsion [Par67]. requirements [Gre66a]. requires [Hel68a, Hel68b]. research [Kom63]. resistance [Boy66, Koi66, Koi67, Lev62, Tom66a]. resolution [Mal65]. resonant [Res69]. resorption [Sha65, WO67]. resources [AB61]. respect [PLB69]. respiration [DH69, SA66]. response [DS66, GK69, Lev64, TP63, Wag68]. responses [Gin67, Hom67]. Resting [MR68]. resulting [Kor67]. results [Cze68c, Noo65]. retinal [Nas69]. retinals [PLB69]. Reversible [Urb68, Bra68, Cen69, DS67, Pul62]. review [CM62, Fre63, Mar63]. revised [Gra64]. rhesus [Alt65]. rhythmicity [Enr65]. rhythms [PZO68, Pav69, Swa69, Wev66, Win67]. ribonuclease [Lan63, Sar65, Wei68]. ribonucleic [Boc67, RS66, Syk66]. Ribosomal [PS68, Rob65, Syk66]. rich [BL69]. ring [PLB69, Shi63, Shi66c]. RNA [Bra68, Gar67c, Geo69, Koc68a, Les69, PS68, Pol63, QZ62, RB63, Sin65, Sin69, Urb68]. RNA-protein [QZ62]. Rod [Shi66c]. Role [Sub68, DL69b, Eyl66, Fit64, Jan64, MM66b, PF67, RB68a, Sam66, Tri69]. root [TI67]. rotating [CS65]. rust [Boy66].

S [PLB69]. S-cis-S-trans [PLB69]. S. [Ano66f]. salicylic [BLT62a]. salt [Cin69, MM66b]. sartorius [Mat68]. scattered [Koc68b]. scattering [ADB61, LMB68]. SCF [IKTN66]. scheme [PR66]. sciences [Pet64]. Scrapie [HKG68]. screws [Jar64]. sea [KCE69]. search [Enr65]. secretion [DiS69, PF67]. Sedimentation [Mel64c]. Segmental [BR66]. seifing [Shi67c]. seining [Pel69]. Selected [Cze67]. Selection [dF62, Arn68, Bla67a, Bod63, Cha66, Eml66, Ham66, Has66b, Lev64, Lur69, Spe64, War67a, RTB67]. selections [Shi67b]. selective [Heb68]. selectivity [Lin67, Tri64]. Self [Fis62d, Ano66f, Cap66a, Deu63, Lau69a, Lew62, RTB67, Sta67, WP63]. self-consistant [WP63]. self-consistency [Lau69a]. self-incompatibility [Lew62]. self-regulated [Ano66f, Cap66a]. self-reproducing [RTB67]. self-reproduction [Sta67]. Self-Sterility [Fis62d]. self-stimulation [Deu63]. selfed [Shi63, Shi64, Shi67b, Shi67d]. selfing [BCB64b, Shi67a]. semantics [Mae65b]. semi [GFR68, IKTN66]. semi-empirical [GFR68, IKTN66]. senescence [Ham66]. sensitized [CS68, Fan69]. separating [EH68]. separation [PR68a]. September [Ano62f, Ano63m, Ano64l, Ano65j, Ano66o, Ano67w, Ano68p, Ano69p]. sequelae [RG69]. Sequence [SZ62, Day65, GD63, Lan63, RB63]. sequences [KS67, KS68b, NN66, TP67, ZES68]. Sequential [Chi65, Kau67]. Serial [SU69]. series [DPW66b, Enr65]. serine [EMM68]. serrate [Daw66]. serum [DL69a]. ses [VP65a]. set [Shi66a]. sex [Ben63, Fis62b, Hor69, RR66, Sin67]. sexuality [DeL67]. shapes [Ehr67]. Shear [Hig68]. sheep [Cla64, WOB62a, WOB62b]. shift [GC69]. shifts [Wev66]. short [GH67, Reh68]. short-circuit [GH67]. short-circuiting [Reh68]. shortening [Mat68]. shown [Vid66]. shrinking [BDF69]. Shunt [D’A63]. sided [Lin68a, Lin68b]. signal [Par66, Par68]. signals [Par66, Par68]. Significance [Ata66, Cin69, GK69, Len62]. similar [AC63, Koc68b]. similar-sized [Koc68b]. similarities [GdlB66, GM68]. Similarity [Ebe69]. simple [DD66, Gar67b, Mar64, RD69, Lin68b]. simplified [Has66b]. simulating [Koc69, PCB68, Zaj68]. Simulation [Fra62, FBM66, RTB67, RD69, SKB69, SA66, Sug66, Gar62, Gar67b, KP67, Pri67a, Pri67b, Pri67c, RP68, ST68, WC69, War67b]. simultaneous [EDT66, Nao68]. single [Cen68, CF69, Hed69, Hed66, JH69]. single-hit [Hed66]. single-stranded [CF69]. sinusoidal [Wil69]. sister [Hed69, Kau67]. site [CS68, Fan69, Fel68, O’S62, SP66, SW66, SW67, Win63]. sites [SP69]. size [Bla66, YSB65]. sized [Koc68b]. sizes [Ber64a, Bro64]. skeletal [Mar64]. skin [Cla64]. slices [Gol69]. sliding [RB69]. slowly [O’S62]. small [Fit64]. smooth [Bar63, Koi66, Koi67, Tom66a, Vay65]. social [Alt65, Ham64a, Ham64b, Lei69b]. Sociobiology [Alt65]. sodium [Kel65, RPS61, Sem67]. sodium-activated [Sem67]. soil [Ids68]. soil-induced [Ids68]. Sokal [Est68]. soluble [Alm69, BBS63]. solute [JW67, Kel65, PGH63, Wal66b]. solutes [GB62b, Wal66a]. Solution [Cal68, WA65, BBKL66, De 67, Est68, Mac69]. solutions [Cin69, DPW66b, HT63, Sav69b, Sne65]. solvation [Lin67]. solvent [PGH63]. solving [Rub67]. Some [Agi63, Agi64, Alm69, CM62, Con69, DKGK66, Doe66, FW65, Hig68, HL64, Kom63, Lei69a, Lew67, Per62a, Ros68, Str62, Vai66a, Ain68, BNW62, Ber64b, Boe63, Gre66a, HHT65a, HOD63, Heb68, Sem67, Sha61, Sav69a, Wei68]. source [Par66, Par68]. spacing [Wil66]. span [Fit64, Kom63, Lan63, MM68b]. spatial [GC69, Wol69]. special [Cla64].

Species [Pie66b, Pel69, PP68]. Species-diversity [Pie66b]. specific [Bur66, Kin64, Mad69, Win63, KS67]. specifically [vO63]. specification [Wue69]. specificity [Des65, Kil66a, Kno65, Tri64, Wer65]. spectra [Bro64, JI69, KSK65]. spectral [Tol62]. spectrometric [Dan69]. spectrum [ADB61]. speculation [RR67]. speculations [Doe66]. Spetner [Has66b]. sphere [Mac69]. spherical [De 67, Hal61]. spider [PP68]. spike [Bra65]. spindle [Cro68, HCS66, Sub68]. Spontaneous [Bur66, Enr67]. spread [Lau69b]. spreading [GB62a]. Squalene [Sol66b]. Stability [VG68, Wal69b, Ell68, Gar67a, HT63, Kau69, Lew67, Mel64b, MHMQ64, Wal69a, WOB62b]. stabilizing [GB62a]. Stable [Mel64a, Mel64b, Mel64c]. Stable-flow [Mel64a, Mel64b, Mel64c]. stacked [CPC66]. stacking [GFR68, NO66]. STAFLO [Mel64c]. State [Ano63b, Cen68, Cen69, Coe68, CD69, Dar69, Gar62, GB66, GGW68, HK66, Hil66, Lew67, Noa68, Ord69, Pas67, PEC69, Sav69b, Sim65c, Vai66a, Wal66c, Wal67]. states [BS69, WW68a]. stationary [Ber65]. Statistical [DD66, NN66, TP67, Vai65b, Vai65c, FBG^{+}66a, GM69, She62a, ZES68]. Steady [Cen68, Cen69, Pas67, Coe68, Dar69, Gar62, GB66, GGW68, HK66, Hil66, Noa68, Ord69, PEC69, Sav69b, Sim65c, Vai66a, WW68a, Wal66c, Wal67]. Steady-state [Cen68, Cen69, Pas67, Coe68, Dar69, Gar62, GGW68, Noa68, PEC69, Sav69b, Wal67]. stem [VNM69]. step [Cze64b, Cze67, Cze68a, Cze68b]. stepwise [CFY68]. stereoisomers [Nas69]. Sterility [Fis62d]. steroid [NO66]. stimulating [Atl68]. stimulation [Deu63, Enr67]. stimulatory [SA66]. stimulus [DS66, Rap65]. stimulus-response-recovery [DS66]. Stochastic [BCB64b, She62b, SU69, VNM69, Ber62a, Ber62b, Bla67b, HH69, JM66, Kor69, LG68, Pel69, Sin69, TL66]. stochastically [Noo68]. Stochastics [Alt65]. stone [FM69]. storage [KT65]. strand [PR68a]. stranded [CF69]. strands [Nao68]. strength [MM66a]. stress [Cur68, Pol65]. Stresses [Doe68, Ids68]. striated [Ull67]. strict [Gar67a]. Structural [Jeh67, Geo69, Kil66a, KS68b, Par67]. Structure [O’B67, WP63, Ata66, Big67, BS69, CRJ66, Dav66, DL69a, Fon67a, HR69a, HR69b, Kil66b, MR63, Sar65, Sne66, Tyl66, VP65a, VP65b, Whi65, Win63]. structures [BL69, JI69, Wol62b]. studied [TS69]. Studies [HK66, Hil66, Wei68, JM66, KS68b, MM68b, Nel63, Pol65, SB65, Tak68, TP67, WOB62a, WOB62b]. Study [RS67, Ain68, Atl68, Gar67b, GC68, Haa69, Hav66, Lau65, Lau69a, Lau69b, MM66b, MLH66, NIT^{+}65, Nas69, Pie66b, RFCH68, TL66, YS67, YP64]. sub [Pla61]. sub-groups [Pla61]. subclinical [RKB67]. subject [Enr67]. subjected [Wil69]. submolecular [SG61]. subsequent [Coh67]. substance [CGTK67b]. substitutions [KS68b]. substrate [Ain68, Cen68, Cen69, DS67, Des65, Dia69, Fis62a, Fuj63, Haa69, HH69, Kat67, O’S62]. substrate-one [DS67]. subunit [RB68b, War64]. subunits [GM69]. succession [Pie66b]. sucrose [Arn68, Koi67]. sucrose-gap [Koi67]. sugars [LL69]. suggested [Boe63, War64]. suggestion [Aid66]. sulfhydryl [Lev62]. sulfhydryl-disulfide [Lev62]. sulfonamides [Nel62]. sulphonamides [HOD63]. super [Cha66]. super-hibernating [Cha66]. suppressor [Boc67]. supra [SG61]. supra- [SG61]. Surface [Bel61, FBG66c, Gin67, Gin68, GGW68, Løv63, Sha65]. surface-volume [GGW68]. surfaces [FBG66b, Wei62]. Survey [Ber64b]. survival [Ber62a, Ber62b, Ber65, Hol67]. suspensions [ADB61]. swelling [BDF69]. swine [Sin67]. switch [Cla63]. symmetry [SS64]. sympathomimetic [Sha61]. synaptic [Bra65]. Synopsis [Els64]. synthesis [Alm69, Bae64, Cam67, CS65, D’A63, Geo69, GL65, Hei64, Hel68a, Hel68b, Hul69, Man69a, Man69b, Pol61, Pol63, Rob65, RS66, Sib64, SR66a, SR66b, Sim67, Sin65, Sin69, Syk66, TB66, YP64]. synthetic [GFR68]. Syrian [Cha66]. system [Ber61b, CS65, CHS67, DP69, GB66, GGW68, Hei64, HR69c, IKTN66, Mae65b, Mal63, PGH63, Psc63, RB68a, Sav69b, SB65, SS67a, Sug66, Wal66c]. systems [BR66, Ber64a, Bri66, Cal68, Cap66b, Cap66c, CFY68, Cze64a, Cze64b, Cze67, DKGK66, Ell68, Fra62, FBM66, FBG^{+}66a, FBG66b, Gar67a, Gar67b, GH67, GC69, HM68, Hig68, Hil66, Hol68, JM66, KP67, LP66, Lin68c, MH69, Mag69a, Mar64, Mor66, Myh69, PEC69, Pol65, Pri67a, Pri67b, Pri67c, Pul62, Rap68, RS62, RP68, RM68, Sav69a, Sav69b, Sem67, She67, SS67b, Ste64, Sug63, SF63, Sug66, TR63, WW68a, WPY67, Wal69a, Wal69b, Wil69].

T [Zár69, Ber62c]. T-even [Zár69]. T4B [Boe63]. table [Wil69]. taking [WO67]. taxonomic [FT68]. taxonomy [Rub67, Sok62]. technique [Reh68]. temperature [CHS67, MM66a, PZO68, RH69, Wil69]. temperatures [Wil69]. template [McM63, ZS65, ZS66]. templates [Wad65]. Temporal [Bos68, GC69, Wad65]. tension [GG67]. tentative [Ber61a]. termination [Sie69]. terms [Fis62a, Pri67a, Pri67b, Pri67c, RP68]. territorial [Gar67a]. tests [Fre62b]. Their [Sem67, Wol62b, Ata66, Cin69, CC65, GM69, Ids68, Pla61]. theorem [De 67, PEC69, She67]. theorems [dLR67]. Theoret [Ano66f, Ano69g]. Theoretical [AKT62, AC63, Ano65n, Man69b, Pol65, Sie69, Sin67, Tak66, Tak68, YS67, Con69, De 69, DKGK66, GC68, Han69, MM66b, Nas69, VWM61, Wev66]. theorie [Jar64]. theories [Cal68]. Théorique [BP62]. Theory [DH69, Koc68b, KSK65, Lev64, LM66, MM68a, Mar64, RS67, Sch62, VPT64, Vai66b, AP65, Ano69g, AW65, Arb67, Atl68, Bue67, Cai61, DS67, Dav65, Ehr67, Els62, Els63, Els64, Est66, Eva62, Fon68, For69, FR64, GdlB66, GM68, Hom67, Jah62, Jar64, Kav63, Kno68, KT66, KT69, LP66, Lew61, Loe68, Lur69, Mae65b, Mar67, MP63, Mil62, Noo66, Ohk67, PA63, Sch68, Ser64, Smi67, SW67, Sta66, Ull67, Vai65b, Vai65c, VP66b, War67a, Whi65, Wil64]. therapeutic [Nel62]. thermal [Atl68]. Thermodynamic [ZZ67, IJ68, Tab67, ZZ69]. thermodynamics [Acl67, Ber64b, DG63, HK66, Hil66]. Thermoregulation [Hen68].

Thickness [Tie67, CC69, Dan66, Ohk69]. Thiemer [Ano69g]. thin [CC69]. thinking [Cai61]. thiopentone [NS61]. thorough [Pet64]. thought [Cai61]. thought-processes [Cai61]. three [Cze67, Cze68a, Myh69]. three-step [Cze67, Cze68a]. threshold [Fri69a, Hol66a]. thresholds [Gau64]. throttle [Low69]. thymine [Les69]. thyroid [DS68, DiS69]. tight [Koc67c]. time [Coh67, Enr65, GM68, Has66a, HE65b, RKB67]. time-series [Enr65]. times [Fri69a, Kal68, Man69a]. Tissue [LM65, Bro63, Gol69, Per62b, PH66, dD64]. tissues [Per62b, Reh68]. tit [Pen69]. TMV [Her62]. tobacco [War64]. tolerance [Fre62b, Len62, Sie69]. tonic [Enr67]. Topaz [Boc67]. topics [Gro69]. tortuosity [TM68]. total [LMB68]. toxicity [HOD63]. Tracer [Ber61b, Ber61a, Ber64a, EKH66, Koc62, Koc68a, Myh69, She62b, SU69]. tracing [Mal65]. training [Ano63b]. trans [PLB69]. transcription [Pai62, Phi69]. transducers [Vai66b]. Transfer [SS67b, SS67a, Acl66, Acl67, Acl69, BS69, Had67, Har63, KS69, Pul62, Res69, RB63, TR63, UE65a, Wol62b]. transform [BBKL66]. Transformation [BP62, Shi66d, Shi69a, Shi69b, Acl69]. Transient [Dar68]. transistor [JH69]. transition [CD69, Lin68c]. transitions [MM68a]. translation [Cam67, Gar67c, Mar67]. translocate [Arn68]. translucens [HWJ69]. transmembrane [Jah62]. transmission [Spe64, Wer65]. transmitters [Gro69]. transpiration [Ids68]. transplantable [TI67]. transport [Bre67, Bri69, CD69, Des65, GB62b, Gro67, HK66, Hol68, Jah62, KK64, Koc67a, LL69, Mil66, Rap68, RW63, Sem67, SKB69, Sne65, UE65a, UE65b, VPT64, Vai65a, Vai65c, Vai66a, Vai69, Vid66, Wan63, YS67]. trapping [Dan64, Kno68]. Treatment [IKTN66, Cen69, Cze68a, Cze68b, Has66b]. treption [Ser64]. Tribolium [Daw66, Sta66]. triggers [SR66b]. Triplet [Gat63, TP67]. tuberculin [CS68, Fan69]. tubular [WGP66]. tubule [FW65]. tumor [Kor67]. tumors [Rot63]. tumour [Gre66a, Gre66b, HOD63]. tumour-host [Gre66b]. tumours [TI67]. tunnels [Sta66]. turning [Sin65]. turnover [Ber64a, LM66]. turtle [KCE69]. twin [Hed69]. twins [Hed69]. Two [Hol66b, Ros67, Ain68, Cen68, Cen69, EH68, Hol66a, Kat67, Lin68b, PGH63, Pel69, PCB68, Wol62a, Cze64b]. two-break [Wol62a]. Two-factor [Ros67]. two-membrane [PGH63]. two-point [Hol66a]. two-sided [Lin68b]. two-species [Pel69]. Two-step-reactions [Cze64b]. two-substrate [Ain68, Kat67]. type [RTB67]. types [Pie66a, Pie67, Wal69a]. Typology [Sok62].

ultraviolet [Lad64]. unbranched [Han67]. uncatalysed [LP66]. unified [GdlB66]. unifying [Eva62, Ser64]. unimolecular [Hil66]. unimpaired [HR69a]. unit [Kno68, Wei63]. unmedullated [Pic66]. unoriented [DKGK66]. unsolvable [Sta65]. unstable [Sin69]. Unwinding [Phi69, Erh69a, ZS66]. uphill [LL69, RW63]. upon [Shi67b]. Uptake [AAB69]. uracil [Les69]. uric [Mil66]. usage [Pet64]. Use [Haa69, AC63, DPW66b, LM66, Wal67]. using [BBKL66, Fis62b, Mal65, Sav69b, Sug63, SF63, Sug66]. Uterine [Cor65]. utilization [FR64].

v [Ano62h, AP65, BLT^{+}62b, Fis62a, Wei68]. v-viii [Ano62h]. value [PK68]. values [Fis62b]. Van’t [Nob69]. vapor [CC65]. variability [Eng68]. Variable [Lan63]. Variable-span [Lan63]. Variance [Shi67d]. variation [HHT65c, UE65b]. various [SZ62]. varying [Coh66, Coh67]. vectors [CS65]. vegetative [DD66]. velocities [JD69, Kos62]. velocity [Ano66f, Cap66a, HH69]. versatile [McM63]. version [Lew66]. vertebrate [Bar63, Vay65]. Vertical [Str69]. vertices [Han67]. Vesicular [SKB69]. vestibulo [Beg63]. vestibulo-ocular [Beg63]. VI [BLT62a]. viii [Ano62h]. viral [Koc64]. virus [Kil66a, War64]. visceral [Bar63, Vay65]. vision [Lei66, Wil61b]. visual [Lei69a, Mal63]. vitalism [Dix68, Ozo69]. vitalistic [vdS69]. vitamine [VP65a]. vitamins [BBS63, PLB69]. vivo [Bro64, Reh68, Sug63, SF63, Sug66]. Volterra [Gar67a]. volume [Ano62c, Ano63c, Ano63d, Ano63h, Ano63i, Ano64a, Ano64g, Ano64h, Ano65a, Ano65b, Ano65f, Ano65g, Ano66g, Ano66h, Ano66m, Ano66n, Ano67b, Ano67c, Ano67j, Ano67k, Ano69h, Ano69i, BDF69, GGW68, JW67, Rob68, Spe67, TM68]. vulgaris [Mal63].

Waals [CPC66]. walks [She62b, SU69]. Water [CC65, Alm69, Dic64, ERH69b, HH66, Ids68, LM66, Løv63, RH69, Wal66b]. Wave [Arn61, Tak66, Tak68]. webs [O’N69]. weight [Spe67]. weighting [BR64]. weightlessness [Tyl66]. weights [AC63]. welches [Hel64]. wheat [Boy66]. which [AP65, RS67]. whose [Hel68a, Hel68b]. Wigglesworth [Mil66]. will [RS67]. within [Bla67a, Par64]. wobble [PR68b]. working [Wei62].

X [FBM66]. xanthine [PVP61, Per62a].

Zeitgeber [Wev66]. Zelman [Agi69]. zone [GB62a]. Zotin [IJ68]. Zotina [IJ68]. Zur [Jar64].

References

Akrigg:1969:UDC

[AAB69] A. Akrigg, S. R. Ayad, and J. Blamire. Uptake of DNA by competent bacteria — a possible mechanism. Journal of Theoretical Biology, 24(3):266–272, September 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

           URL http://www.sciencedirect.com/science/article/pii/
           S0022519369800512.

Andrewartha:1961:AIR

[AB61] H. G. Andrewartha and T. O. Browning. An analysis of the idea of “resources” in animal ecology. Journal of Theoretical Biology, 1(1): 83–97, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900285.

Angervall:1963:TCU

[AC63] Lennart Angervall and Esbjörn Carlström. Theoretical criteria for the use of relative organ weights and similar ratios in biology. Journal of Theoretical Biology, 4(3):254–259, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900043.

Acland:1966:ATC

[Acl66] J. D. Acland. Active transfer and the Curie Principle. Journal of Theoretical Biology, 13(??):318–323, December 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366900245.

Acland:1967:PCI

[Acl67] J. D. Acland. Predictive character of irreversible thermodynamics in active transfer. Journal of Theoretical Biology, 14(3):323– 324, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900847.

Acland:1969:ETA

[Acl69] J. D. Acland. Energy transformation in active transfer. Journal of Theoretical Biology, 22(3):535–536, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900204.

Amesz:1961:MMC

[ADB61] J. Amesz, L. N. M. Duysens, and D. C. Brandt. Methods for measuring and correcting the absorption spectrum of scattering suspensions. Journal of Theoretical Biology, 1(1):59–74, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519361900261.

Agin:1963:SCH

[Agi63] D. Agin. Some comments on the Hodgkin–Huxley equations. Journal of Theoretical Biology, 5(2):161–170, September 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519363900572.

Agin:1964:SCH

[Agi64] D. Agin. Some comments on the Hodgkin–Huxley equations — corrections. Journal of Theoretical Biology, 7(2):388, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900827.

Agin:1969:CZL

[Agi69] D. Agin. Comments on Zelman’s letter to the editor. Journal of Theoretical Biology, 22(3):533–534, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900198.

Aidley:1966:SRC

[Aid66] D. J. Aidley. A suggestion regarding the cellular basis of learning. Journal of Theoretical Biology, 11(2):343–345, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901718.

Ainsworth:1968:CSS

[Ain68] S. Ainsworth. A comparative study of some mechanisms describing interaction effects in a dimeric, two-substrate enzyme catalysed reaction. Journal of Theoretical Biology, 19(1):1–28, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900027.

Aladjem:1962:TAI

[AKT62] Frederick Aladjem, Henry Klostergaard, and Robert W. Taylor. Theoretical analysis of immunoelectrophoresis. Journal of Theoretical Biology, 3(1):134–145, July 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800101.

Almen:1969:CAD

[Alm69] Torsten Almen. Contrast agent design: Some aspects on the synthesis of water soluble contrast agents of low osmolality. Journal of Theoretical Biology, 24(2):216–226, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800470.

Altmann:1965:SRM

[Alt65] Stuart A. Altmann. Sociobiology of rhesus monkeys. II: Stochastics of social communication. Journal of Theoretical Biology, 8(3): 490–522, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936590024X.

Anonymous:1961:EB

[Ano61a] Anonymous. Editorial Board. Journal of Theoretical Biology, 1 (1):ifc, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900212.

Anonymous:1961:PJa

[Ano61b] Anonymous. Pages 1–106 (January 1961). Journal of Theoretical Biology, 1(1):??, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1961:PA

[Ano61c] Anonymous. Pages 107–287 (April 1961). Journal of Theoretical Biology, 1(2):??, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1961:PJb

[Ano61d] Anonymous. Pages 289–409 (July 1961). Journal of Theoretical Biology, 1(3):??, July 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1962:EBa

[Ano62a] Anonymous. Editorial Board. Journal of Theoretical Biology, 2 (1):i, January 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800216.

Anonymous:1962:EBb

[Ano62b] Anonymous. Editorial Board. Journal of Theoretical Biology, 3(1):i, July 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800010.

Anonymous:1962:IV

[Ano62c] Anonymous. Index to volume 3. Journal of Theoretical Biology, 3(3):v–vii, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519362800447.

Anonymous:1962:PJb

[Ano62d] Anonymous. Pages 1–147 (July 1962). Journal of Theoretical Biology, 3(1):??, July 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1962:PJa

[Ano62e] Anonymous. Pages 1–73 (January 1962). Journal of Theoretical Biology, 2(1):??, January 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1962:PS

[Ano62f] Anonymous. Pages 149–334 (September 1962). Journal of Theoretical Biology, 3(2):??, September 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1962:PMb

[Ano62g] Anonymous. Pages 193–346 (May 1962). Journal of Theoretical Biology, 2(3):??, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1962:PVV

[Ano62h] Anonymous. Pages 335–514, v-viii (November 1962). Journal of Theoretical Biology, 3(3):??, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1962:PMa

[Ano62i] Anonymous. Pages 75–191 (March 1962). Journal of Theoretical Biology, 2(2):??, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1963:A

[Ano63a] Anonymous. Announcement. Journal of Theoretical Biology, 4 (2):228, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900328.

Anonymous:1963:BTN

[Ano63b] Anonymous. Biomathematics training at North Carolina state College. Journal of Theoretical Biology, 4(1):145, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363901073.

Anonymous:1963:CVa

[Ano63c] Anonymous. Contents of volume 4. Journal of Theoretical Biology, 4(3):i–iv, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900092.

Anonymous:1963:CVb

[Ano63d] Anonymous. Contents of volume 5. Journal of Theoretical Biology, 5(3):III–IV, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936390095X.

Anonymous:1963:EBa

[Ano63e] Anonymous. Editorial Board. Journal of Theoretical Biology, 4 (1):ifc, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900961.

Anonymous:1963:EBb

[Ano63f] Anonymous. Editorial Board. Journal of Theoretical Biology, 5 (1):ifc, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936390033X.

Anonymous:1963:E

[Ano63g] Anonymous. Erratum. Journal of Theoretical Biology, 4(1): 146, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363901085.

Anonymous:1963:IVa

[Ano63h] Anonymous. Index to volume 4. Journal of Theoretical Biology, 4 (3):297–299, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900080.

Anonymous:1963:IVb

[Ano63i] Anonymous. Index to volume 5. Journal of Theoretical Biology, 5(3):500–502, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900948.

Anonymous:1963:PJa

[Ano63j] Anonymous. Pages 1–146 (January 1963). Journal of Theoretical Biology, 4(1):??, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1963:PJb

[Ano63k] Anonymous. Pages 1–160 (July 1963). Journal of Theoretical Biology, 5(1):??, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1963:PMa

[Ano63l] Anonymous. Pages 147–228 (March 1963). Journal of Theoretical Biology, 4(2):??, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1963:PS

[Ano63m] Anonymous. Pages 161–339 (September 1963). Journal of Theoretical Biology, 5(2):??, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1963:PMb

[Ano63n] Anonymous. Pages 229–299 (May 1963). Journal of Theoretical Biology, 4(3):??, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1963:PN

[Ano63o] Anonymous. Pages 341–502 (November 1963). Journal of Theoretical Biology, 5(3):??, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1964:CV

[Ano64a] Anonymous. Contents of volume. Journal of Theoretical Biology, 6(3):iii–iv, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900608.

Anonymous:1964:EBa

[Ano64b] Anonymous. Editorial Board. Journal of Theoretical Biology, 6 (1):ifc, January 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936490061X.

Anonymous:1964:EBb

[Ano64c] Anonymous. Editorial Board. Journal of Theoretical Biology, 7 (1):ifc, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900372.

Anonymous:1964:Ea

[Ano64d] Anonymous. Errat. Journal of Theoretical Biology, 6(2):306, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900360.

Anonymous:1964:Ec

[Ano64e] Anonymous. Errata. Journal of Theoretical Biology, 7(2):390– 391, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900840.

Anonymous:1964:Ed

[Ano64f] Anonymous. Errata. Journal of Theoretical Biology, 7(3):565, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900232.

Anonymous:1964:IVa

[Ano64g] Anonymous. Index to volume 6. Journal of Theoretical Biology, 6 (3):471–473, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900591.

Anonymous:1964:IVb

[Ano64h] Anonymous. Index to volume 7. Journal of Theoretical Biology, 7(3):567–569, 571–573, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900244.

Anonymous:1964:PJa

[Ano64i] Anonymous. Pages 1–136 (January 1964). Journal of Theoretical Biology, 6(1):??, January 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1964:PJb

[Ano64j] Anonymous. Pages 1–170 (July 1964). Journal of Theoretical Biology, 7(1):??, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1964:PMa

[Ano64k] Anonymous. Pages 137–309 (March 1964). Journal of Theoretical Biology, 6(2):??, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1964:PS

[Ano64l] Anonymous. Pages 171–391 (September 1964). Journal of Theoretical Biology, 7(2):??, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1964:PMb

[Ano64m] Anonymous. Pages 307–473 (May 1964). Journal of Theoretical Biology, 6(3):??, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1964:PN

[Ano64n] Anonymous. Pages 393–573 (November 1964). Journal of Theoretical Biology, 7(3):??, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1965:CVa

[Ano65a] Anonymous. Contents of volume 8. Journal of Theoretical Biology, 8(3):III–V, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900263.

Anonymous:1965:CVb

[Ano65b] Anonymous. Contents of volume 9. Journal of Theoretical Biology, 9(3):III–V, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900512.

Anonymous:1965:EBa

[Ano65c] Anonymous. Editorial Board. Journal of Theoretical Biology, 8 (1):ifc, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900858.

Anonymous:1965:EBb

[Ano65d] Anonymous. Editorial Board. Journal of Theoretical Biology, 9 (1):ifc, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900524.

Anonymous:1965:E

[Ano65e] Anonymous. Erratum. Journal of Theoretical Biology, 9(3):504, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900482.

Anonymous:1965:IVa

[Ano65f] Anonymous. Index to volume 8. Journal of Theoretical Biology, 8 (3):523–527, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900251.

Anonymous:1965:IVb

[Ano65g] Anonymous. Index to volume 9. Journal of Theoretical Biology, 9(3):507–510, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900500.

Anonymous:1965:PJb

[Ano65h] Anonymous. Pages 1–123 (July 1965). Journal of Theoretical Biology, 9(1):??, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1965:PJa

[Ano65i] Anonymous. Pages 1–220 (January 1965). Journal of Theoretical Biology, 8(1):??, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1965:PS

[Ano65j] Anonymous. Pages 125–350 (September 1965). Journal of Theoretical Biology, 9(2):??, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1965:PMa

[Ano65k] Anonymous. Pages 221–369 (March 1965). Journal of Theoretical Biology, 8(2):??, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1965:PN

[Ano65l] Anonymous. Pages 351–510 (November 1965). Journal of Theoretical Biology, 9(3):??, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1965:PMb

[Ano65m] Anonymous. Pages 371–527 (May 1965). Journal of Theoretical Biology, 8(3):??, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1965:TBC

[Ano65n] Anonymous. Theoretical biology colloquium. Journal of Theoretical Biology, 9(3):505, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900494.

Anonymous:1966:Ab

[Ano66a] Anonymous. Announcement. Journal of Theoretical Biology, 10 (3):508, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901408.

Anonymous:1966:Ac

[Ano66b] Anonymous. Announcement. Journal of Theoretical Biology, 11 (3):499, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901111.

Anonymous:1966:Ad

[Ano66c] Anonymous. Announcement. Journal of Theoretical Biology, 12(3): 442, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901573.

Anonymous:1966:Ae

[Ano66d] Anonymous. Announcement. Journal of Theoretical Biology, 13(??):433, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900312.

Anonymous:1966:Aa

[Ano66e] Anonymous. Announcements. Journal of Theoretical Biology, 10(2):384–385, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901342.

Anonymous:1966:CSR

[Ano66f] Anonymous. A characteristic of self-regulated linear energy converters the hill force–velocity relation for muscle: S. R. Caplan J. Theoret. Biol. (1966) 11, 63–86. Journal of Theoretical Biology, 12(1):155, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901949.

Anonymous:1966:CVa

[Ano66g] Anonymous. Contents of volume 10. Journal of Theoretical Biology, 10(3):I–IV, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901421.

Anonymous:1966:CVb

[Ano66h] Anonymous. Contents of volume 12. Journal of Theoretical Biology, 12(3):i–v, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901597.

Anonymous:1966:EBa

[Ano66i] Anonymous. Editorial Board. Journal of Theoretical Biology, 10 (1):ifc, January 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901731.

Anonymous:1966:EBb

[Ano66j] Anonymous. Editorial Board. Journal of Theoretical Biology, 11 (1):ifc, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900324.

Anonymous:1966:EBc

[Ano66k] Anonymous. Editorial Board. Journal of Theoretical Biology, 12 (1):ifc, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901822.

Anonymous:1966:EBd

[Ano66l] Anonymous. Editorial Board. Journal of Theoretical Biology, 13(??):ifc, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900014.

Anonymous:1966:IVa

[Ano66m] Anonymous. Index to volume 10. Journal of Theoretical Biology, 10(3):509–512, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936690141X.

Anonymous:1966:IVb

[Ano66n] Anonymous. Index to volume 12. Journal of Theoretical Biology, 12(3):443–446, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901585.

Anonymous:1966:PS

[Ano66o] Anonymous. Pages 1–155 (September 1966). Journal of Theoretical Biology, 12(1):??, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PJa

[Ano66p] Anonymous. Pages 1–158 (January 1966). Journal of Theoretical Biology, 10(1):??, January 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PM

[Ano66q] Anonymous. Pages 1–179 (May 1966). Journal of Theoretical Biology, 11(1):??, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PDb

[Ano66r] Anonymous. Pages 1–433 (December 1966). Journal of Theoretical Biology, 13(??):1–433, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PN

[Ano66s] Anonymous. Pages 157–290 (November 1966). Journal of Theoretical Biology, 12(2):??, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PF

[Ano66t] Anonymous. Pages 159–385 (February 1966). Journal of Theoretical Biology, 10(2):??, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PJb

[Ano66u] Anonymous. Pages 181–347 (July 1966). Journal of Theoretical Biology, 11(2):??, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PDa

[Ano66v] Anonymous. Pages 291–446 (December 1966). Journal of Theoretical Biology, 12(3):??, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PAb

[Ano66w] Anonymous. Pages 349–499 (August 1966). Journal of Theoretical Biology, 11(3):??, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1966:PAa

[Ano66x] Anonymous. Pages 387–512 (April 1966). Journal of Theoretical Biology, 10(3):??, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:A

[Ano67a] Anonymous. Announcement. Journal of Theoretical Biology, 14 (3):329, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900872.

Anonymous:1967:CVa

[Ano67b] Anonymous. Contents of volume 15. Journal of Theoretical Biology, 15(3):i–iv, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901452.

Anonymous:1967:CVb

[Ano67c] Anonymous. Contents of volume 16. Journal of Theoretical Biology, 16(3):i–v, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900690.

Anonymous:1967:EBa

[Ano67d] Anonymous. Editorial Board. Journal of Theoretical Biology, 14 (1):ifc, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900884.

Anonymous:1967:EBb

[Ano67e] Anonymous. Editorial Board. Journal of Theoretical Biology, 15 (1):ifc, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900392.

Anonymous:1967:EBc

[Ano67f] Anonymous. Editorial Board. Journal of Theoretical Biology, 16 (1):ifc, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900495.

Anonymous:1967:EBd

[Ano67g] Anonymous. Editorial Board. Journal of Theoretical Biology, 17 (1):ifc, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900136.

Anonymous:1967:Ea

[Ano67h] Anonymous. Errata. Journal of Theoretical Biology, 14(1): 102, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900963.

Anonymous:1967:Ec

[Ano67i] Anonymous. Errata. Journal of Theoretical Biology, 14(3): 328, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900860.

Anonymous:1967:IVa

[Ano67j] Anonymous. Index to volume 15. Journal of Theoretical Biology, 15 (3):377–380, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901440.

Anonymous:1967:IVb

[Ano67k] Anonymous. Index to volume 16. Journal of Theoretical Biology, 16(3):443–447, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900689.

Anonymous:1967:PJa

[Ano67l] Anonymous. Pages 1–102 (January 1967). Journal of Theoretical Biology, 14(1):??, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PJc

[Ano67m] Anonymous. Pages 1–150 (July 1967). Journal of Theoretical Biology, 16(1):??, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PAa

[Ano67n] Anonymous. Pages 1–177 (April 1967). Journal of Theoretical Biology, 15(1):??, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PO

[Ano67o] Anonymous. Pages 1–180 (October 1967). Journal of Theoretical Biology, 17(1):??, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PF

[Ano67p] Anonymous. Pages 103–224 (February 1967). Journal of Theoretical Biology, 14(2):??, February 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PAb

[Ano67q] Anonymous. Pages 151–331 (August 1967). Journal of Theoretical Biology, 16(2):??, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PMb

[Ano67r] Anonymous. Pages 179–272 (May 1967). Journal of Theoretical Biology, 15(2):??, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PN

[Ano67s] Anonymous. Pages 181–318 (November 1967). Journal of Theoretical Biology, 17(2):??, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PMa

[Ano67t] Anonymous. Pages 225–329 (March 1967). Journal of Theoretical Biology, 14(3):??, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PJb

[Ano67u] Anonymous. Pages 273–380 (June 1967). Journal of Theoretical Biology, 15(3):??, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PD

[Ano67v] Anonymous. Pages 319–497 (December 1967). Journal of Theoretical Biology, 17(3):??, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1967:PS

[Ano67w] Anonymous. Pages 333–447 (September 1967). Journal of Theoretical Biology, 16(3):??, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:EBa

[Ano68a] Anonymous. Editorial Board. Journal of Theoretical Biology, 18 (1):ifc, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901653.

Anonymous:1968:EBb

[Ano68b] Anonymous. Editorial Board. Journal of Theoretical Biology, 19 (1):ifc, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900015.

Anonymous:1968:EBc

[Ano68c] Anonymous. Editorial Board. Journal of Theoretical Biology, 20 (1):ifc, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936890088X.

Anonymous:1968:EBd

[Ano68d] Anonymous. Editorial Board. Journal of Theoretical Biology, 21 (1):ifc, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900544.

Anonymous:1968:E

[Ano68e] Anonymous. Errata. Journal of Theoretical Biology, 20(2): I, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901963.

Anonymous:1968:PJc

[Ano68f] Anonymous. Pages 1–127 (July 1968). Journal of Theoretical Biology, 20(1):??, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PO

[Ano68g] Anonymous. Pages 1–131 (October 1968). Journal of Theoretical Biology, 21(1):??, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PAa

[Ano68h] Anonymous. Pages 1–132 (April 1968). Journal of Theoretical Biology, 19(1):??, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PJa

[Ano68i] Anonymous. Pages 1–156 (January 1968). Journal of Theoretical Biology, 18(1):??, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PAb

[Ano68j] Anonymous. Pages 129–270 (August 1968). Journal of Theoretical Biology, 20(2):??, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PMb

[Ano68k] Anonymous. Pages 133–240 (May 1968). Journal of Theoretical Biology, 19(2):??, May 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PN

[Ano68l] Anonymous. Pages 133–292 (November 1968). Journal of Theoretical Biology, 21(2):??, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PF

[Ano68m] Anonymous. Pages 157–261 (February 1968). Journal of Theoretical Biology, 18(2):??, February 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PJb

[Ano68n] Anonymous. Pages 241–344 (June 1968). Journal of Theoretical Biology, 19(3):??, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PMa

[Ano68o] Anonymous. Pages 263–398 (March 1968). Journal of Theoretical Biology, 18(3):??, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PS

[Ano68p] Anonymous. Pages 271–358 (September 1968). Journal of Theoretical Biology, 20(3):??, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1968:PD

[Ano68q] Anonymous. Pages 293–452 (December 1968). Journal of Theoretical Biology, 21(3):??, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:Ab

[Ano69a] Anonymous. Announcement. Journal of Theoretical Biology, 25(3): 505, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800391.

Anonymous:1969:Aa

[Ano69b] Anonymous. Announcements. Journal of Theoretical Biology, 22(3):543–545, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936990023X.

Anonymous:1969:EBa

[Ano69c] Anonymous. Editorial Board. Journal of Theoretical Biology, 22 (1):ifc, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900745.

Anonymous:1969:EBb

[Ano69d] Anonymous. Editorial Board. Journal of Theoretical Biology, 23 (1):ifc, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900629.

Anonymous:1969:EBc

[Ano69e] Anonymous. Editorial Board. Journal of Theoretical Biology, 24(1):i, July 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800019.

Anonymous:1969:EBd

[Ano69f] Anonymous. Editorial Board. Journal of Theoretical Biology, 25 (1):i, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800147.

Anonymous:1969:FTD

[Ano69g] Anonymous. Formal theory of drug dosage regimens. I: E. Krüger– Thiemer, J. Theoret. Biol. (1966). 13, 212–235. Journal of Theoretical Biology, 22(1):I, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900885.

Anonymous:1969:IVa

[Ano69h] Anonymous. Index to volume 24. Journal of Theoretical Biology, 24 (1):351–354, July 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800020.

Anonymous:1969:IVb

[Ano69i] Anonymous. Index to volume 25. Journal of Theoretical Biology, 25(1):509–513, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519369800159.

Anonymous:1969:PAa

[Ano69j] Anonymous. Pages 1–168 (April 1969). Journal of Theoretical Biology, 23(1):??, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PJa

[Ano69k] Anonymous. Pages 1–207 (January 1969). Journal of Theoretical Biology, 22(1):??, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PAb

[Ano69l] Anonymous. Pages 131–245 (August 1969). Journal of Theoretical Biology, 24(2):??, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PMb

[Ano69m] Anonymous. Pages 169–346 (May 1969). Journal of Theoretical Biology, 23(2):??, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PN

[Ano69n] Anonymous. Pages 187–342 (November 1969). Journal of Theoretical Biology, 25(2):??, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PF

[Ano69o] Anonymous. Pages 209–379 (February 1969). Journal of Theoretical Biology, 22(2):??, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PS

[Ano69p] Anonymous. Pages 247–350 (September 1969). Journal of Theoretical Biology, 24(3):??, September 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PD

[Ano69q] Anonymous. Pages 343–505 (December 1969). Journal of Theoretical Biology, 25(3):??, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PJb

[Ano69r] Anonymous. Pages 347–495 (June 1969). Journal of Theoretical Biology, 23(3):??, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PJc

[Ano69s] Anonymous. Pages 351–354, 1-130 (July 1969). Journal of Theoretical Biology, 24(1):??, July 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PMa

[Ano69t] Anonymous. Pages 381–545 (March 1969). Journal of Theoretical Biology, 22(3):??, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1969:PO

[Ano69u] Anonymous. Pages 509–513, 1-185 (October 1969). Journal of Theoretical Biology, 25(1):??, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Aladjem:1965:AAR

[AP65] Frederick Aladjem and Michael T. Palmiter. The antigen-antibody reaction: V. A quantitative theory of antigen-antibody reactions which allows for heterogeneity of antibodies. Journal of Theoretical Biology, 8(1):8–21, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900871.

Arbib:1967:ATD

[Arb67] Michael A. Arbib. Automata theory and development: Part I. Journal of Theoretical Biology, 14(2):131–156, February 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901105.

Arnold:1961:WMM

[Arn61] William Arnold. Wave mechanics and muscular contraction. Journal of Theoretical Biology, 1(3):404–406, July 1961. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251936190039X.

Arnold:1968:SST

[Arn68] W. N. Arnold. The selection of sucrose as the translocate of higher plants. Journal of Theoretical Biology, 21(1):13–20, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900568.

Atassi:1966:SAA

[Ata66] M. Z. Atassi. Significance of the amino acid composition of proteins 1. Composition of hemoglobins and myoglobins in relation to their structure, function and evolution. Journal of Theoretical Biology, 11(2):227–241, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901639.

Atlan:1968:AIT

[Atl68] Henri Atlan. Application of information theory to the study of the stimulating effects of ionizing radiation, thermal energy, and other environmental factors preliminary ideas for a theory of organization. Journal of Theoretical Biology, 21(1):45– 70, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900593.

Apter:1965:CDI

[AW65] M. J. Apter and Lewis Wolpert. Cybernetics and development I. Information theory. Journal of Theoretical Biology, 8(2):244– 257, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900755.

Bowman:1963:MDH

[BA63] James D. Bowman and Frederick Aladjem. A method for the determination of heterogeneity of antibodies. Journal of Theoretical Biology, 4(3):242–253, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900031.

Baer:1964:PMH

[Bae64] Dee Baer. A possible mode of histone synthesis. Journal of Theoretical Biology, 6(2):282–289, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900347.

Barr:1962:NGP

[Bar62] H. J. Barr. A new genetical particle: The hypotheson. Journal of Theoretical Biology, 3(3):514, November 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800435.

Barr:1963:PVV

[Bar63] Lloyd Barr. Propagation in vertebrate visceral smooth muscle. Journal of Theoretical Biology, 4(1):73–85, January 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519363901012.

Barr:1965:MPP

[Bar65] L. Barr. Membrane potential profiles and the Goldman equation. Journal of Theoretical Biology, 9(3):351–356, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900366.

Bassham:1963:ECC

[Bas63] J. A. Bassham. Energy capture and conversion by photosynthesis. Journal of Theoretical Biology, 4(1):52–72, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363901000.

Bellman:1966:CSE

[BBKL66] R. Bellman, J. Buell, R. Kalaba, and J. Lockett. On the computational solution of an equation arising in chemotherapy using numerical inversion of the Laplace transform. Journal of Theoretical Biology, 11(2):334–337, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936690169X.

Brown:1963:HCM

[BBS63] R. Glenn Brown, Grace M. Button, and John T. Smith. A hypothesis for a common mode of action of the fat soluble vitamins A, D, and E. Journal of Theoretical Biology, 5(3):489– 492, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900912.

Baker:1964:AGC

[BCB64a] George A. Baker, J. Christy, and G. A. Baker. Analysis of genetic changes in finite populations composed of mixtures of pure lines. Journal of Theoretical Biology, 7(1):68–85, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900414.

Baker:1964:SPG

[BCB64b] George A. Baker, J. Christy, and G. A. Baker. Stochastic processes and genotypic frequencies under mixed selfing and random mating. Journal of Theoretical Biology, 7(1):86–97, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900426.

Bartley:1969:MMV

[BDF69] W. Bartley, B. Dean, and W. Ferdinand. Maintenance of mitochondrial volume and the effects of phosphate and ATP in producing swelling and shrinking. Journal of Theoretical Biology, 24(2):192– 202, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800457.

Begbie:1963:ECV

[Beg63] G. H. Begbie. Eye-centering and the vestibulo-ocular reflex in man. Journal of Theoretical Biology, 4(2):215–222, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900298.

Bell:1961:SEM

[Bel61] L. G. E. Bell. Surface extension as the mechanism of cellular movement and cell division. Journal of Theoretical Biology, 1(1):104– 106, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900303.

Bell:1962:PCM

[Bel62] L. G. E. Bell. Polysaccharide and cell membranes. Journal of Theoretical Biology, 3(1):132–133, July 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800095.

Bennett:1963:RMS

[Ben63] J. H. Bennett. Random mating and sex linkage. Journal of Theoretical Biology, 4(1):28–36, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900985.

Bergner:1961:TDT

[Ber61a] Per-Erik E. Bergner. Tracer dynamics: I. A tentative approach and definition of fundamental concepts. Journal of Theoretical Biology, 1(2):120–140, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936190042X.

Bergner:1961:TDI

[Ber61b] Per-Erik E. Bergner. Tracer dynamics: II. The limiting properties of the tracer system. Journal of Theoretical Biology, 1(3): 359–381, July 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900376.

Bergner:1962:SICa

[Ber62a] Per-Erik E. Bergner. On the stochastic interpretation of cell survival curves. Journal of Theoretical Biology, 2(3):279–295, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519362900310.

Bergner:1962:SICb

[Ber62b] Per-Erik E. Bergner. On the stochastic interpretation of cell survival curves — errata. Journal of Theoretical Biology, 3(3): 354, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800319.

Bernstein:1962:MIR

[Ber62c] Harris Bernstein. On the mechanism of intragenic recombination. I. The r II region of bacteriophage T 4. Journal of Theoretical Biology, 3(3):335–353, November 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800307.

Berman:1963:PAM

[Ber63] Mones Berman. A postulate to aid in model building. Journal of Theoretical Biology, 4(3):229–236, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900018.

Bergner:1964:TDD

[Ber64a] Per-Erik E. Bergner. Tracer dynamics and the determination of pool-sizes and turnover factors in metabolic systems. Journal of Theoretical Biology, 6(2):137–158, March 1964. CODEN JTBIAP.

ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900256.

Bernhard:1964:SSB

[Ber64b] R. Bernhard. Survey of some biological aspects of irreversible thermodynamics. Journal of Theoretical Biology, 7(3):532–557, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900207.

Bernhard:1964:TCK

[Ber64c] R. Bernhard. Towards a chemical kinetic brain model. Journal of Theoretical Biology, 6(2):244–257, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900311.

Bernstein:1964:MIR

[Ber64d] Harris Bernstein. On the mechanism of intragenic recombination II. Neurospora crassa. Journal of Theoretical Biology, 6(3): 347–370, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900529.

Bergner:1965:SNS

[Ber65] Per-Erik E. Bergner. On stationary and non-stationary red cell survival curves. Journal of Theoretical Biology, 9(3):366–388, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936590038X.

Bigelow:1967:AHP

[Big67] Charles C. Bigelow. On the average hydrophobicity of proteins and the relation between it and protein structure. Journal of Theoretical Biology, 16(2):187–211, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900045.

Bellman:1966:IPE

[BKK66] R. Bellman, H. Kagiwada, and R. Kalaba. Inverse problems in ecology. Journal of Theoretical Biology, 11(1):164–167, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900464.

Boyd:1969:ESE

[BL69] Donald B. Boyd and William N. Lipscomb. Electronic structures for energy-rich phosphates. Journal of Theoretical Biology, 25(3):403–420, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519369800299.

Black:1966:MEI

[Bla66] Francis L. Black. Measles endemicity in insular populations: Critical community size and its evolutionary implication. Journal of Theoretical Biology, 11(2):207–211, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901615.

Blakley:1967:DNS

[Bla67a] G. R. Blakley. Darwinian natural selection acting within populations. Journal of Theoretical Biology, 17(2):252–281, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901713.

Blatt:1967:EBP

[Bla67b] John M. Blatt. Enzymatic break-up of polypeptides as a stochastic process. Journal of Theoretical Biology, 17(2):282–303, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901725.

Bleecken:1969:DBC

[Ble69] S. Bleecken. Duplication of the bacterial cell and its initiation. Journal of Theoretical Biology, 25(1):137–158, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800214.

Baker:1962:NCAb

[BLT62a] B. R. Baker, William W. Lee, and Ethel Tong. Non-classical antimetabolites. VI. 4-(iodoacetamido)-salicylic acid, an exoalkylating irreversible inhibitor. Journal of Theoretical Biology, 3(3):459–483, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251936280037X.

Baker:1962:NCAa

[BLT^{+}62b] B. R. Baker, William W. Lee, Ethel Tong, Leonard O. Ross, and Abelardo P. Martinez. Non-classical antimetabolites. V. Further factors in the design of exo-alkylating enzyme inhibitors, particularly of lactic dehydrogenase. Journal of Theoretical Biology, 3(3):446–458, November 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800368.

Blumenthal:1962:CCP

[Blu62] Linda K. Blumenthal. A contribution to the coding problem. Journal of Theoretical Biology, 2(1):72–73, January 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800290.

Blumberg:1968:LGR

[Blu68] A. A. Blumberg. Logistic growth rate functions. Journal of Theoretical Biology, 21(1):42–44, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900581.

Bailey:1962:IBH

[BNW62] V. A. Bailey, A. J. Nicholson, and E. J. Williams. Interaction between hosts and parasites when some host individuals are more difficult to find than others. Journal of Theoretical Biology, 3 (1):1–18, July 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800022.

Bennett:1965:ARA

[BO65] J. H. Bennett and Carolyn R. Oertel. The approach to a random association of genotypes with random mating. Journal of Theoretical Biology, 9(1):67–76, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900573.

Bock:1967:PTS

[Boc67] R. M. Bock. Prediction of a topaz suppressor ribonucleic acid. Journal of Theoretical Biology, 16(3):438–439, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900665.

Bodmer:1963:NSM

[Bod63] Walter F. Bodmer. Natural selection for modifiers of heterozygote fitness. Journal of Theoretical Biology, 4(1):86–97, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363901024.

Boettcher:1963:SES

[Boe63] B. Boettcher. A suggested explanation of some cases of high negative interference and certain mapping anomalies in phage t4b. Journal of Theoretical Biology, 5(1):108–113, July 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519363900390.

Bossert:1968:TPO

[Bos68] William H. Bossert. Temporal patterning in olfactory communication. Journal of Theoretical Biology, 18(2):157–170, February 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901586.

Boyden:1962:CDB

[Boy62] Stephen V. Boyden. Cellular discrimination between indigenous and foreign matter. Journal of Theoretical Biology, 3(1):123– 131, July 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800083.

Boyd:1966:PII

[Boy66] W. J. R. Boyd. Problems in the interpretation of the inheritance of leaf rust resistance in the common wheat. Journal of Theoretical Biology, 13(??):283–294, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900221.

Balassa:1962:ETT

[BP62] Georges Balassa and Georges Prevost. Etude Théorique de la transformation Bactérienne. Journal of Theoretical Biology, 3(2):315– 334, September 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800204.

Beck:1964:DPO

[BR64] James S. Beck and Aldo Rescigno. Determination of precursor order and particular weighting functions from kinetic data. Journal of Theoretical Biology, 6(1):1–12, January 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900621.

Bellman:1966:SDA

[BR66] R. Bellman and R. S. Roth. Segmental differential approximation and biological systems: An analysis of a metabolic process. Journal of Theoretical Biology, 11(1):168–176, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900476.

Braitenberg:1965:WCL

[Bra65] V. Braitenberg. What can be learned from spike interval histograms about synaptic mechanisms. Journal of Theoretical Biology, 8(3):419–425, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900202.

Brannen:1968:CRD

[Bra68] J. P. Brannen. Concerning reversible denaturation of RNA and memory. Journal of Theoretical Biology, 20(3):358, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901367.

Bresler:1967:CAT

[Bre67] Emanuel H. Bresler. On criteria for active transport. Journal of Theoretical Biology, 16(1):135–146, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900574.

Britton:1966:FPM

[Bri66] H. G. Britton. Fluxes in passive, monovalent and polyvalent carrier systems. Journal of Theoretical Biology, 10(1):28–52, January 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901779.

Bright:1969:CAC

[Bri69] Peter B. Bright. Concerning “anisotropic” contributions of chemical reactions to flow equations and active transport. Journal of Theoretical Biology, 23(1):135–157, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900721.

Brlow:1964:INB

[Brl64] John S. Brlow. Inertial navigation as a basis for animal navigation. Journal of Theoretical Biology, 6(1):76–117, January 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900670.

Browning:1962:EAP

[Bro62] T. O. Browning. The environments of animals and plants. Journal of Theoretical Biology, 2(1):63–68, January 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800277.

Brown:1963:CGC

[Bro63] A. P. Brown. Cohesion and growth of cells in a tissue. Journal of Theoretical Biology, 5(3):372–388, November 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519363900845.

Brody:1964:EES

[Bro64] S. S. Brody. An estimate of the effective sizes of chlorophyll a aggregates in vivo as determined from emission spectra. Journal of Theoretical Biology, 7(2):352–359, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900797.

Bronk:1969:RLP

[Bro69] B. V. Bronk. On radioactive labeling of proliferating cells: The graph of labeled mitosis. Journal of Theoretical Biology, 22(3): 468–492, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900162.

Blizzard:1969:CCE

[BS69] A. C. Blizzard and D. P. Santry. CNDO/2 calculations of the electronic structure of the guanine-cytosine base pair: Proton transfer in excited states. Journal of Theoretical Biology, 25(3):461– 472, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800330.

Bingley:1962:BPR

[BT62] M. S. Bingley and C. M. Thompson. Bioelectric potentials in relation to movement in amoebae. Journal of Theoretical Biology, 2(1): 16–32, January 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800241.

Buerger:1966:MPI

[Bue66] A. A. Buerger. A model for the penetration of integuments by nonelectrolytes. Journal of Theoretical Biology, 11(1):131– 139, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900439.

Buerger:1967:TIP

[Bue67] A. A. Buerger. A theory of integumental penetration. Journal of Theoretical Biology, 14(1):66–73, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936790094X.

Burch:1966:SAI

[Bur66] P. R. J. Burch. Spontaneous auto-immunity: Equations for age-specific prevalence and initiation-rates. Journal of Theoretical Biology, 12(3):397–409, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901524.

Bossert:1963:AOC

[BW63] William H. Bossert and Edward O. Wilson. The analysis of olfactory communication among animals. Journal of Theoretical Biology, 5(3):443–469, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900894.

Boulton:1969:ICM

[BW69] D. M. Boulton and C. S. Wallace. The information content of a multistate distribution. Journal of Theoretical Biology, 23(2): 269–278, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900411.

Caianiello:1961:OTT

[Cai61] E. R. Caianiello. Outline of a theory of thought-processes and thinking machines. Journal of Theoretical Biology, 1(2):204– 235, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900467.

Calvin:1961:QCP

[Cal61] Melvin Calvin. Quantum conversion in photosynthesis. Journal of Theoretical Biology, 1(2):258–262, IN1–IN2, 263–287, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519361900492.

Calogero:1968:SIS

[Cal68] F. Calogero. Solution of infinite systems of differential equations of possible interest in evolutionary theories. Journal of Theoretical Biology, 19(1):79–89, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900064.

Campbell:1967:PTN

[Cam67] John Campbell. Permissive translation of nucleic acid: a mechanism for antibody synthesis. Journal of Theoretical Biology, 16(2): 321–331, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900124.

Caplan:1966:CSR

[Cap66a] S. R. Caplan. A characteristic of self-regulated linear energy converters the Hill force–velocity relation for muscle. Journal of Theoretical Biology, 11(1):63–86, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900403.

Caplan:1966:DCR

[Cap66b] S. R. Caplan. The degree of coupling and its relation to efficiency of energy conversion in multiple-flow systems. Journal of Theoretical Biology, 11(2):346–347, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936690172X.

Caplani:1966:DCR

[Cap66c] S. R. Caplani. The degree of coupling and its relation to efficiency of energy conversion in multiple-flow systems. Journal of Theoretical Biology, 10(2):209–235, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936690124X.

Carlson:1964:MCM

[Car64] Elof Axel Carlson. A method for calculating the maximum number of mutational events in Drosophila melanogaster. Journal of Theoretical Biology, 6(3):432–440, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900578.

Claringbold:1961:ERF

[CB61] P. J. Claringbold and J. S. F. Barker. The estimation of relative fitness of drosophila populations. Journal of Theoretical Biology, 1(2): 190–203, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900455.

Coe:1965:HRO

[CC65] Elmon L. Coe and Mary H. Coe. A hypothesis relating oil: Water partition coefficients and vapor pressures of nonelectrolytes to their penetration rates through biological membranes. Journal of Theoretical Biology, 8(2):327–343, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900810.

Cherry:1969:ODT

[CC69] R. J. Cherry and D. Chapman. Optical determination of the thickness of thin lipid films. Journal of Theoretical Biology, 24(2):137– 146, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251936980041X.

Cullinan:1969:TSM

[CD69] H. T. Cullinan and G. B. DiPirro. A transition state model for active transport. Journal of Theoretical Biology, 23(1):11– 22, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900642.

Cennamo:1968:SSK

[Cen68] C. Cennamo. Steady-state kinetics of one-substrate enzymic mechanisms involving two enzyme conformations I. Effects of modifiers on a mechanism postulating a single enzyme-substrate complex. Journal of Theoretical Biology, 21(2):260–277, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936890074X.

Cennamo:1969:SSK

[Cen69] C. Cennamo. Steady-state kinetics of one-substrate enzymic mechanisms involving two enzyme conformations: II. Kinetic treatment of a reversible mechanism: Effects of modifiers and product inhibition. Journal of Theoretical Biology, 23(1):53–71, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900678.

Chen:1969:HCS

[CF69] R. L. W. Chen and P. Fong. Helical configuration of singlestranded DNA. Journal of Theoretical Biology, 22(1):180–187, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900861.

Chun:1968:DEC

[CFY68] P. W. Chun, M. Fried, and K. S. Yee. Determination of the equilibrium constants of associating protein systems I. the application of stepwise polynomial regression analysis. Journal of Theoretical Biology, 19(2):147–158, May 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901100.

Cho:1967:KAIa

[CGTK67a] Chai Moo Cho, Joseph Greco, Peter J. Talso, and Ervin Kaplan. Kinetic analysis of isotope exchange from a cylinder. Journal of Theoretical Biology, 16(2):151–158, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936790001X.

Cho:1967:KAIb

[CGTK67b] Chai Moo Cho, Joseph Greco, Peter J. Talso, and Ervin Kaplan. Kinetic analysis of isotope exchange from a cylinder containing dissociable non-permeable interacting substance. Journal of Theoretical Biology, 16(2):159–165, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900021.

Chaffee:1966:ESS

[Cha66] R. R. J. Chaffee. On experimental selection for super-hibernating and non-hibernating lines of Syrian hamsters. Journal of Theoretical Biology, 12(1):151–154, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901937.

Chiaraviglio:1965:SMC

[Chi65] Lucio Chiaraviglio. Sequential machines and the coding of polypeptides. Journal of Theoretical Biology, 8(1):130–140, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900998.

Cornew:1967:FCH

[CHS67] Ronald W. Cornew, James C. Houk, and Lawrence Stark. Fine control in the human temperature regulation system. Journal of Theoretical Biology, 16(3):406–426, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900641.

Cinatl:1969:IOM

[Cin69] Jaroslav Cinátl. Inorganic-organic multimolecular complexes of salt solutions, culture media and biological fluids and their possible significance for the origin of life. Journal of Theoretical Biology, 23(1):1–8, IN1–IN6, 9–10, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900630.

Clayton:1963:AES

[Cla63] Roderick K. Clayton. An automatic energy switch in photosynthesis. Journal of Theoretical Biology, 5(3):497–499, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900936.

Claxton:1964:DPS

[Cla64] J. H. Claxton. The determination of patterns with special reference to that of the central primary skin follicles in sheep. Journal of Theoretical Biology, 7(2):302–317, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900748.

Clayton:1967:ARB

[Cla67] Roderick K. Clayton. An analysis of the relations between fluorescence and photochemistry during photosynthesis. Journal of Theoretical Biology, 14(2):173–186, February 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901129.

Chalvet:1962:SAQ

[CM62] O. Chalvet and R. Mason. Some aspects of quantum biology: a review. Journal of Theoretical Biology, 3(1):51–81, July 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519362800046.

Coe:1968:MLS

[Coe68] Fredric L. Coe. Mean life in steady-state populations. Journal of Theoretical Biology, 18(2):171–180, February 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368901598.

Cohen:1966:ORR

[Coh66] Dan Cohen. Optimizing reproduction in a randomly varying environment. Journal of Theoretical Biology, 12(1):119–129, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901883.

Cohen:1967:ORR

[Coh67] Dan Cohen. Optimizing reproduction in a randomly varying environment when a correlation may exist between the conditions at the time a choice has to be made and the subsequent outcome. Journal of Theoretical Biology, 16(1):1–14, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900501.

Cohen:1969:NMI

[Coh69] Dan Cohen. A necessary mechanism for the intracellular degradation of metabolically active macromolecules. Journal of Theoretical Biology, 24(1):126–127, July 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800123.

Cone:1969:STA

[Con69] C. D. Cone. Some theoretical aspects of intercellular bridges as a potential mechanism of cancerous proliferation. Journal of Theoretical Biology, 22(2):365–379, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900101.

Coren:1965:UCI

[Cor65] Richard L. Coren. Uterine contraction and intra-amniotic pressure. Journal of Theoretical Biology, 9(1):1–15, July 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365900536.

Coulson:1969:RPL

[Cou69] A. S. Coulson. Recognition pathway in lymphocytes. Journal of Theoretical Biology, 25(1):127–136, October 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800202.

Claverie:1966:VWL

[CPC66] Pierre Claverie, Bernard Pullman, and Jacqueline Caillet. Van der Waals–London interactions between stacked purines and pyrimidines. Journal of Theoretical Biology, 12(3):419–434, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901548.

Cocordano:1966:RBE

[CRJ66] Maurice Cocordano, Jacques Ricard, and André Julg. Relations between the electronic structure and auxinic activity of chlorinated benzoic acids. Journal of Theoretical Biology, 12(3):291– 296, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901433.

Crosby:1963:END

[Cro63] J. L. Crosby. The evolution and nature of dominance. Journal of Theoretical Biology, 5(1):35–51, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900353.

Crowe:1968:MMM

[Cro68] Alan Crowe. A mechanical model of the mammalian muscle spindle. Journal of Theoretical Biology, 21(1):21–41, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936890057X.

Case:1965:ASF

[CS65] Carl T. Case and Manuel Schwartz. The application of a system of fixed-rotating vectors to circuit analysis and synthesis. Journal of Theoretical Biology, 8(2):221–232, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900731.

Cairns-Smith:1966:OLN

[CS66] A. G. Cairns-Smith. The origin of life and the nature of the primitive gene. Journal of Theoretical Biology, 10(1):53–88, January 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901780.

Coulson:1968:TRS

[CS68] A. S. Coulson and N. V. Smith. The tuberculin receptor site on sensitized human lymphocytes. Journal of Theoretical Biology, 20(1): 112–116, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900969.

Changeux:1967:MAC

[CT67] J. P. Changeux and J. Thiéry. On the mode of action of colicins: a model of regulation at the membrane level. Journal of Theoretical Biology, 17(2):315–318, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901750.

Currey:1968:ABS

[Cur68] J. D. Currey. The adaptation of bones to stress. Journal of Theoretical Biology, 20(1):91–106, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900945.

Czerlinski:1964:ACRa

[Cze64a] Georg Czerlinski. Application of chemical relaxation to biochemical systems: I. Isolated reactions. Journal of Theoretical Biology, 7(3):435–462, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900165.

Czerlinski:1964:ACRb

[Cze64b] Georg Czerlinski. Application of chemical relaxation to biochemical systems: II. Two-step-reactions. Journal of Theoretical Biology, 7(3):463–484, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900177.

Czerlinski:1967:ACR

[Cze67] George Czerlinski. Application of chemical relaxation to biological systems: III. Selected three-step reactions. Journal of Theoretical Biology, 17(3):343–382, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900987.

Czerlinski:1968:CRCa

[Cze68a] G. H. Czerlinski. Chemical relaxation of cyclic enzyme reactions: I. General kinetic treatment of three-step mechanisms. Journal of Theoretical Biology, 21(3):387–397, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901227.

Czerlinski:1968:CRCb

[Cze68b] G. H. Czerlinski. Chemical relaxation of cyclic enzyme reactions: II. General kinetic treatment of four-step mechanisms. Journal of Theoretical Biology, 21(3):398–407, December 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368901239.

Czerlinski:1968:CRCc

[Cze68c] G. H. Czerlinski. Chemical relaxation of cyclic enzyme reactions: III. Experimental implication of previous results. Journal of Theoretical Biology, 21(3):408–420, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901240.

DAdamo:1963:SPH

[D’A63] Amedeo F. D’Adamo. Shunt pathways and high energy phosphate synthesis. Journal of Theoretical Biology, 4(1):142–144, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363901061.

Danielli:1961:P

[Dan61] J. F. Danielli. Preface. Journal of Theoretical Biology, 1(1): i, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900224.

Daniel:1964:FCI

[Dan64] Vera Daniel. Field controlled ion trapping as a mechanism in nerve membranes. Journal of Theoretical Biology, 6(3):375–385, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900542.

Danielli:1966:TLM

[Dan66] T. F. Danielli. On the thickness of lipid membranes. Journal of Theoretical Biology, 12(3):439–441, December 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366901561.

Danielli:1967:Ea

[Dan67a] J. F. Danielli. Editorial. Journal of Theoretical Biology, 15 (2):179, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901981.

Danielli:1967:Eb

[Dan67b] J. F. Danielli. Editorial. Journal of Theoretical Biology, 15(3): 273, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901361.

Danchin:1969:BMI

[Dan69] A. Danchin. Binding of metal ions to macromolecules through an NMR spectrometric method of investigation. Journal of Theoretical Biology, 25(2):317–330, November 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800664.

Darvey:1968:TPK

[Dar68] I. G. Darvey. Transient phase kinetics of enzyme reactions. Journal of Theoretical Biology, 19(2):215–231, May 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251936890115X.

Darvey:1969:PFS

[Dar69] I. G. Darvey. Prediction of the form of steady-state rate equations from kinetic data. Journal of Theoretical Biology, 25(1):109– 112, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800184.

Davies:1965:TQO

[Dav65] J. T. Davies. A theory of the quality of odours. Journal of Theoretical Biology, 8(1):1–7, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936590086X.

Davidovits:1966:QFS

[Dav66] J. Davidovits. The quaternary and fibrillar structure of native collagen. Journal of Theoretical Biology, 12(1):1–6, IN2, 7–11, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901834.

Dawson:1966:IAD

[Daw66] P. S. Dawson. Insect antennal development and the serrate gene in tribolium castaneum. Journal of Theoretical Biology, 12(1):133– 139, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901901.

Dayhoff:1965:CAP

[Day65] M. O. Dayhoff. Computer aids to protein sequence determination. Journal of Theoretical Biology, 8(1):97–112, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900962.

deDuve:1964:PTF

[dD64] Christian de Duve. Principles of tissue fractionation. Journal of Theoretical Biology, 6(1):33–40, IN2, 41–59, January 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900657.

Drobnik:1966:SME

[DD66] Jaroslav Drobník and Jarmila Dlouhá. Statistical model of evolution of haploid organisms during simple vegetative reproduction. Journal of Theoretical Biology, 11(3):418–435, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901020.

DeRocco:1967:GTH

[De 67] Andrew G. De Rocco. On a generalization of the theorem of hamaker and de Boer: The interaction of spherical colloids in solution. Journal of Theoretical Biology, 15(1):34–40, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900410.

DeLey:1969:CND

[De 69] J. De Ley. Compositional nucleotide distribution and the theoretical prediction of homology in bacterial DNA. Journal of Theoretical Biology, 22(1):89–116, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900824.

DeLong:1967:SB

[DeL67] Richard DeLong. On sexuality in bacteria. Journal of Theoretical Biology, 16(3):333–336, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900598.

Despopoulos:1965:DSS

[Des65] Agamemnon Despopoulos. A definition of substrate specificity in renal transport of organic anions. Journal of Theoretical Biology, 8(1):163–186, IN1–IN3, 187–192, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365901013.

Deutsch:1963:LES

[Deu63] J. A. Deutsch. Learning and electrical self-stimulation of the brain. Journal of Theoretical Biology, 4(2):193–214, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900286.

Dewan:1964:NOE

[Dew64] Edmond M. Dewan. Nonlinear oscillations and electroencephalography. Journal of Theoretical Biology, 7(1):141–159, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519364900475.

Lima-de-Faria:1962:SML

[dF62] A. Lima de Faria. Selection at the molecular level. Journal of Theoretical Biology, 2(1):7–15, January 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251936280023X.

Dainty:1963:ITF

[DG63] J. Dainty and B. Z. Ginzburg. Irreversible thermodynamics and frictional models of membrane processes, with particular reference to the cell membrane. Journal of Theoretical Biology, 5(2):256– 265, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900638.

Degn:1969:TOR

[DH69] H. Degn and D. E. F. Harrisson. Theory of oscillations of respiration rate in continuous culture of klebsiella aerogenes. Journal of Theoretical Biology, 22(2):238–248, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900034.

Diamondstone:1969:DBB

[Dia69] Thomas I. Diamondstone. On the distinction between binding of substrate and kinetic order of addition in enzyme-catalyzed reactions. Journal of Theoretical Biology, 25(3):347–364, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800251.

Dick:1964:PCW

[Dic64] D. A. T. Dick. The permeability coefficient of water in the cell membrane and the diffusion coefficient in the cell interior. Journal of Theoretical Biology, 7(3):504–531, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900190.

DiStefano:1969:MNT

[DiS69] J. J. DiStefano. A model of the normal thyroid hormone glandular secretion mechanism. Journal of Theoretical Biology, 22(3): 412–417, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900137.

Dix:1968:DV

[Dix68] Douglas Dix. A defense of vitalism. Journal of Theoretical Biology, 20(3):338–340, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901331.

Deering:1966:STC

[DKGK66] R. A. Deering, Alfred A. Kraus, D. M. Gray, and Rein Kilkson. Some theoretical considerations of polarized light effects on unoriented biological systems. Journal of Theoretical Biology, 13(??):295–317, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900233.

Day:1969:MSS

[DL69a] C. E. Day and R. S. Levy. Molecular structure of serum lipoproteins. Journal of Theoretical Biology, 23(3):387–399, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900277.

Day:1969:GFR

[DL69b] Charles E. Day and Robert S. Levy. A generalized functional role for phospholipids. Journal of Theoretical Biology, 22(3):541– 542, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900228.

deLuca:1967:FNP

[dLR67] A. de Luca and L. M. Ricciardi. Formalized neuron: Probabilistic description and asymptotic theorems. Journal of Theoretical Biology, 14(2):206–217, February 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901142.

Doetsch:1966:SSA

[Doe66] R. N. Doetsch. Some speculations accounting for the movement of bacterial flagella. Journal of Theoretical Biology, 11(3):411– 417, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901019.

Doerner:1967:PGC

[Doe67] Karl Doerner. Pressure gradients in cells. Journal of Theoretical Biology, 14(3):284–292, March 1967. CODEN JTBIAP.

ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900811.

Doerner:1968:SDC

[Doe68] Karl Doerner. Stresses and duration of cellular contact. Journal of Theoretical Biology, 19(3):275–286, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901392.

Doerner:1969:PFC

[Doe69] Karl F. Doerner. Propagation of forces in cells. Journal of Theoretical Biology, 25(2):226–235, November 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800615.

Dounce:1962:NHN

[Dou62] A. L. Dounce. A new hypothesis for nucleic acid replication. Journal of Theoretical Biology, 2(2):152–158, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519362900437.

Dilworth:1969:DNF

[DP69] M. J. Dilworth and C. A. Parker. Development of the nitrogenfixing system in legumes. Journal of Theoretical Biology, 25(2):208– 218, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800597.

Darvey:1966:DNI

[DPW66a] I. G. Darvey, S. J. Prokhovnik, and J. F. Williams. The determination of the number of intermediates in enzyme catalysed reactions. Journal of Theoretical Biology, 13(??):90–99, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900105.

Darvey:1966:UPS

[DPW66b] I. G. Darvey, S. J. Prokhovnik, and J. F. Williams. The use of power series solutions in the kinetics of enzyme reactions with one intermediate. Journal of Theoretical Biology, 11(3):459– 464, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901056.

Dikstein:1966:BSR

[DS66] S. Dikstein and F. G. Sulman. Biochemistry of stimulus-response-recovery cycle and its relationship to the pharmacological receptor. Journal of Theoretical Biology, 12(3):385–396, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901512.

Darvey:1967:ATM

[DS67] I. G. Darvey and P. J. Staff. The application of the theory of Markov processes to the reversible one substrate-one intermediateone product enzymic mechanism. Journal of Theoretical Biology, 14(2):157–172, February 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901117.

DiStefano:1968:IHH

[DS68] Joseph J. DiStefano and Edwin B. Stear. On identification of hypothalamo-hypophysial control and feedback relationships with the thyroid gland. Journal of Theoretical Biology, 19(1): 29–50, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900039.

Dalrymple:1968:DCM

[DSB68] G. V. Dalrymple, J. L. Sanders, and M. L. Baker. Do cultured mammalian cells repair radiation injury by the “cut-and-patch” mechanism? Journal of Theoretical Biology, 21(3):368– 386, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901215.

Duncan:1963:EMC

[Dun63] C. J. Duncan. Excitatory mechanisms in chemo- and mechanoreceptors. Journal of Theoretical Biology, 5(1):114–126, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900407.

Duncan:1965:CPC

[Dun65] C. J. Duncan. Cation-permeability control and depolarization in excitable cells. Journal of Theoretical Biology, 8(3):403–418, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900196.

Drobnik:1966:NMD

[DV66] J. Drobník and V. Vondrejs. A new model of DNA replication. Journal of Theoretical Biology, 12(2):251–259, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901160.

Eberhardt:1969:SAF

[Ebe69] L. L. Eberhardt. Similarity, allometry and food chains. Journal of Theoretical Biology, 24(1):43–55, July 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800056.

Eck:1962:PCN

[Eck62] Richard V. Eck. The protein cryptogram I. Non-random occurrence of amino acid “alleles”. Journal of Theoretical Biology, 2(2): 139–151, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519362900425.

Evans:1966:IDM

[EDT66] Robert L. Evans, Raymond L. Duncan, and John V. Tyberg. Indicator dilution measurements of almost-simultaneous regional blood flows in the dog. Journal of Theoretical Biology, 10(3): 490–507, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901391.

Everitt:1968:ECL

[EH68] C. T. Everitt and D. A. Haydon. Electrical capacitance of a lipid membrane separating two aqueous phases. Journal of Theoretical Biology, 18(3):371–379, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900842.

Everitt:1969:IDL

[EH69] C. T. Everitt and D. A. Haydon. Influence of diffusion layers during osmotic flow across bimolecular lipid membranes. Journal of Theoretical Biology, 22(1):9–19, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900769.

Ehret:1967:PTS

[Ehr67] C. F. Ehret. Paratene theory of the shapes of cells. Journal of Theoretical Biology, 15(2):263–272, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936790207X.

Essig:1966:NFT

[EKH66] A. Essig, O. Kedem, and T. L. Hill. Net flow and tracer flow in lattice and carrier models. Journal of Theoretical Biology, 13(??): 72–89, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900099.

Elliott:1968:FBS

[Ell68] G. F. Elliott. Force-balances and stability in hexagonally-packed polyelectrolyte systems. Journal of Theoretical Biology, 21(1):71– 87, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936890060X.

Elsasser:1961:QCO

[Els61] Walter M. Elsasser. Quanta and the concept of organismic law. Journal of Theoretical Biology, 1(1):27–58, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936190025X.

Elsasser:1962:PAN

[Els62] Walter M. Elsasser. Physical aspects of non-mechanistic biological theory. Journal of Theoretical Biology, 3(2):164–191, September 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800137.

Elsasser:1963:NEO

[Els63] Walter M. Elsasser. Note on evolution in organismic theory. Journal of Theoretical Biology, 4(2):166–174, March 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519363900250.

Elsasser:1964:SOT

[Els64] Walter M. Elsasser. Synopsis of organismic theory. Journal of Theoretical Biology, 7(1):53–67, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900402.

Elsasser:1969:MEG

[Els69] Walter M. Elsasser. The mathematical expression of generalized complementarity. Journal of Theoretical Biology, 25(2):276– 296, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800640.

Emlen:1966:NSH

[Eml66] J. Merritt Emlen. Natural selection and human behavior. Journal of Theoretical Biology, 12(3):410–418, December 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366901536.

Epstein:1968:NSE

[EMM68] J. Epstein, H. O. Michel, and W. A. Mosher. On the nucleophilicity of serine in enzymes. Journal of Theoretical Biology, 19(3): 320–326, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901434.

Engelberg:1962:ARI

[Eng62] Joseph Engelberg. Analysis of a radiation-induced mutation in E. coli. Journal of Theoretical Biology, 2(3):312–325, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519362900346.

Engelberg:1968:DOM

[Eng68] Joseph Engelberg. On deterministic origins of mitotic variability. Journal of Theoretical Biology, 20(2):249–259, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936890194X.

Enright:1965:SRB

[Enr65] J. T. Enright. The search for rhythmicity in biological time-series. Journal of Theoretical Biology, 8(3):426–468, May 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365900214.

Enright:1967:SNS

[Enr67] J. T. Enright. The spontaneous neuron subject to tonic stimulation. Journal of Theoretical Biology, 16(1):54–77, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900537.

Erhan:1969:UDM

[Erh69a] Semih Erhan. The unwinding of the DNA molecule during replication. Journal of Theoretical Biology, 23(2):339–341, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900447.

Everitt:1969:PBL

[ERH69b] C. T. Everitt, W. R. Redwood, and D. A. Haydon. Problem of boundary layers in the exchange diffusion of water across bimolecular lipid membranes. Journal of Theoretical Biology, 22(1):20– 32, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900770.

Essig:1966:IIA

[Ess66] A. Essig. Isotope interaction and “abnormal” flux ratios — a frictional model. Journal of Theoretical Biology, 13(??):63–71, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900087.

Estabrook:1966:MMG

[Est66] G. F. Estabrook. A mathematical model in graph theory for biological classification. Journal of Theoretical Biology, 12(3):297– 310, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901445.

Estabrook:1968:GSP

[Est68] G. F. Estabrook. A general solution in partial orders for the Camin–Sokal model in phylogeny. Journal of Theoretical Biology, 21(3):421–438, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901252.

Ehret:1967:MMC

[ET67] C. F. Ehret and E. Trucco. Molecular models for the circadian clock: I. The chronon concept. Journal of Theoretical Biology, 15 (2):240–262, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367902068.

Evans:1962:UAB

[Eva62] Robert L. Evans. A unifying approach to blood flow theory. Journal of Theoretical Biology, 3(3):392–411, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519362800332.

Eylar:1966:BRG

[Eyl66] Edwin H. Eylar. On the biological role of glycoproteins. Journal of Theoretical Biology, 10(1):89–113, January 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366901792.

Fang:1969:ATR

[Fan69] P. H. Fang. Analysis of tuberculin receptor site on sensitized human lymphocytes. Journal of Theoretical Biology, 24(1):128– 130, July 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800135.

Friedenberg:1966:EMSa

[FBG^{+}66a] R. Friedenberg, A. Blatt, V. Gallucci, J. F. Danielli, and I. Shames. Electrostatics of membrane systems — a non-statistical approach to cellular membrane systems. Journal of Theoretical Biology, 11(3):465–477, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901068.

Friedenberg:1966:EMSb

[FBG66b] R. Friedenberg, A. I. Blatt, and V. Galluccl. Electrostatics of membrane systems — II Fixed charge and dipole planar surfaces of finite dimensions. Journal of Theoretical Biology, 11(3):478– 484, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936690107X.

Friedenberg:1966:PEF

[FBG66c] R. Friedenberg, A. J. Blatt, and V. Gallucci. The potential energy functions of idealized models of fixed charge and dipole distributions as related to surface chemical phenomena. Journal of Theoretical Biology, 11(3):485–489, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901081.

Fraser:1966:SGS

[FBM66] Alex Fraser, Donald Burnell, and David Miller. Simulation of genetic systems X. Inversion polymorphism. Journal of Theoretical Biology, 13(??):1–14, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900026.

Filmer:1967:MDR

[FCR67] David L. Filmer, John R. Cannon, and N. Reiss. A method for determination of rate constants in enzyme reactions. Journal of Theoretical Biology, 16(2):280–293, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900094.

Feibleman:1967:PG

[Fei67] James K. Feibleman. The philosophy of gestation. Journal of Theoretical Biology, 16(3):356–365, September 1967. CODEN JTBIAP.

ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900616.

Fellman:1968:CAQ

[Fel68] J. H. Fellman. Comments on the action of quaternary ammonium compounds at the receptor site. Journal of Theoretical Biology, 20 (1):124–125, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900982.

Fisher:1962:ASB

[Fis62a] Harvey F. Fisher. The appearance of substrate binding terms in the V_{max} expression of enzyme reactions. Journal of Theoretical Biology, 3(3):484–486, November 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800381.

Fisher:1962:DSD

[Fis62b] R. A. Fisher. The detection of a sex difference in recombination values using double heterozygotes. Journal of Theoretical Biology, 3(3):509–513, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800423.

Fisher:1962:ECP

[Fis62c] Ronald A. Fisher. Enumeration and classification in polysomic inheritance. Journal of Theoretical Biology, 2(3):309–311, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519362900334.

Fisher:1962:LES

[Fis62d] Ronald A. Fisher. Letter to the Editor: Self-sterility alleles: a reply to Professor D. Lewis. Journal of Theoretical Biology, 3(1): 146–147, July 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800113.

Fitzgerald:1964:IRL

[Fit64] P. H. Fitzgerald. The immunological role and long life-span of small lymphocytes. Journal of Theoretical Biology, 6(1):13– 25, January 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900633.

Finlayson:1969:AEF

[FM69] Birdwell Finlayson and George H. Miller. An analysis of experimental foreign body stone formation. Journal of Theoretical Biology, 23 (2):344–346, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900460.

Fong:1967:DBH

[Fon67a] Peter Fong. Dynamical basis of the helical structure of DNA. Journal of Theoretical Biology, 15(2):225–229, May 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367902032.

Fong:1967:PDM

[Fon67b] Peter Fong. Packing of the DNA molecule. Journal of Theoretical Biology, 15(2):230–235, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367902044.

Fong:1968:PTL

[Fon68] P. Fong. Phenomenological theory of life. Journal of Theoretical Biology, 21(2):133–152, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900672.

Forsdyke:1969:TI

[For69] D. R. Forsdyke. A theory of immunity. Journal of Theoretical Biology, 25(1):173–185, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519369800238.

Franck:1964:TLU

[FR64] James Franck and J. L. Rosenberg. A theory of light utilization in plant photosynthesis. Journal of Theoretical Biology, 7(2):276– 301, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900736.

Fraser:1962:SGS

[Fra62] A. S. Fraser. Simulation of genetic systems. Journal of Theoretical Biology, 2(3):329–346, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936290036X.

Freese:1962:EBC

[Fre62a] Ernst Freese. On the evolution of the base composition of DNA. Journal of Theoretical Biology, 3(1):82–101, July 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800058.

Frenster:1962:MDM

[Fre62b] J. H. Frenster. The magnitude of disease as measured by tolerance tests. Journal of Theoretical Biology, 2(2):159–164, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519362900449.

Fredericq:1963:NCF

[Fre63] Pierre Fredericq. On the nature of colicinogenic factors: a review. Journal of Theoretical Biology, 4(2):159–165, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900249.

Fried:1969:CTE

[Fri69a] Jerrold Fried. Correction for threshold error in the determination of generation times by the grain-count halving method. Journal of Theoretical Biology, 24(1):108–120, July 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800093.

Friedman:1969:ACF

[Fri69b] M. H. Friedman. On an analysis of the constant field equation. Journal of Theoretical Biology, 25(3):502–504, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251936980038X.

Frumento:1965:EEI

[Fru65] Antonio S. Frumento. The electrical effects of an ionic pump. Journal of Theoretical Biology, 9(2):253–262, September 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365901116.

Flake:1968:NCT

[FT68] R. H. Flake and B. L. Turner. Numerical classification for taxonomic problems. Journal of Theoretical Biology, 20(2):260– 270, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901951.

Fujimoto:1963:KMG

[Fuj63] Yoshinori Fujimoto. Kinetics of microbial growth and substrate consumption. Journal of Theoretical Biology, 5(2):171–191, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900584.

Friedlander:1965:SAF

[FW65] S. K. Friedlander and Mackenzie Walser. Some aspects of flow and diffusion in the proximal tubule of the kidney. Journal of Theoretical Biology, 8(1):87–96, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900950.

Garfinkel:1962:CSS

[Gar62] David Garfinkel. Computer simulation of steady-state glutamate metabolism in rat brain. Journal of Theoretical Biology, 3(3):412– 422, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800344.

Garfinkel:1967:ESL

[Gar67a] David Garfinkel. Effect on stability of Lotka–Volterra ecological systems of imposing strict territorial limits on populations. Journal of Theoretical Biology, 14(3):325–327, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900859.

Garfinkel:1967:SSE

[Gar67b] David Garfinkel. A simulation study of the effect on simple ecological systems of making rate of increase of population densitydependent. Journal of Theoretical Biology, 14(1):46–58, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900926.

Garrick:1967:KTM

[Gar67c] Michael D. Garrick. The kinetics of the translation of messenger RNA into protein. Journal of Theoretical Biology, 17(1):19– 30, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900161.

Gatlin:1963:TFD

[Gat63] L. L. Gatlin. Triplet frequencies in DNA and the genetic program. Journal of Theoretical Biology, 5(3):360–371, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900833.

Gatlin:1964:BCH

[Gat64] L. L. Gatlin. Base composition hyperbolic functional relationships in DNA. Journal of Theoretical Biology, 7(1):129–140, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900463.

Gatlin:1966:ICD

[Gat66] L. L. Gatlin. The information content of DNA. Journal of Theoretical Biology, 10(2):281–300, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901275.

Gatlin:1968:ICD

[Gat68] L. L. Gatlin. The information content of DNA. II. Journal of Theoretical Biology, 18(2):181–194, February 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901604.

Gaunt:1964:ITD

[Gau64] Paul Gaunt. Increment thresholds and dark adaptation. Journal of Theoretical Biology, 7(1):160–167, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900487.

Giddings:1962:ESM

[GB62a] J. Calvin Giddings and James R. Boyack. Effect of stabilizing media on zone mobility and spreading in electrophoresis. Journal of Theoretical Biology, 2(1):1–6, January 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800228.

Gosselin:1962:DTS

[GB62b] R. E. Gosselin and W. O. Berndt. Diffusional transport of solutes through mesentery and peritoneum. Journal of Theoretical Biology, 3(3):487–495, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800393.

Grainger:1966:MGS

[GB66] J. N. R. Grainger and L. Bass. A model of a growing steady state system. Journal of Theoretical Biology, 10(3):387–398, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901354.

Gilbert:1968:TSE

[GC68] M. Gilbert and P. Claverie. A theoretical study of the electrostatic interactions in the intercalation model of the DNA-dye complex. Journal of Theoretical Biology, 18(3):330–349, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900829.

Goodwin:1969:PSM

[GC69] Brian C. Goodwin and Morrel H. Cohen. A phase-shift model for the spatial and temporal organization of developing systems. Journal of Theoretical Biology, 25(1):49–107, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800172.

Gatlin:1963:FRB

[GD63] L. L. Gatlin and J. C. Davis. A functional relationship between base sequence in DNA and amino acid composition of protein. Journal of Theoretical Biology, 5(2):249–255, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900626.

Gunther:1966:UTB

[GdlB66] B. Günther and B. León de la Barra. A unified theory of biological similarities. Journal of Theoretical Biology, 13(??):48– 59, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900063.

Geduldig:1968:AMP

[Ged68] Donald Geduldig. Analysis of membrane permeability coefficient ratios and internal ion concentrations from a constant field equation. Journal of Theoretical Biology, 19(1):67–78, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900052.

Georgiev:1969:SOO

[Geo69] G. P. Georgiev. On the structural organization of operon and the regulation of RNA synthesis in animal cells. Journal of Theoretical Biology, 25(3):473–490, December 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800342.

Goel:1968:SEA

[GFR68] Narendra S. Goel, Nobuo Fukuda, and Robert Rein. A semiempirical analysis of the melting curves of synthetic DNA molecules and the calculation of the stacking and pairing energies and entropies in DNA. Journal of Theoretical Biology, 18(3):350– 370, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900830.

Gershfeld:1967:LTP

[GG67] N. L. Gershfeld and R. J. Good. Line tension and the penetration of a cell membrane by an oil drop. Journal of Theoretical Biology, 17(2):246, IN6, 247–251, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901701.

Grainger:1968:MGS

[GGW68] J. N. R. Grainger, P. E. Gaffney, and T. T. West. A model of a growing steady-state system with a changing surface-volume ratio. Journal of Theoretical Biology, 21(1):123–130, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900659.

Ginzburg:1967:WDS

[GH67] B. Z. Ginzburg and J. Hogg. What does a short-circuit current measure in biological systems? Journal of Theoretical Biology, 14(3): 316–322, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900835.

Ghent:1966:MCB

[Ghe66] Arthur W. Ghent. A model of crossing-over based on alternatelyoriented dipoles. Journal of Theoretical Biology, 11(1):87–111, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900415.

Gilvarry:1964:PPL

[Gil64] J. J. Gilvarry. The possibility of a pristine lunar life. Journal of Theoretical Biology, 6(3):325–346, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900517.

Gilbert:1968:DOC

[Gil68] D. A. Gilbert. Differentiation, oncogenesis and cellular periodicities. Journal of Theoretical Biology, 21(1):113–122, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900647.

Gingell:1967:MSP

[Gin67] D. Gingell. Membrane surface potential in relation to a possible mechanism for intercellular interactions and cellular responses: a physical basis. Journal of Theoretical Biology, 17(3):451–482, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901075.

Gingell:1968:CSP

[Gin68] D. Gingell. Computed surface potential changes with membrane interaction. Journal of Theoretical Biology, 19(3):340– 344, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901471.

Gifford:1969:IDR

[GK69] George E. Gifford and Arthur L. Koch. The interferon dose response curve and its possible significance. Journal of Theoretical Biology, 22(2):271–283, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900058.

Gerst:1965:KPS

[GL65] I. Gerst and S. N. Levine. Kinetics of protein synthesis by polyribosomes. Journal of Theoretical Biology, 9(1):16–36, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900548.

Gunther:1968:OTT

[GM68] B. Günther and C. Martinoya. Operational time and theory of biological similarities. Journal of Theoretical Biology, 20(1): 107–111, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900957.

Goel:1969:EIM

[GM69] Narendra S. Goel and Samaresh C. Maitra. On the existence of independently melting subunits in a copolymeric DNA and their statistical distribution. Journal of Theoretical Biology, 23 (1):87–98, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900691.

Goel:1967:CPG

[Goe67a] Narendra S. Goel. Comments on a paper by Gatlin “On the base composition hyperbolic functional relationship in DNA”. Journal of Theoretical Biology, 16(3):440–441, September 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367900677.

Goel:1967:KRC

[Goe67b] Narendra S. Goel. Kinetics of replication of copolymeric DNA. Journal of Theoretical Biology, 15(3):325–345, June 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367901415.

Goffman:1966:MMD

[Gof66] William Goffman. A mathematical model for describing the compatibility of infectious diseases. Journal of Theoretical Biology, 11(3):349–361, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936690097X.

Gold:1969:KFD

[Gol69] Harvey J. Gold. Kinetics of facilitated diffusion of oxygen in tissue slices. Journal of Theoretical Biology, 23(3):455–462, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900319.

Grabe:1964:RMD

[Gra64] Barbro Grabe. A revised and more detailed hypothesis for the mechanism of oxidative phosphorylation at the diphosphopyridine nucleotide–flavin level. Journal of Theoretical Biology, 7(1): 112–122, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936490044X.

Green:1966:ASC

[Gre66a] William A. Green. Aspects of some co-factor requirements in tumour metabolism. Journal of Theoretical Biology, 13(??):412– 428, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900294.

Green:1966:LMF

[Gre66b] William A. Green. Is lipid metabolism a fundamental feature in tumour-host relationship? Journal of Theoretical Biology, 11(2): 212–226, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901627.

Griffith:1968:MCCa

[Gri68a] J. S. Griffith. Mathematics of cellular control processes I. Negative feedback to one gene. Journal of Theoretical Biology, 20(2): 202–208, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901896.

Griffith:1968:MCCb

[Gri68b] J. S. Griffith. Mathematics of cellular control processes II. Positive feedback to one gene. Journal of Theoretical Biology, 20(2): 209–216, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901902.

Gross:1967:AMA

[Gro67] Leo Gross. Active membranes for active transport. Journal of Theoretical Biology, 15(3):298–306, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901397.

Grossberg:1969:PRC

[Gro69] S. Grossberg. On the production and release of chemical transmitters and related topics in cellular control. Journal of Theoretical Biology, 22(2):325–364, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900095.

Haarhoff:1969:UMN

[Haa69] K. N. Haarhoff. Use of multivariate non-linear regression analysis in fitting enzyme kinetic models: An empirical study of the inhibition of aspartate aminotransferase by dicarboxylic acid substrate analogues. Journal of Theoretical Biology, 22(1):117– 150, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900836.

Hadjiolov:1967:EHM

[Had67] Asen A. Hadjiolov. An endomessenger hypothesis on the mechanism of information transfer in animal cells. Journal of Theoretical Biology, 16(2):229–250, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900069.

Hall:1961:PTS

[Hal61] G. G. Hall. Permeation through a spherical membrane. Journal of Theoretical Biology, 1(1):18–26, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900248.

Hamilton:1964:GESa

[Ham64a] W. D. Hamilton. The genetical evolution of social behaviour. I. Journal of Theoretical Biology, 7(1):1–16, July 1964. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519364900384.

Hamilton:1964:GESb

[Ham64b] W. D. Hamilton. The genetical evolution of social behaviour. II. Journal of Theoretical Biology, 7(1):17–52, July 1964. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519364900396.

Hamilton:1966:MSN

[Ham66] W. D. Hamilton. The moulding of senescence by natural selection. Journal of Theoretical Biology, 12(1):12–45, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901846.

Hancock:1967:OUP

[Han67] J. E. H. Hancock. Open unbranched pathways through the vertices of an icosahedron. Journal of Theoretical Biology, 14(1):35– 45, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900914.

Hansche:1969:TBE

[Han69] P. E. Hansche. A theoretical basis for the entrainment of chemostat populations. Journal of Theoretical Biology, 24(3):335–350, September 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800573.

Hartley:1963:CMD

[Har63] G. S. Hartley. A contribution to the mathematics of dosage transfer. Journal of Theoretical Biology, 5(1):57–101, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900377.

Hasofer:1966:CTM

[Has66a] A. M. Hasofer. A continuous-time model in population genetics. Journal of Theoretical Biology, 11(1):150–163, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900452.

Hasofer:1966:STS

[Has66b] A. M. Hasofer. A simplified treatment of Spetner’s natural selection model. Journal of Theoretical Biology, 11(2):338–342, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901706.

Havsteen:1966:SCB

[Hav66] B. H. Havsteen. A study of the correlation between the amino acid composition and the helical content of proteins. Journal of Theoretical Biology, 10(1):1–10, January 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901743.

Houk:1966:MAA

[HCS66] James C. Houk, Ronald W. Cornew, and Lawrence Stark. A model of adaptation in amphibian spindle receptors. Journal of Theoretical Biology, 12(2):196–215, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901135.

Hirsch:1965:ACC

[HE65a] Henry R. Hirsch and Joseph Engelberg. An analog of cellculture growth. Journal of Theoretical Biology, 9(2):303–312, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365901153.

Hirsch:1965:DCD

[HE65b] Henry R. Hirsch and Joseph Engelberg. Determination of the cell doubling-time distribution from culture growth-rate data. Journal of Theoretical Biology, 9(2):297–302, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365901141.

Hebborn:1968:SAA

[Heb68] P. Hebborn. On the selective antitumor activity of some alkylating carbamates. Journal of Theoretical Biology, 21(3):449– 452, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901276.

Hedges:1966:ESH

[Hed66] A. J. Hedges. An examination of single-hit and multi-hit hypotheses in relation to the possible kinetics of colicin adsorption. Journal of Theoretical Biology, 11(3):383–410, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901007.

Heddle:1969:IFT

[Hed69] John A. Heddle. Influence of false twins on the ratios of twin and single sister chromatid exchanges. Journal of Theoretical Biology, 22(1):151–162, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900848.

Heinmets:1964:ACA

[Hei64] F. Heinmets. Analog computer analysis of a model-system for the induced enzyme synthesis. Journal of Theoretical Biology, 6(1): 60–75, January 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900669.

Hellwig:1964:EPW

[Hel64] Gerhard Hellwig. Über ein einfaches prinzip, welches die entropieerzeugung von lebewesen betrifft. Journal of Theoretical Biology, 6(2):258–274, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900323.

Hellwig:1968:CBCa

[Hel68a] Gerhard Hellwig. The connection between the cell environment and the production of molecules whose synthesis requires energy I. The formulation of a general law. Journal of Theoretical Biology, 20 (1):1–29, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900891.

Hellwig:1968:CBCb

[Hel68b] Gerhard Hellwig. The connection between the cell environment and the production of molecules whose synthesis requires energy II. The discussion of the general law. Journal of Theoretical Biology, 20 (1):30–55, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900908.

Henshaw:1968:TDH

[Hen68] Robert E. Henshaw. Thermoregulation during hibernation: Application of Newton’s law of cooling. Journal of Theoretical Biology, 20(1):79–90, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900933.

Hersh:1962:MTC

[Her62] R. T. Hersh. Mutants of TMV and the commaless code. Journal of Theoretical Biology, 2(3):326–328, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519362900358.

Herman:1969:CAD

[Her69] G. T. Herman. Computing ability of a developmental model for filamentous organisms. Journal of Theoretical Biology, 25(3):421– 435, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800305.

Hanai:1966:PWB

[HH66] T. Hanai and D. A. Haydon. The permeability to water of bimolecular lipid membranes. Journal of Theoretical Biology, 11(3): 370–382, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900993.

Heyde:1969:SAO

[HH69] C. C. Heyde and Elizabeth Heyde. A stochastic approach to a one substrate, one product enzyme reaction in the initial velocity phase. Journal of Theoretical Biology, 25(1):159–172, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800226.

Hanai:1965:ILC

[HHT65a] T. Hanai, D. A. Haydon, and Janet Taylor. The influence of lipid composition and of some adsorbed proteins on the capacitance of black hydrocarbon membranes. Journal of Theoretical Biology, 9(3):422–432, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936590041X.

Hanai:1965:PGO

[HHT65b] T. Hanai, D. A. Haydon, and Janet Taylor. Polar group orientation and the electrical properties of lecithin bimolecular leaflets. Journal of Theoretical Biology, 9(2):278–296, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936590113X.

Hanai:1965:VCC

[HHT65c] T. Hanai, D. A. Haydon, and Janet Taylor. The variation of capacitance and conductance of bimolecular lipid membranes with area. Journal of Theoretical Biology, 9(3):433–443, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900421.

Higgins:1968:SRS

[Hig68] J. Higgins. Some remarks on Shear’s Liapunov function for systems of chemical reactions. Journal of Theoretical Biology, 21(3):293– 304, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901173.

Hill:1966:SITb

[Hil66] Terrell L. Hill. Studies in irreversible thermodynamics IV. diagrammatic representation of steady state fluxes for unimolecular systems. Journal of Theoretical Biology, 10(3):442–459, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901378.

Hill:1966:SITa

[HK66] Terrell L. Hill and Ora Kedem. Studies in irreversible thermodynamics III. Models for steady state and active transport across membranes. Journal of Theoretical Biology, 10(3):399– 441, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901366.

Hunter:1968:SMM

[HKG68] G. D. Hunter, R. H. Kimberlin, and R. A. Gibbons. Scrapie: a modified membrane hypothesis. Journal of Theoretical Biology, 20(3):355–357, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901355.

Hoffmann:1964:SRH

[HL64] T. A. Hoffmann and J. Ladik. Some remarks on the hypochromicity of polynucleotides. Journal of Theoretical Biology, 6(1):26– 32, January 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900645.

Hartl:1968:PEN

[HM68] D. L. Hartl and T. Maruyama. Phenogram enumeration: The number of regular genotype-phenotype correspondences in genetic systems. Journal of Theoretical Biology, 20(2):129–163, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901860.

Hawkins:1963:TAT

[HOD63] Rosemary Hawkins, L. N. Owen, and J. F. Danielli. The toxicity and anti-tumour action of some alkylating sulphonamides. Journal of Theoretical Biology, 5(2):236–248, September 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519363900614.

Holden:1966:RFT

[Hol66a] A. L. Holden. Receptive fields and the two-point threshold. Journal of Theoretical Biology, 11(1):22–29, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936690035X.

Holgate:1966:TLD

[Hol66b] P. Holgate. Two limit distributions in evolutionary genetics. Journal of Theoretical Biology, 11(3):362–369, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900981.

Holgate:1967:PSL

[Hol67] P. Holgate. Population survival and life history phenomena. Journal of Theoretical Biology, 14(1):1–10, January 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367900896.

Holden:1968:ETS

[Hol68] J. T. Holden. Evolution of transport systems. Journal of Theoretical Biology, 21(1):97–102, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900623.

Homer:1967:ROT

[Hom67] Louis D. Homer. The receptor occupation theory of drug responses. Journal of Theoretical Biology, 17(3):399–409, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901002.

Horrobin:1969:FSC

[Hor69] D. F. Horrobin. The female sex cycle. Journal of Theoretical Biology, 22(1):80–88, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900812.

Hill:1962:C

[HPW62] J. A. Hill, J. M. Pratt, and R. J. P. Williams. The corphyrins. Journal of Theoretical Biology, 3(3):423–445, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519362800356.

Hausmann:1966:ENF

[HR66] Ernest Hausmann and Douglas S. Riggs. The effectiveness of negative feedback in regulating plasma calcium in the dog. Journal of Theoretical Biology, 12(3):350–363, December 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366901482.

Hilgartner:1969:PLSa

[HR69a] C. A. Hilgartner and John F. Randolph. Psycho-logics 2. The structure of ‘unimpaired’ human behavior. Journal of Theoretical Biology, 23(3):347–374, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900241.

Hilgartner:1969:PLSb

[HR69b] C. A. Hilgartner and John F. Randolph. Psycho-logics: 3. The structure of empathy. Journal of Theoretical Biology, 24(1): 1–29, July 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800032.

Hilgartner:1969:PLA

[HR69c] C. A. Hilgartner and John F. Randolph. Psycho-logics: An axiomatic system describing human behavior. Journal of Theoretical Biology, 23(2):285–338, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900435.

Haydon:1963:SPB

[HT63] D. A. Haydon and Janet Taylor. The stability and properties of bimolecular lipid leaflets in aqueous solutions. Journal of Theoretical Biology, 4(3):281–296, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900079.

Hughes:1968:CPL

[Hug68] A. Hughes. Comments on the paper by lea “permeation through long narrow pores”. Journal of Theoretical Biology, 21(1):131, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900660.

Hulett:1969:LPS

[Hul69] H. R. Hulett. Limitations on prebiological synthesis. Journal of Theoretical Biology, 24(1):56–72, July 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800068.

Hagemeijer:1969:LGA

[HvR69] F. Hagemeijer and P. van Remoortere. Limitations of the graphical analysis of elution curves. Journal of Theoretical Biology, 25(2):236–254, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519369800627.

Hogg:1969:MEP

[HWJ69] J. Hogg, E. J. Williams, and R. J. Johnston. The membrane electrical parameters of nitella translucens. Journal of Theoretical Biology, 24(3):317–334, September 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800561.

Idso:1968:ASI

[Ids68] S. B. Idso. Atmospheric- and soil-induced water stresses in plants and their effects on transpiration and photosynthesis. Journal of Theoretical Biology, 21(1):1–12, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900556.

Idso:1968:CPZ

[IJ68] Sherwood B. Idso and Ray D. Jackson. Comments on paper by zotin and zotina “thermodynamic aspects of developmental biology”. Journal of Theoretical Biology, 20(1):126–127, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900994.

Imamura:1966:SES

[IKTN66] Akira Imamura, Masahiko Kodama, Yusaku Tagashira, and Chikayoshi Nagata. The semi-empirical SCF treatment of σ–π system I. Peptide molecule. Journal of Theoretical Biology, 10(2):356– 369, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901329.

Jahn:1962:TEC

[Jah62] Theodore Louis Jahn. A theory of electronic conduction through membranes, and of active transport of ions, based on redox transmembrane potentials. Journal of Theoretical Biology, 2(2):129– 138, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519362900413.

Janoff:1964:RIM

[Jan64] Aaron Janoff. The role of iron in macrophages. Journal of Theoretical Biology, 7(1):168–170, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900499.

Jarosch:1964:QTP

[Jar64] R. Jarosch. Zur quantitativen theorie der Proteinschrauben– Ganghöhen von Bakteriengeißeln. (German) [The quantitative theory of protein screws — bacterial flagella pitches]. Journal of Theoretical Biology, 7(2):171–180, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900694.

Johnston:1969:IRE

[JD69] M. M. Johnston and W. F. Diven. An integrated rate equation for determining initial velocities. Journal of Theoretical Biology, 25(2):331–338, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800676.

Jehle:1967:SAC

[Jeh67] Herbert Jehle. Structural aspects of the coding problem. Journal of Theoretical Biology, 15(2):236–238, IN2–IN4, 239, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367902056.

Johnson:1969:MMS

[JH69] R. N. Johnson and G. R. Hanna. Membrane model: a single transistor analog of excitable membrane. Journal of Theoretical Biology, 22(3):401–411, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900125.

Johansen:1969:CES

[JI69] Helge Johansen and Lloyd L. Ingraham. Calculation of the electronic structures and spectra for corrin and related compounds. Journal of Theoretical Biology, 23(2):191–204, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900368.

Jacquez:1966:BSC

[JM66] John A. Jacquez and Frances J. Mather. Balance studies on compartmental systems with stochastic inputs. Journal of Theoretical Biology, 11(3):446–458, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901044.

Jennrich:1969:MNC

[JT69] R. I. Jennrich and F. B. Turner. Measurement of non-circular home range. Journal of Theoretical Biology, 22(2):227–237, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900022.

Johnson:1967:OVC

[JW67] John A. Johnson and Theodore A. Wilson. Osmotic volume changes induced by a permeable solute. Journal of Theoretical Biology, 17(2):304–311, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901737.

Kallenbach:1968:RDB

[Kal68] Neville R. Kallenbach. Replication of DNA in bacteria with heterogeneous generation times. Journal of Theoretical Biology, 18(1): 19–33, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901689.

Karlin:1967:APM

[Kar67] Arthur Karlin. On the application of “a plausible model” of allosteric proteins to the receptor for acetylcholine. Journal of Theoretical Biology, 16(2):306–320, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900112.

Katsumata:1967:RIM

[Kat67] Masuyuki Katsumata. Rearrangement of the Ingraham–Makower equation for a two-substrate enzyme reaction for practical application. Journal of Theoretical Biology, 17(1):108–120, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900239.

Katsumata:1969:IEA

[Kat69] Masuyuki Katsumata. Influence of enzyme activators and inhibitors, present in an enzyme preparation, on the relations between reaction rate and enzyme concentration. Journal of Theoretical Biology, 24(3):294–306, September 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800548.

Kauffman:1967:SDR

[Kau67] Stuart A. Kauffman. Sequential DNA replication and the control of differences in gene activity between sister chromatids — a possible factor in cell differentiation. Journal of Theoretical Biology, 17(3):483–497, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901087.

Kauffman:1969:MSE

[Kau69] S. A. Kauffman. Metabolic stability and epigenesis in randomly constructed genetic nets. Journal of Theoretical Biology, 22(3): 437–467, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900150.

Kavanau:1963:NTA

[Kav63] J. Lee Kavanau. A new theory of amoeboid locomotion. Journal of Theoretical Biology, 4(1):124–141, January 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251936390105X.

Koch:1969:POS

[KCE69] A. L. Koch, A. Carr, and D. W. Ehrenfeld. The problem of open-sea navigation: The migration of the green turtle to ascension Island. Journal of Theoretical Biology, 22(1):163–179, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936990085X.

Kelman:1965:MAS

[Kel65] R. B. Kelman. Mathematical analysis of sodium reabsorption in proximal part of nephron in presence of nonreabsorbed solute. Journal of Theoretical Biology, 8(1):22–26, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900883.

Keleti:1967:L

[Kel67] T. Keleti. The liberator. Journal of Theoretical Biology, 16(3):337– 355, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900604.

Kiefer:1968:MFC

[Kie68] Jürgen Kiefer. A model of feedback-controlled cell populations. Journal of Theoretical Biology, 18(3):263–279, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900787.

Kilkson:1966:ASV

[Kil66a] Rein Kilkson. Antigenic specificity of virus structural components. Journal of Theoretical Biology, 12(3):435–438, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936690155X.

Kilkson:1966:SEC

[Kil66b] Rein Kilkson. The structure and evolution of chromosomes. Journal of Theoretical Biology, 13(??):357–378, December 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366900270.

Kimizuka:1966:ICP

[Kim66] Hideo Kimizuka. Ion current and potential across membranes. Journal of Theoretical Biology, 13(??):145–163, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366900142.

King:1964:PNM

[Kin64] M. H. King. A possible new method for producing specific antisera. Journal of Theoretical Biology, 7(3):562–564, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900220.

Kimizuka:1964:ITT

[KK64] H. Kimizuka and K. Koketsu. Ion transport through cell membrane. Journal of Theoretical Biology, 6(2):290–305, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900359.

Klachko:1966:RIM

[Kla66] David M. Klachko. Regulation of intermediary metabolism by nucleotides: The mechanism of action of insulin. Journal of Theoretical Biology, 12(2):267–275, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901184.

Klima:1966:IBN

[Kli66] Jörg Klima. Informationsfluß beim nachbau von kettenmolekülen. (German) [Information flow in the reconstruction of chain molecules]. Journal of Theoretical Biology, 12(3):342–349, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901470.

Koch:1965:MKM

[KM65] Arthur L. Koch and Constance Miller. A mechanism for keeping mutations in check. Journal of Theoretical Biology, 8(1):71– 80, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900937.

Kimelberg:1965:MIH

[KMT65] H. Kimelberg, J. F. Moran, and D. J. Triggle. The mechanism of interaction of 2-halogenoethylamines at the noradrenaline recentor. Journal of Theoretical Biology, 9(3):502–503, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900470.

Knowles:1965:ESC

[Kno65] J. R. Knowles. Enzyme specificity: α-chymotrypsin. Journal of Theoretical Biology, 9(2):213–228, September 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365901086.

Knox:1968:TTE

[Kno68] R. S. Knox. On the theory of trapping of excitation in the photosynthetic unit. Journal of Theoretical Biology, 21(2):244– 259, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900738.

Koch:1962:ERB

[Koc62] Arthur L. Koch. The evaluation of the rates of biological processes from tracer kinetic data. I. The influence of labile metabolic pools. Journal of Theoretical Biology, 3(2):283–303, September 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519362800186.

Koch:1964:GVP

[Koc64] Arthur L. Koch. The growth of viral plaques during the enlargement phase. Journal of Theoretical Biology, 6(3):413–431, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900566.

Koch:1966:LBM

[Koc66] Arthur L. Koch. The logarithm in biology 1. Mechanisms generating the log-normal distribution exactly. Journal of Theoretical Biology, 12(2):276–290, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901196.

Koch:1967:KPC

[Koc67a] Arthur L. Koch. Kinetics of permease catalyzed transport. Journal of Theoretical Biology, 14(2):103–130, February 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367901099.

Koch:1967:MCT

[Koc67b] Arthur L. Koch. Metabolic control through reflexive enzyme action. Journal of Theoretical Biology, 15(1):75–102, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900458.

Koch:1967:TBN

[Koc67c] Arthur L. Koch. The tight binding nature of the repressoroperon interaction. Journal of Theoretical Biology, 16(2):166– 186, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900033.

Koch:1968:ERB

[Koc68a] Arthur L. Koch. The evaluation of the rates of biological processes from tracer kinetic data: II. RNA metabolism in growing bacteria. Journal of Theoretical Biology, 18(1):105–132, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901732.

Koch:1968:TAD

[Koc68b] Arthur L. Koch. Theory of the angular dependence of light scattered by bacteria and similar-sized biological objects. Journal of Theoretical Biology, 18(1):133–156, January 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368901744.

Koch:1969:LBI

[Koc69] A. L. Koch. The logarithm in biology: II. Distributions simulating the log-normal. Journal of Theoretical Biology, 23(2): 251–268, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936990040X.

Koide:1966:DIB

[Koi66] F. T. Koide. The determination of the intercellular bridge resistance between smooth muscle cells. Journal of Theoretical Biology, 12(1):89–118, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901871.

Koide:1967:DIB

[Koi67] Frank T. Koide. Determination of intercellular bridge resistance in smooth muscle by measurement of injury and sucrose-gap potentials. Journal of Theoretical Biology, 16(2):268–279, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900082.

Komarov:1963:SPA

[Kom63] L. V. Komarov. Some potential areas for research on means of drastically increasing the human life span. Journal of Theoretical Biology, 4(2):147–158, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900237.

Kordan:1967:LIF

[Kor67] Herbert A. Kordan. Localized interfacial forces resulting from implanted plastics as possible physical factors involved in tumor formation. Journal of Theoretical Biology, 17(1):1–11, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900148.

Korn:1969:SAD

[Kor69] Robert W. Korn. A stochastic approach to the development of coleocheate. Journal of Theoretical Biology, 24(2):147–158, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800421.

Koshland:1962:CNE

[Kos62] D. E. Koshland. The comparison of non-enzymic and enzymic reaction velocities. Journal of Theoretical Biology, 2(2):75–86, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519362900371.

King:1967:DMS

[KP67] Charles E. King and G. J. Paulik. Dynamic models and the simulation of ecological systems. Journal of Theoretical Biology, 16(2): 251–267, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900070.

Krzywicki:1967:FAP

[KS67] André Krzywicki and Piotr P. Slonimski. Formal analysis of protein sequences: I. Specific long-range constraints in pair associations of amino acids. Journal of Theoretical Biology, 17(1):136– 158, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900252.

Kruckenberg:1968:AMA

[KS68a] Peter Kruckenberg and Rainer Sandweg. An analog model for acetylcholine release by motor nerve endings. Journal of Theoretical Biology, 19(3):327–332, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901446.

Krzywicki:1968:FAP

[KS68b] A. Krzywicki and P. P. Slonimski. Formal analysis of protein sequences: II. Method for structural studies of homologous proteins amino acid substitutions in cytochromes c. Journal of Theoretical Biology, 21(3):305–330, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901185.

Kirschbaum:1969:MMP

[KS69] Thomas H. Kirschbaum and Norman Z. Shapiro. A mathematical model of placental oxygen transfer. Journal of Theoretical Biology, 25(3):380–402, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800287.

Kustin:1965:TRS

[KSK65] Kenneth Kustin, David Shear, and Daniel Kleitman. Theory of relaxation spectra; the kinetics of coupled chemical reactions. Journal of Theoretical Biology, 9(2):186–212, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365901074.

Kimelberg:1965:RHN

[KT65] H. Kimelberg and D. J. Triggle. The reactions of 2-halogenoethylamines at the noradrenaline receptor-storage complex. Journal of Theoretical Biology, 9(2):313–322, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365901165.

Kruger-Thiemer:1966:FTD

[KT66] Ekkehard Krüger-Thiemer. Formal theory of drug dosage regimens. I. Journal of Theoretical Biology, 13(??):212–235, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936690018X.

Kruger-Thiemer:1969:FTD

[KT69] E. Krüger-Thiemer. Formal theory of drug dosage regimens: II. The exact plateau effect. Journal of Theoretical Biology, 23(2): 169–190, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900356.

Ladik:1964:PIM

[Lad64] J. Ladik. Possible interpretation of the mutagenic effect of ultraviolet radiation. Journal of Theoretical Biology, 6(2):201– 207, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900281.

Lanni:1963:ASP

[Lan63] Frank Lanni. Analysis of sequence patterns in ribonuclease. III. Variable-span pair-order analysis. Journal of Theoretical Biology, 4 (1):1–27, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900973.

Lauria:1965:MAS

[Lau65] Francesco E. Lauria. Mathematical approach to the study of a cerebral cortex: I. Histological methods of calculating the coupling coefficients in the neuronic equations. Journal of Theoretical Biology, 8(1):54–70, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900925.

Lauria:1969:MASa

[Lau69a] F. E. Lauria. Mathematical approach to the study of a cerebral cortex: II. The self-consistency of the model and the coupling coefficients of the neuronic equations. Journal of Theoretical Biology, 23 (1):72–86, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936990068X.

Lauria:1969:MASb

[Lau69b] F. E. Lauria. Mathematical approach to the study of a cerebral cortex: III. The columnar organization in the cortex and the spread of activity in the model. Journal of Theoretical Biology, 24(3):307– 316, September 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251936980055X.

Lea:1963:PTL

[Lea63] E. J. A. Lea. Permeation through long narrow pores. Journal of Theoretical Biology, 5(1):102–107, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900389.

Lea:1965:PTL

[Lea65] E. J. A. Lea. Permeation through long narrow pores — corrections. Journal of Theoretical Biology, 9(2):350, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365901190.

Leibovic:1966:MIP

[Lei66] K. N. Leibovic. A model for information processing with reference to vision. Journal of Theoretical Biology, 11(1):112–130, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900427.

Leibowitz:1968:FF

[Lei68] Lila Leibowitz. Founding families. Journal of Theoretical Biology, 21(2):153–169, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900684.

Leibovic:1969:SPI

[Lei69a] K. N. Leibovic. Some problems of information processing and models of the visual pathway. Journal of Theoretical Biology, 22(1): 62–79, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900800.

Leibowitz:1969:DSE

[Lei69b] Lila Leibowitz. Dilemma for social evolution: The impact of Darwin. Journal of Theoretical Biology, 25(2):255–275, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800639.

Lengerova:1962:SII

[Len62] A. Lengerová. The significance of immunological immaturity for the induction of tolerance. Journal of Theoretical Biology, 3(3):503–508, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519362800411.

Lesk:1969:WDD

[Les69] Arthur M. Lesk. Why does DNA contain thymine and RNA uracil? Journal of Theoretical Biology, 22(3):537–540, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900216.

Levitt:1962:SDH

[Lev62] J. Levitt. A sulfhydryl-disulfide hypothesis of frost injury and resistance in plants. Journal of Theoretical Biology, 3(3):355– 391, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800320.

Levins:1964:TFH

[Lev64] Richard Levins. Theory of fitness in a heterogeneous environment: Part III. The response to selection. Journal of Theoretical Biology, 7(2):224–240, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900712.

Levine:1966:MRE

[Lev66] Sumner N. Levine. A model for renal-electrolyte regulation. Journal of Theoretical Biology, 11(2):242–256, July 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366901640.

Lewontin:1961:ETG

[Lew61] R. C. Lewontin. Evolution and the theory of games. Journal of Theoretical Biology, 1(3):382–403, July 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900388.

Lewis:1962:GSI

[Lew62] D. Lewis. The generation of self-incompatibility alleles. Journal of Theoretical Biology, 2(1):69–71, January 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800289.

Lewis:1966:NPD

[Lew66] E. R. Lewis. Neuroelectric potentials derived from an extended version of the Hodgkin–Huxley model. Journal of Theoretical Biology, 10(1):125–152, IN1–IN3, 153–158, January 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901810.

Lewin:1967:SAH

[Lew67] S. Lewin. Some aspects of hydration and stability of the native state of DNA. Journal of Theoretical Biology, 17(2):181– 188, IN1–IN3, 189–212, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936790166X.

Lewin:1969:ILP

[Lew69] S. Lewin. Ionic linkages in protein interactions. Journal of Theoretical Biology, 23(2):279–280, IN1–IN2, 281–284, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900423.

Levy:1968:SMB

[LG68] P. S. Levy and G. M. Green. A stochastic model of the bactericidal activity of the lung. Journal of Theoretical Biology, 21(1): 103–112, October 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900635.

Lindegren:1961:BFD

[Lin61] Carl C. Lindegren. The biological function of deoxyribonucleic acid. Journal of Theoretical Biology, 1(2):107–119, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519361900418.

Lindegren:1963:RHI

[Lin63] Carl C. Lindegren. The receptor-hypothesis of induction of gene-controlled adaptive enzymes. Journal of Theoretical Biology, 5(2):192–210, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900596.

Lindley:1967:MSB

[Lin67] Barry D. Lindley. Membrane solvation as a basis for ionic selectivity. Journal of Theoretical Biology, 17(2):213–228, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901671.

Lindenmayer:1968:MMCa

[Lin68a] Aristid Lindenmayer. Mathematical models for cellular interactions in development I. Filaments with one-sided inputs. Journal of Theoretical Biology, 18(3):280–299, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900799.

Lindenmayer:1968:MMCb

[Lin68b] Aristid Lindenmayer. Mathematical models for cellular interactions in development II. Simple and branching filaments with two-sided inputs. Journal of Theoretical Biology, 18(3):300– 315, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900805.

Lindley:1968:MST

[Lin68c] B. D. Lindley. Membrane systems as transition assemblies — a basis for phenomenological descriptions. Journal of Theoretical Biology, 20(1):56–64, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936890091X.

Lwoff:1962:ECM

[LL62] André Lwoff and Marguerite Lwoff. Événements cycliques et Molécules Métastables. Journal of Theoretical Biology, 2(1):48– 62, January 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800265.

Levine:1969:KIU

[LL69] M. Levine and S. Levine. Kinetics of induced uphill transport of sugars in human erythrocytes. Journal of Theoretical Biology, 24 (1):85–107, July 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800081.

Longmuir:1965:TAO

[LM65] I. S. Longmuir and M. McCabe. Tissue adaptation to oxygen lack. Journal of Theoretical Biology, 8(1):124–129, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900986.

Lifson:1966:TUT

[LM66] N. Lifson and Ruth McClintock. Theory of use of the turnover rates of body water for measuring energy and material balance. Journal of Theoretical Biology, 12(1):46–74, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901858.

Latimer:1968:CTL

[LMB68] Paul Latimer, Douglas M. Moore, and F. Dudley Bryant. Changes in total light scattering and absorption caused by changes in particle conformation. Journal of Theoretical Biology, 21(3):348– 367, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901203.

Lockhart:1965:AIP

[Loc65] James A. Lockhart. An analysis of irreversible plant cell elongation. Journal of Theoretical Biology, 8(2):264–275, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900779.

Loewy:1968:TCB

[Loe68] Ariel G. Loewy. A theory of covalent bonding in muscle contraction. Journal of Theoretical Biology, 20(2):164–172, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901872.

Lovtrup:1963:RWE

[Løv63] Søren Løvtrup. On the rate of water exchange across the surface of animal cells. Journal of Theoretical Biology, 5(3):341– 359, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900821.

Lovrien:1969:ADE

[Lov69] Rex Lovrien. Activation diagrams in enzyme catalysis. Journal of Theoretical Biology, 24(3):247–265, September 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800500.

Lowenstein:1961:ROM

[Low61] John M. Lowenstein. Reductions and oxidations in mammalian biosyntheses. Journal of Theoretical Biology, 1(1):98–103, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519361900297.

Lowenhaupt:1969:BTM

[Low69] B. Lowenhaupt. Bioelectrochemical throttle of metabolism. Journal of Theoretical Biology, 25(2):187–202, IN1–IN2, 203–207, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800585.

Levey:1966:NTU

[LP66] H. C. Levey and C. J. Perret. A note on the theory of uncatalysed linear expanding systems. Journal of Theoretical Biology, 13(??): 15–17, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900038.

Lebowitz:1969:GCD

[LR69] J. L. Lebowitz and S. I. Rubinow. Grain count distributions in labeled cell populations. Journal of Theoretical Biology, 23(1): 99–123, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900708.

Lucy:1964:GLM

[Luc64] J. A. Lucy. Globular lipid micelles and cell membranes. Journal of Theoretical Biology, 7(2):360–373, September 1964. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519364900803.

Lurie:1969:DST

[Lur69] M. Lurie. The Darwinian selection theory of antibody formation. Journal of Theoretical Biology, 23(3):380–386, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900265.

MacGillivray:1969:ASP

[Mac69] A. D. MacGillivray. Asymptotic solution of the Poisson–Boltzmann equation for a charged sphere, and implications. Journal of Theoretical Biology, 23(2):205–217, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936990037X.

Maddaiah:1969:MHS

[Mad69] V. T. Maddaiah. A model for hormone specific activation of adenyl cyclase. Journal of Theoretical Biology, 25(3):495–498, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800366.

Maelzer:1965:DCE

[Mae65a] D. A. Maelzer. A discussion of components of environment in ecology. Journal of Theoretical Biology, 8(1):141–162, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365901001.

Maelzer:1965:ESS

[Mae65b] D. A. Maelzer. Environment, semantics, and system theory in ecology. Journal of Theoretical Biology, 8(3):395–402, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900184.

Magar:1969:AAS

[Mag69a] M. E. Magar. Analysis of aggregating systems. Journal of Theoretical Biology, 22(2):291–313, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900071.

Magar:1969:DRC

[Mag69b] Magar E. Magar. On the determination of rate constants and the number of intermediates in enzyme-catalyzed reactions. Journal of Theoretical Biology, 25(3):343–346, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251936980024X.

Maldonado:1963:VAL

[Mal63] H. Maldonado. The visual attack learning system in Octopus vulgaris. Journal of Theoretical Biology, 5(3):470–488, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900900.

Malmon:1965:HRI

[Mal65] A. G. Malmon. High resolution isotope tracing in electron microscopy using induced nuclear reactions. Journal of Theoretical Biology, 9(1):77–92, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900585.

Maniloff:1969:CLR

[Man69a] Jack Maniloff. Collision lifetimes and recognition times for macromolecule synthesis. Journal of Theoretical Biology, 25(2):339– 342, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800688.

Maniloff:1969:TCB

[Man69b] Jack Maniloff. Theoretical considerations of biopolymer synthesis. Journal of Theoretical Biology, 23(3):441–454, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900307.

Marler:1961:LAA

[Mar61] Peter Marler. The logical analysis of animal communication. Journal of Theoretical Biology, 1(3):295–317, July 1961. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519361900327.

Martin:1963:BPR

[Mar63] H. H. Martin. Bacterial protoplasts — a review. Journal of Theoretical Biology, 5(1):1–16, IN1–IN4, 17–34, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900341.

Marshall:1964:TAE

[Mar64] John H. Marshall. Theory of alkaline earth metabolism: The power function makes possible a simple but comprehensive model of skeletal systems. Journal of Theoretical Biology, 6(3):386– 412, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900554.

Marshak:1967:HOT

[Mar67] A. Marshak. The harmonic oscillator theory applied to the code translation problem. Journal of Theoretical Biology, 17(1):12– 18, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936790015X.

Matsumoto:1968:NSR

[Mat68] Yorimi Matsumoto. Normalization of shortening in rana pipiens sartorius muscle. Journal of Theoretical Biology, 18(3):387– 395, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900866.

McMullen:1963:PDH

[McM63] A. I. McMullen. The polynucleotide double helix as a versatile assembly template during polymer biogenesis. Journal of Theoretical Biology, 5(1):127–134, IN5, 135–136, IN7–IN10, 137– 141, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900419.

Mel:1964:SFFa

[Mel64a] Howard C. Mel. Stable-flow free boundary migration and fractionation of cell mixtures: I. Apparatus and hydrodynamic feedback principles. Journal of Theoretical Biology, 6(2):159–166, IN1, 167–180, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900268.

Mel:1964:SFFb

[Mel64b] Howard C. Mel. Stable-flow free boundary migration and fractionation of cell mixtures: II. Laminar flow and density gradient stability. Journal of Theoretical Biology, 6(2):181–188, IN3, 189– 200, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936490027X.

Mel:1964:SFFc

[Mel64c] Howard C. Mel. Stable-flow free boundary (STAFLO) migration and fractionation of cell mixtures: III. Migration principles — sedimentation and electrophoresis. Journal of Theoretical Biology, 6(3):307–322, IN1, 323–324, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900505.

MacGillivray:1969:AGC

[MH69] A. D. MacGillivray and Daphne Hare. Applicability of Goldman’s constant field assumption to biological systems. Journal of Theoretical Biology, 25(1):113–126, October 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800196.

Morowitz:1964:PSM

[MHMQ64] H. J. Morowitz, W. A. Higinbotham, S. W. Matthysse, and H. Quastler. Passive stability in a metabolic network. Journal of Theoretical Biology, 7(1):98–111, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900438.

Milne:1962:TNC

[Mil62] A. Milne. On a theory of natural control of insect population. Journal of Theoretical Biology, 3(1):19–50, July 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800034.

Miles:1966:MWM

[Mil66] P. W. Miles. A modification of Wigglesworth’s model for the excretion of uric acid in insects, in the light of modern hypotheses of ion transport. Journal of Theoretical Biology, 12(1):130– 132, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901895.

Moses:1966:SMC

[MLH66] V. Moses and K. K. Lonberg-Holm. The study of metabolic compartmentalization. Journal of Theoretical Biology, 10(2):336– 355, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901317.

Macgillivray:1966:IIS

[MM66a] A. D. Macgillivray and A. I. McMullen. On the influence of ionic strength on the melting temperature of DNA. Journal of Theoretical Biology, 12(2):260–265, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901172.

MacGillivray:1966:TSR

[MM66b] A. D. MacGillivray and A. I. McMullen. A theoretical study of the role of salt concentration in the melting of DNA. Journal of Theoretical Biology, 12(1):75–88, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936690186X.

MacGillivray:1968:THC

[MM68a] A. D. MacGillivray and A. I. McMullen. Theory of helix-coil transitions based on the non-linear Poisson–Boltzmann equation. Journal of Theoretical Biology, 19(2):159–168, May 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368901112.

Moore:1968:LSC

[MM68b] G. E. Moore and W. F. McLimans. The life span of the cultured normal cell: Concepts derived from studies of human lymphoblasts. Journal of Theoretical Biology, 20(2):217–226, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901914.

Morowitz:1966:PBC

[Mor66] Harold J. Morowitz. Physical background of cycles in biological systems. Journal of Theoretical Biology, 13(??):60–62, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900075.

Morowitz:1969:MAF

[Mor69] Harold J. Morowitz. A mechanism for the amplification of fluctuations in racemic mixtures. Journal of Theoretical Biology, 25(3):491–494, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519369800354.

Miller:1963:TPM

[MP63] D. S. Miller and P. R. Payne. A theory of protein metabolism. Journal of Theoretical Biology, 5(3):398–411, November 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519363900869.

Mueller:1963:IER

[MR63] Paul Mueller and Donald O. Rudin. Induced excitability in reconstituted cell membrane structure. Journal of Theoretical Biology, 4 (3):268–280, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900067.

Mueller:1968:RAP

[MR68] Paul Mueller and Donald O. Rudin. Resting and action potentials in experimental bimolecular lipid membranes. Journal of Theoretical Biology, 18(2):222–224, IN1, 225–258, February 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936890163X.

Murakami:1966:EAM

[Mur66] Hideo Murakami. Electron aspects of the mode of action of the mitomycin molecule. Journal of Theoretical Biology, 10(2):236– 250, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901251.

Myhill:1969:IED

[Myh69] J. Myhill. Investigation of the effect of data error in the analysis of biological tracer data from three compartment systems. Journal of Theoretical Biology, 23(2):218–231, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900381.

Naora:1968:PMS

[Nao68] Hiroto Naora. A plausible mechanism of simultaneous read-out of genetic information from both strands of DNA. Journal of Theoretical Biology, 19(2):183–195, May 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901136.

Nash:1969:SRT

[Nas69] H. A. Nash. The stereoisomers of retinal — a theoretical study of energy differences. Journal of Theoretical Biology, 22(2):314– 324, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900083.

Nelson:1962:KES

[Nel62] Eino Nelson. Kinetics of the excretion of sulfonamides during therapeutic dosage regimens. Journal of Theoretical Biology, 2(3): 193–200, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519362900243.

Nelson:1963:DPS

[Nel63] Eino Nelson. Design of pharmacogenetic studies of drug metabolism. Journal of Theoretical Biology, 5(3):493–496, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900924.

Nicklas:1966:NOM

[Nic66] R. Bruce Nicklas. A note on orientation in mitosis and meiosis. Journal of Theoretical Biology, 12(1):147–150, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901925.

Nagata:1965:QMS

[NIT^{+}65] C. Nagata, A. Imamura, Y. Tagashira, M. Kodama, and N. Fukuda. Quantum-mechanical study on the photodimerization of aromatic molecules. Journal of Theoretical Biology, 9(3):357– 365, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900378.

Nagata:1968:MMA

[NM68] C. Nagata and O. Mårtensson. On the mechanism of mutagenic action of hydroxylamine. Journal of Theoretical Biology, 19(2): 133–146, May 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901094.

Nakahama:1966:SDB

[NN66] Hiroshi Nakahama and Shinko Nishioka. Statistical dependency between intervals in neuronal impulse sequences. Journal of Theoretical Biology, 12(1):140–146, September 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901913.

Norton:1966:ISM

[NO66] Dorita A. Norton and Jean M. Ohrt. Introduction to steroid molecular stacking. Journal of Theoretical Biology, 11(1):140– 149, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900440.

Noack:1968:RMS

[Noa68] D. Noack. A regulatory model for steady-state conditions in populations of lysogenic bacteria. Journal of Theoretical Biology, 18(1): 1–8, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901665.

Nobel:1969:BVH

[Nob69] Park S. Nobel. The Boyle–Van’t Hoff relation. Journal of Theoretical Biology, 23(3):375–379, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900253.

Nooney:1965:MMR

[Noo65] Grove C. Nooney. Mathematical models, reality and results. Journal of Theoretical Biology, 9(2):239–252, September 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365901104.

Nooney:1966:AET

[Noo66] Grove C. Nooney. Allocation of errors to theory and experiment. Journal of Theoretical Biology, 11(1):46–53, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900373.

Nooney:1968:ADS

[Noo68] G. C. Nooney. Age distributions in stochastically dividing populations. Journal of Theoretical Biology, 20(3):314–320, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901318.

Norton:1968:DNB

[Nor68] Stata Norton. On the discontinuous nature of behavior. Journal of Theoretical Biology, 21(2):229–243, November 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368900726.

Naftalin:1961:ECR

[NS61] L. Naftalin and A. Stephens. The effect of concentration and rate of intravenous injection of thiopentone. Journal of Theoretical Biology, 1(3):407–409, July 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900406.

OBrien:1967:CMC

[O’B67] John S. O’Brien. Cell membranes — composition: Structure: Function. Journal of Theoretical Biology, 15(3):307–312, IN3, 313–324, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901403.

Ohki:1967:IIL

[OF67] Shinpei Ohki and Nobuo Fukuda. Interlayer-interaction for a lipid bilayer model. Journal of Theoretical Biology, 15(3):362– 375, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901439.

Ohki:1967:ALG

[Ohk67] Shinpei Ohki. Application of the lattice gas theory to lipid monolayers. Journal of Theoretical Biology, 15(3):346–361, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901427.

Ohki:1968:DCR

[Ohk68] Shinpei Ohki. Dielectric constant and refractive index of lipid bilayers. Journal of Theoretical Biology, 19(1):97–115, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900088.

Ohki:1969:PLB

[Ohk69] Shinpei Ohki. Properties of lipid bilayer membranes: Determination of membrane thickness. Journal of Theoretical Biology, 23(1): 158–168, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900733.

ONeill:1969:IEE

[O’N69] Robert V. O’Neill. Indirect estimation of energy fluxes in animal food webs. Journal of Theoretical Biology, 22(2):284–290, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936990006X.

Ordon:1969:PDS

[Ord69] Richard Ordon. Polyribosome dynamics at steady state. Journal of Theoretical Biology, 22(3):515–532, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900186.

Orloci:1968:IAP

[Orl68] Laszlo Orloci. Information analysis in phytosociology: Partition, classification and prediction. Journal of Theoretical Biology, 20(3):271–284, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901288.

OSullivan:1962:QPE

[O’S62] D. G. O’Sullivan. Quantitative potentialities in enzyme cytochemistry. Modified Michaelis–Menten rate law applicable when a substrate diffuses slowly into an enzyme site. Journal of Theoretical Biology, 2(2):117–128, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519362900401.

Ozonoff:1969:ADV

[Ozo69] David Ozonoff. An attack on “a defense of vitalism”. Journal of Theoretical Biology, 24(1):121–122, July 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251936980010X.

Palmiter:1963:AAR

[PA63] Michael T. Palmiter and Frederick Aladjem. The antigen-antibody reaction. IV. a quantitative theory of antigen-antibody reactions. Journal of Theoretical Biology, 5(2):211–235, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900602.

Palmiter:1968:CIA

[PA68] Michael T. Palmiter and Frederick Aladjem. On the composition of insoluble antigen-antibody complexes. Journal of Theoretical Biology, 18(1):34–52, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901690.

Paigen:1962:RDT

[Pai62] Kenneth Paigen. On the regulation of DNA transcription. Journal of Theoretical Biology, 3(2):268–282, September 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800174.

Parsons:1964:IWB

[Par64] P. A. Parsons. Interactions within and between chromosomes. Journal of Theoretical Biology, 6(2):208–216, March 1964. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519364900293.

Partridge:1966:PSN

[Par66] Lloyd D. Partridge. A possible source of nerve signal distortion arising in pulse rate encoding of signals. Journal of Theoretical Biology, 11(2):257–281, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901652.

Parsegian:1967:ELR

[Par67] V. A. Parsegian. Evidence for long-range repulsion between bimolecular leaflets of lecithin: An examination of structural data. Journal of Theoretical Biology, 15(1):70–74, April 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367900446.

Partridge:1968:PSN

[Par68] Lloyd D. Partridge. A possible source of nerve signal distortion arising in pulse rate encoding of signals. Journal of Theoretical Biology, 21(2):292, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900763.

Parry:1969:ACC

[Par69] David A. D. Parry. An alternative to the coiled-coil for αfibrous proteins. Journal of Theoretical Biology, 24(1):73–84, July 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251936980007X.

Passow:1967:SSD

[Pas67] H. Passow. Steady-state diffusion of non-electrolytes through epithelial brush borders. Journal of Theoretical Biology, 17(3):383– 384, IN2, 385–398, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900999.

Pattee:1967:QMH

[Pat67] H. H. Pattee. Quantum mechanics, heredity and the origin of life. Journal of Theoretical Biology, 17(3):410–420, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901014.

Pavlidis:1969:PIO

[Pav69] T. Pavlidis. Populations of interacting oscillators and circadian rhythms. Journal of Theoretical Biology, 22(3):418–436, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900149.

Pennycuick:1968:CMS

[PCB68] C. J. Pennycuick, R. M. Compton, and Linda Beckingham. A computer model for simulating the growth of a population, or of two interacting populations. Journal of Theoretical Biology, 18(3): 316–329, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900817.

Perl:1969:IET

[PEC69] W. Perl, R. M. Effros, and F. P. Chinard. Indicator equivalence theorem for input rates and regional masses in multiinlet steady-state systems with partially labeled input. Journal of Theoretical Biology, 25(2):297–316, November 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800652.

Pella:1969:SMP

[Pel69] Jerome J. Pella. A stochastic model for purse seining in a two-species fishery. Journal of Theoretical Biology, 22(2):209– 226, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900010.

Pennycuick:1969:CMO

[Pen69] Linda Pennycuick. A computer model of the Oxford great tit population. Journal of Theoretical Biology, 22(3):381–400, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900113.

Perault:1962:SEA

[Per62a] Anne-Marie Perault. Some electronic aspects of the mechanism of action of clostridial xanthine dehydrogenase. Journal of Theoretical Biology, 2(3):263–265, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519362900292.

Perl:1962:HMD

[Per62b] W. Perl. Heat and matter distribution in body tissues and the determination of tissue blood flow by local clearance methods. Journal of Theoretical Biology, 2(3):201–235, May 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519362900255.

Pettersson:1964:CMT

[Pet64] Max Pettersson. On the convenience of more thorough metric usage in the biological and allied sciences. Journal of Theoretical Biology, 6(2):217–243, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936490030X.

Patton:1967:RPM

[PF67] Stuart Patton and Frederick M. Fowkes. The role of the plasma membrane in the secretion of milk fat. Journal of Theoretical Biology, 15(3):274–276, IN1, 277–281, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901373.

Patlak:1963:FSS

[PGH63] C. S. Patlak, D. A. Goldstein, and J. F. Hoffman. The flow of solute and solvent across a two-membrane system. Journal of Theoretical Biology, 5(3):426–442, November 1963. CODEN JTBIAP.

ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900882.

Perl:1966:LBF

[PH66] W. Perl and R. L. Hirsch. Local blood flow in kidney tissue by heat clearance measurement. Journal of Theoretical Biology, 10(2):251– 280, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901263.

Phillips:1969:UMD

[Phi69] A. P. Phillips. Unwinding models for double-helical DNA during its replication and transcription. Journal of Theoretical Biology, 24(3):273–278, September 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800524.

Pickard:1966:PNI

[Pic66] William F. Pickard. On the propagation of the nervous impulse down medullated and unmedullated fibers. Journal of Theoretical Biology, 11(1):30–45, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900361.

Pielou:1966:MDD

[Pie66a] E. C. Pielou. The measurement of diversity in different types of biological collections. Journal of Theoretical Biology, 13(??):131– 144, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900130.

Pielou:1966:SDP

[Pie66b] E. C. Pielou. Species-diversity and pattern-diversity in the study of ecological succession. Journal of Theoretical Biology, 10(2):370– 383, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901330.

Pielou:1967:MDD

[Pie67] E. C. Pielou. The measurement of diversity in different types of biological collections. Journal of Theoretical Biology, 15(1): 177, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900483.

Packard:1968:IFA

[PK68] G. C. Packard and D. L. Kilgore. Internal fertilization: Adaptive value to early amniotes. Journal of Theoretical Biology, 20(2):245– 248, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901938.

Parker:1967:PDK

[PL67] Rodger D. Parker and Thomas L. Lincoln. Parameters deducible from the kinetics of enzyme induction and repression. Journal of Theoretical Biology, 15(2):218–224, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367902020.

Platt:1961:PLM

[Pla61] John R. Platt. Properties of large molecules that go beyond the properties of their chemical sub-groups. Journal of Theoretical Biology, 1(3):342–358, July 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900364.

Pullman:1969:CTI

[PLB69] Bernard Pullman, Jacqueline Langlet, and Hélène Berthod. On the S-cis-S-trans isomerism of the cyclohexene ring with respect to the polyene chain in retinals, carotenoids, vitamins a and related compounds. Journal of Theoretical Biology, 23(3):492– 495, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900344.

Pollard:1961:PAP

[Pol61] Ernest Pollard. Physical aspects of protein and DNA synthesis. Journal of Theoretical Biology, 1(3):328–341, July 1961. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519361900352.

Pollard:1963:CKA

[Pol63] Ernest Pollard. Collision kinetics applied to phage synthesis, messenger RNA, and glucose metabolism. Journal of Theoretical Biology, 4(1):98–112, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363901036.

Pollard:1965:TSL

[Pol65] Ernest C. Pollard. Theoretical studies on living systems in the absence of mechanical stress. Journal of Theoretical Biology, 8(1): 113–123, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900974.

Pielou:1967:DDD

[PP67] D. P. Pielou and E. C. Pielou. The detection of different degrees of coexistence. Journal of Theoretical Biology, 16(3):427– 437, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900653.

Pielou:1968:AAS

[PP68] D. P. Pielou and E. C. Pielou. Association among species of infrequent occurrence: The insect and spider fauna of polyporus betulinus (bulliard) fries. Journal of Theoretical Biology, 21(2):202– 216, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900702.

Pollak:1966:HBS

[PR66] Michael Pollak and Robert Rein. On the hydrogen bonding scheme between purines and pyrimidines. Journal of Theoretical Biology, 11(3):490–494, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901093.

Pollak:1968:EMS

[PR68a] M. Pollak and Robert Rein. An electrical mechanism for strand separation of DNA. Journal of Theoretical Biology, 19(3):333– 336, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901458.

Pollak:1968:EEC

[PR68b] Michael Pollak and Robert Rein. An examination of the energetics of Crick’s wobble hypothesis. Journal of Theoretical Biology, 19(3):

241–246, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901379.

Pratt:1962:FN

[Pra62] J. M. Pratt. The fixation of nitrogen. Journal of Theoretical Biology, 2(3):251–258, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519362900279.

Pring:1967:SADa

[Pri67a] M. Pring. The simulation and analysis by digital computer of biochemical systems in terms of kinetic models: I. The choice of integration method. Journal of Theoretical Biology, 17(3):421– 429, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901026.

Pring:1967:SADb

[Pri67b] M. Pring. The simulation and analysis by digital computer of biochemical systems in terms of kinetic models: II. Curvefitting procedures. Journal of Theoretical Biology, 17(3):430– 435, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901038.

Pring:1967:SADc

[Pri67c] M. Pring. The simulation and analysis by digital computer of biochemical systems in terms of kinetic models: III. Generator programming. Journal of Theoretical Biology, 17(3):436–440, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936790104X.

Pandharipande:1968:RRK

[PS68] K. R. Pandharipande and U. N. Singh. Ribosomal RNA: Kinetic analysis of incorporation of labelled precursors. Journal of Theoretical Biology, 20(3):285–296, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936890129X.

Pscheidt:1963:BCP

[Psc63] Gordon R. Pscheidt. Biochemical correlates with phyletic division of the nervous system. Journal of Theoretical Biology, 5(1): 52–56, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900365.

Pullman:1962:PCR

[Pul62] Alberte Pullman. On the possibilities of correlating the redox potential of reversible systems to the molecular orbitals involved in the electron transfer. Journal of Theoretical Biology, 2(3): 259–262, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519362900280.

Perault:1961:EAM

[PVP61] A-M. Perault, C. Valdemoro, and B. Pullman. Electronic aspects of the mechanism of action of xanthine oxidase. Journal of Theoretical Biology, 1(2):180–189, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900443.

Pavlidis:1968:MMT

[PZO68] T. Pavlidis, W. F. Zimmerman, and J. Osborn. A mathematical model for the temperature effects on circadian rhythms. Journal of Theoretical Biology, 18(2):210–221, February 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368901628.

Quastler:1962:RPC

[QZ62] H. Quastler and G. Zubay. An RNA-protein code based on replacement data. II. Adjustment and extension. Journal of Theoretical Biology, 3(3):496–502, November 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251936280040X.

Rappaport:1965:GRC

[Rap65] R. Rappaport. Geometrical relations of the cleavage stimulus in invertebrate eggs. Journal of Theoretical Biology, 9(1): 51–66, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900561.

Rapoport:1968:ATI

[Rap68] Stanley I. Rapoport. Active transport in isotropic and anisotropic systems: a critical analysis. Journal of Theoretical Biology, 19(3): 247–274, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901380.

Rice:1963:PSD

[RB63] William E. Rice and Robert M. Bock. The problem of sequence determination in transfer RNA. Journal of Theoretical Biology, 4(3): 260–267, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900055.

Ralph:1966:RPB

[RB66] R. K. Ralph and B. C. Baguley. Regulation of protein biosynthesis. Journal of Theoretical Biology, 13(??):164–186, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900154.

Raff:1968:PRA

[RB68a] Elizabeth C. Raff and J. J. Blum. A possible role for adenylate kinase in cilia: Concentration profiles in a geometrically constrained dual enzyme system. Journal of Theoretical Biology, 18(1):53– 71, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901707.

Rouslin:1968:AFA

[RB68b] William Rouslin and Emory Braswell. Analysis of factors affecting lactic dehydrogenase subunit composition determinations. Journal of Theoretical Biology, 19(2):169–182, May 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901124.

Rinaldi:1969:SFM

[RB69] R. A. Rinaldi and W. R. Baker. A sliding filament model of amoeboid motion. Journal of Theoretical Biology, 23(3):463– 474, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900320.

Rohlf:1969:SSM

[RD69] F. James Rohlf and Demorest Davenport. Simulation of simple models of animal behavior with a digital computer. Journal of Theoretical Biology, 23(3):400–424, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900289.

Rehm:1968:ASC

[Reh68] W. S. Rehm. An analysis of the short-circuiting technique applied to in vivo tissues. Journal of Theoretical Biology, 20(3):341– 354, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901343.

Rendel:1962:RBG

[Ren62] J. M. Rendel. The relationship between gene and phenotype. Journal of Theoretical Biology, 2(3):296–308, May 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519362900322.

Ressler:1969:CCP

[Res69] Newton Ressler. Control of cellular processes by the coupling of resonant energy to hydrogen transfer reactions I. Application to enzyme mechanisms. Journal of Theoretical Biology, 23(3): 425–440, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900290.

Rein:1968:IEH

[RFCH68] Robert Rein, Nobuo Fukuda, George A. Clarke, and Frank E. Harris. Iterative extended Hückel study of nucleic acid bases. Journal of Theoretical Biology, 21(1):88–96, October 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368900611.

Rose:1969:OCS

[RG69] S. Rose and P. Glow. Ontogeny of chimaerism and its sequelae. Journal of Theoretical Biology, 22(1):43–61, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900794.

Redwood:1969:ITM

[RH69] W. R. Redwood and D. A. Haydon. Influence of temperature and membrane composition on the water permeability of lipid bilayers. Journal of Theoretical Biology, 22(1):1–8, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900757.

Rickards:1965:ACO

[Ric65] G. K. Rickards. An analysis of co-orientation in mitosis and meiosis. Journal of Theoretical Biology, 9(2):332–333, IN3, 334–349, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365901189.

Riggs:1966:QHC

[Rig66] Douglas S. Riggs. A quantitative hypothesis concerning the action of the parathyroid hormone. Journal of Theoretical Biology, 12(3):364–372, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901494.

Rinaldini:1965:AMD

[Rin65] L. M. Rinaldini. Antigenesis and molecular differentiation. Journal of Theoretical Biology, 9(2):125–155, September 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365901050.

Robson:1967:BMR

[RKB67] D. S. Robson, R. F. Kahrs, and J. A. Baker. Bounds on the mean recurrence time of subclinical epidemics in dairy herds. Journal of Theoretical Biology, 17(1):47–56, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900197.

Rider:1968:MPC

[RM68] K. Rider and H. J. Morowitz. The most probable covalent bond distribution in non-equilibrium systems of an atomic composition characteristic of the biosphere. Journal of Theoretical Biology, 21(2):278–291, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900751.

Roberts:1965:SRP

[Rob65] Richard B. Roberts. The synthesis of ribosomal protein. Journal of Theoretical Biology, 8(1):49–53, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900913.

Robinson:1968:RIP

[Rob68] Joseph D. Robinson. Regulating ion pumps to control cell volume. Journal of Theoretical Biology, 19(1):90–96, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900076.

Rosenberg:1964:PJF

[Ros64] Jerome L. Rosenberg. Professor James Franck. Journal of Theoretical Biology, 7(2):389, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900839.

Rosen:1967:TFM

[Ros67] Robert Rosen. Two-factor models, neural nets and biochemical automata. Journal of Theoretical Biology, 15(3):282–297, June 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901385.

Rosen:1968:SCP

[Ros68] Robert Rosen. Some comments on the physico-chemical description of biological activity. Journal of Theoretical Biology, 18(3): 380–386, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368900854.

Roth:1963:HCP

[Rot63] Jay S. Roth. A hypothesis concerning patterns of deoxynucleotide metabolism in tumors in relation to rate of cell proliferation. Journal of Theoretical Biology, 4(1):113–123, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363901048.

Rhoads:1968:SAD

[RP68] D. G. Rhoads and M. Pring. The simulation and analysis by digital computer of biochemical systems in terms of kinetic models IV. automatic derivation of enzymic rate laws. Journal of Theoretical Biology, 20(3):297–313, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901306.

Rummel:1961:EEH

[RPS61] W. Rummel, K. Pfleger, and E. Seifen. Experimental examination of the hypothesis of coupling between the sodium and potassium pump in erythrocytes. Journal of Theoretical Biology, 1(3): 318–323, July 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900339.

Rosado:1966:GCS

[RR66] J. M. Campos Rosado and Alan Robertson. The genetic control of sex ratio. Journal of Theoretical Biology, 13(??):324– 329, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900257.

Reanney:1967:SOG

[RR67] D. C. Reanney and R. K. Ralph. A speculation on the origin of the genetic code. Journal of Theoretical Biology, 15(1): 41–52, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900422.

Reanney:1968:GCE

[RR68] D. C. Reanney and R. K. Ralph. Genetic circularity and evolution. Journal of Theoretical Biology, 21(2):217–228, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900714.

Rescigno:1962:AMB

[RS62] A. Rescigno and G. Segre. Analysis of multicompartmented biological systems. Journal of Theoretical Biology, 3(2):149–163, September 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800125.

Ruckenstein:1966:RSL

[RS66] E. Ruckenstein and Z. Simon. Regulation and synthesis in the living cell I. Kinetics of ribonucleic acid synthesis. Journal of Theoretical Biology, 11(2):282–298, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901664.

Renneboog-Squilbin:1967:TCS

[RS67] C. Renneboog-Squilbin. Theory of the chemostat — study of the conditions under which a population of micro-organisms and its mutant will coexist in the chemostat. Journal of Theoretical Biology, 14(1):74–101, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900951.

Reed:1967:SBE

[RTB67] Jon Reed, Robert Toombs, and Nils Aall Barricelli. Simulation of biological evolution and machine learning: I. Selection of selfreproducing numeric patterns by data processing machines, effects of hereditary control, mutation type and crossing. Journal of Theoretical Biology, 17(3):319–342, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900975.

Rubin:1967:OCG

[Rub67] Jerrold Rubin. Optimal classification into groups: An approach for solving the taxonomy problem. Journal of Theoretical Biology, 15(1):103–144, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936790046X.

Russell:1968:MMO

[Rus68a] C. S. Russell. Mechanisms as models for oxidative phosphorylation. Journal of Theoretical Biology, 19(3):337–339, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936890146X.

Russell:1968:ABM

[Rus68b] Charles D. Russell. Acid-base metabolism and the proton condition. Journal of Theoretical Biology, 19(2):196–214, May 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901148.

Rosenberg:1963:CTU

[RW63] T. Rosenberg and W. Wilbrandt. Carrier transport uphill. I. General. Journal of Theoretical Biology, 5(2):288–305, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519363900651.

Strickland:1966:SAS

[SA66] E. Hardin Strickland and Eugene Ackerman. Simulation of the action of stimulatory agents on mitochondrial respiration. Journal of Theoretical Biology, 10(1):114–124, January 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366901809.

Swanson:1966:BGO

[SAFT66] C. H. Swanson, R. Aris, A. G. Fredrickson, and H. M. Tsuchiya. Bacterial growth as an optimal process. Journal of Theoretical Biology, 12(2):228–250, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901159.

Sagan:1967:OMC

[Sag67] Lynn Sagan. On the origin of mitosing cells. Journal of Theoretical Biology, 14(3):225–274, IN1–IN6, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900793.

Saidel:1968:BCPb

[Sai68a] G. M. Saidel. Bacterial cell populations in a continuously changing environment. Journal of Theoretical Biology, 21(2):292, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900775.

Saidel:1968:BCPa

[Sai68b] Gerald M. Saidel. Bacterial cell populations in a continuously changing environment. Journal of Theoretical Biology, 19(3):287– 296, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901409.

Samis:1966:CBA

[Sam66] H. V. Samis. A concept of biological aging: The role of compensatory processes. Journal of Theoretical Biology, 13(??):236– 250, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900191.

Saroff:1965:SRD

[Sar65] H. A. Saroff. The structure of ribonuclease derived from the clustering of its ionizable groups. Journal of Theoretical Biology, 9(2):229– 238, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365901098.

Sattler:1967:PIP

[Sat67] R. Sattler. Petal inception and the problem of pattern detection. Journal of Theoretical Biology, 17(1):31–39, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900173.

Savageau:1969:BSAa

[Sav69a] Michael A. Savageau. Biochemical systems analysis: I. Some mathematical properties of the rate law for the component enzymatic reactions. Journal of Theoretical Biology, 25(3):365–369, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800263.

Savageau:1969:BSAb

[Sav69b] Michael A. Savageau. Biochemical systems analysis: II. The steady-state solutions for an n-pool system using a power-law approximation. Journal of Theoretical Biology, 25(3):370–379, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800275.

Shea:1965:NSC

[SB65] Stephen M. Shea and Anthony F. Bartholomay. In numero studies in a cell renewal system periodically adjusted cell renewal process. Journal of Theoretical Biology, 9(3):389–413, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900391.

Schwimmer:1962:TDP

[Sch62] Sigmund Schwimmer. Theory of double pH optima of enzymes. Journal of Theoretical Biology, 3(1):102–110, July 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251936280006X.

Scheuplein:1968:ART

[Sch68] Robert J. Scheuplein. On the application of rate theory to complex multibarrier flow co-ordinates: Membrane permeability. Journal of Theoretical Biology, 18(1):72–89, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901719.

Segal:1967:ECI

[Seg67] John R. Segal. Electrical capacitance of ion-exchanger membranes. Journal of Theoretical Biology, 14(1):11–34, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900902.

Segal:1969:IPB

[Seg69] John R. Segal. The “independence principle” of biological membranes: Its misuse. Journal of Theoretical Biology, 24(2):159– 170, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800433.

Semenza:1967:RES

[Sem67] G. Semenza. Rate equations of some cases of enzyme inhibition and activation — their application to sodium-activated membrane transport systems. Journal of Theoretical Biology, 15(1): 145–176, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900471.

Serra:1964:GCT

[Ser64] J. A. Serra. The genetic concept of treption (and brief reference to a unifying theory of cancer). Journal of Theoretical Biology, 6(3): 371–374, May 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900530.

Sugita:1963:FACb

[SF63] Motoyosi Sugita and Nobuo Fukuda. Functional analysis of chemical systems in vivo using a logical circuit equivalent: III. Analysis using a digital circuit combined with an analogue computer. Journal of Theoretical Biology, 5(3):412–418, IN1, 419– 425, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900870.

Szent-Gyorgyi:1961:SSB

[SG61] Albert Szent-Györgyi. The supra- and submolecular in biology. Journal of Theoretical Biology, 1(1):75–82, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519361900273.

Stahl:1963:MA

[SG63] Walter R. Stahl and Harry E. Goheen. Molecular algorithms. Journal of Theoretical Biology, 5(2):266–287, September 1963. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251936390064X.

Sibatani:1964:MCD

[SH64] A. Sibatani and S. Hiai. A model of chromosome duplication in Escherichia coli. Journal of Theoretical Biology, 7(3):393– 411, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936490013X.

Shapiro:1961:RBI

[Sha61] Herbert Shapiro. On the relation between ionization and biological activity in some sympathomimetic amines. Journal of Theoretical Biology, 1(3):289–294, July 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900315.

Shaffer:1965:MCM

[Sha65] B. M. Shaffer. Mechanical control of the manufacture and resorption of cell surface in collective amoebae. Journal of Theoretical Biology, 8(1):27–40, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900895.

Shea:1962:SAH

[She62a] Stephen M. Shea. The statistical approach in histology. Journal of Theoretical Biology, 3(1):111–122, July 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800071.

Sheppard:1962:SMT

[She62b] C. W. Sheppard. Stochastic models for tracer experiments in the circulation: Parallel random walks. Journal of Theoretical Biology, 2(1):33–47, January 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519362800253.

Shear:1967:ABH

[She67] David Shear. An analog of the Boltzmann h-theorem (a Liapunov function) for systems of coupled chemical reactions. Journal of Theoretical Biology, 16(2):212–228, August 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367900057.

Shikata:1963:RCS

[Shi63] Morikazu Shikata. Representation and calculation of selfed population by group ring. Journal of Theoretical Biology, 5(1):142– 160, July 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900420.

Shikata:1964:IEC

[Shi64] Morikazu Shikata. Interference effect of cross-overs in selfed populations. Journal of Theoretical Biology, 7(2):181–223, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900700.

Shikata:1966:CIC

[Shi66a] Morikazu Shikata. Calculations of inbreeding coefficient and geneset probability in diploid populations. Journal of Theoretical Biology, 10(1):15–27, January 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901767.

Shikata:1966:CEH

[Shi66b] Morikazu Shikata. Cross-over effect in homozygosity by descent. Journal of Theoretical Biology, 10(2):196–208, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901238.

Shikata:1966:RRC

[Shi66c] Morikazu Shikata. Rod and ring chromosome cross-overs described by braid group. Journal of Theoretical Biology, 10(2):184– 195, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901226.

Shikata:1966:TGI

[Shi66d] Morikazu Shikata. Transformation of generalized inbreeding coefficient in components of pedigree. Journal of Theoretical Biology, 10(1):11–14, January 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901755.

Shikata:1967:CIC

[Shi67a] Morikazu Shikata. Change of inbreeding coefficient in mixed selfing and random mating populations. Journal of Theoretical Biology, 17(2):229–239, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367901683.

Shikata:1967:ESU

[Shi67b] Morikazu Shikata. Effect of selections upon recombination probabilities in selfed populations. Journal of Theoretical Biology, 14(1): 59–65, January 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900938.

Shikata:1967:GIC

[Shi67c] Morikazu Shikata. Generalized inbreeding coefficient in seifing obtained from group king. Journal of Theoretical Biology, 17(3):447– 450, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901063.

Shikata:1967:VHD

[Shi67d] Morikazu Shikata. Variance of homozygosity by descent in selfed populations. Journal of Theoretical Biology, 17(2):240– 245, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367901695.

Shikata:1969:TMIa

[Shi69a] Morikazu Shikata. Transformation of mictic index through a diploid individual. Journal of Theoretical Biology, 24(2):131– 136, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800408.

Shikata:1969:TMIb

[Shi69b] Morikazu Shikata. Transformation of mictic indices in components of pedigree. Journal of Theoretical Biology, 24(3):279–293, September 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800536.

Sibatani:1964:MDS

[Sib64] Atuhiro Sibatani. The mechanism of DNA synthesis by DNA polymerase. Journal of Theoretical Biology, 7(3):558–561, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900219.

Siegel:1969:TAT

[Sie69] Israel Siegel. Theoretical aspects of termination of immune tolerance by cross-reacting antigens. Journal of Theoretical Biology, 24(2):171–191, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519369800445.

Simon:1965:LMH

[Sim65a] Z. Simon. Lethal mutation hypothesis for the mechanism of action of the cytostatic alkylating agents. Journal of Theoretical Biology, 8(1):193–197, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365901025.

Simon:1965:RBB

[Sim65b] Z. Simon. Recognition between biological molecules. Journal of Theoretical Biology, 9(3):414–421, November 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365900408.

Simon:1965:MSS

[Sim65c] Zeno Simon. Multi-steady-state model for cell differentiation. Journal of Theoretical Biology, 8(2):258–263, March 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365900767.

Simon:1967:RSP

[Sim67] Z. Simon. Regulation and synthesis processes in the living cell IV. mitotic and functional homeostasis by competition between derepressed operons. Journal of Theoretical Biology, 16(2):294– 305, August 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900100.

Singh:1965:RTR

[Sin65] U. N. Singh. Rapidly turning over RNA as a limiting factor in protein synthesis: Effect of actinomycin D. Journal of Theoretical Biology, 8(3):480–489, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900238.

Sink:1967:TAS

[Sin67] J. D. Sink. Theoretical aspects of sex odor in swine. Journal of Theoretical Biology, 17(1):174–180, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900276.

Singh:1969:PUM

[Sin69] U. N. Singh. Polyribosomes and unstable messenger RNA: a stochastic model of protein synthesis. Journal of Theoretical Biology, 25(3):444–460, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800329.

Shea:1969:VTA

[SKB69] Stephen M. Shea, Morris J. Karnovsky, and William H. Bossert. Vesicular transport across endothelium: Simulation of a diffusion model. Journal of Theoretical Biology, 24(1):30–32, IN1, 33–42, July 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800044.

Smith:1966:AAC

[Smi66] Malcolm H. Smith. The amino acid composition of proteins. Journal of Theoretical Biology, 13(??):261–282, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936690021X.

Smith:1967:CTH

[Smi67] C. Lavett Smith. Contribution to a theory of hermaphroditism. Journal of Theoretical Biology, 17(1):76–90, October 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367900215.

Snell:1965:FTO

[Sne65] Fred M. Snell. Facilitated transport of oxygen through solutions of hemoglobin. Journal of Theoretical Biology, 8(3):469– 479, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900226.

Sneath:1966:RBC

[Sne66] P. H. A. Sneath. Relations between chemical structure and biological activity in peptides. Journal of Theoretical Biology, 12(2):157– 195, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901123.

Sokal:1962:TET

[Sok62] Robert R. Sokal. Typology and empiricism in taxonomy. Journal of Theoretical Biology, 3(2):230–267, September 1962. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800162.

Soloway:1966:AHN

[Sol66a] A. H. Soloway. Aromatic hydroxylation — a new biosynthetic mechanism. Journal of Theoretical Biology, 13(??):100–105, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900117.

Soloway:1966:SLP

[Sol66b] A. H. Soloway. Squalene to lanosterol-possible biosynthetic pathway. Journal of Theoretical Biology, 13(??):429–432, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900300.

Savage:1966:SDP

[SP66] John R. K. Savage and D. G. Papworth. The site distribution parameter for radiation induced chromosome exchanges. Journal of Theoretical Biology, 11(1):1–9, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900336.

Savage:1969:DHA

[SP69] J. R. K. Savage and D. G. Papworth. Distortion hypothesis: An alternative to a limited number of sites for radiation-induced chromosome exchange. Journal of Theoretical Biology, 22(3):493– 514, March 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900174.

Spetner:1964:NSI

[Spe64] L. M. Spetner. Natural selection: An information-transmission mechanism for evolution. Journal of Theoretical Biology, 7(3):412– 429, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900141.

Spencer:1967:BVH

[Spe67] Richard P. Spencer. A blood volume heart weight relationship. Journal of Theoretical Biology, 17(3):441–446, December 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901051.

Sprent:1967:EMG

[Spr67] P. Sprent. Estimation of mean growth curves for groups of organisms. Journal of Theoretical Biology, 17(1):159–173, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900264.

Simon:1966:RSPa

[SR66a] Z. Simon and E. Ruckenstein. Regulation and synthesis processes in the living cell II. Kinetics of protein synthesis. Journal of Theoretical Biology, 11(2):299–313, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901676.

Simon:1966:RSPb

[SR66b] Z. Simon and E. Ruckenstein. Regulation and synthesis processes in the living cell III. Interrelated operon triggers as elements of the cellular automaton. Journal of Theoretical Biology, 11(2): 314–333, July 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901688.

Satir:1964:MNS

[SS64] P. Satir and B. Satir. A model for ninefold symmetry in α keratin and cilia. Journal of Theoretical Biology, 7(1):123–124, IN1, 125–128, July 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900451.

Snell:1966:ORR

[SS66] Fred M. Snell and Barry Stein. Onsager reciprocal relations in complex membranes. Journal of Theoretical Biology, 10(2):177– 183, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901214.

Spangler:1967:TFAb

[SS67a] Robert A. Spangler and Fred M. Snell. Transfer function analysis of an oscillatory model chemical system. Journal of Theoretical Biology, 16(3):381–405, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936790063X.

Spangler:1967:TFAa

[SS67b] Robert A. Spangler and Fred M. Snell. Transfer function analysis of chemical kinetic systems. Journal of Theoretical Biology, 16(3):366–380, September 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900628.

Sendov:1968:CSR

[ST68] Bl. Sendov and R. Tsanev. Computer simulation of the regenerative processes in the liver. Journal of Theoretical Biology, 18(1): 90–104, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901720.

Stahl:1965:AUP

[Sta65] Walter R. Stahl. Algorithmically unsolvable problems for a cell automaton. Journal of Theoretical Biology, 8(3):371–394, May 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900172.

Stanley:1966:MTG

[Sta66] John Stanley. A mathematical theory of the growth of populations of the flour beetle Tribolium confusum, Duval IX. The persistence of tunnels in the flour mass. Journal of Theoretical Biology, 13(??):379–411, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900282.

Stahl:1967:CMC

[Sta67] Walter R. Stahl. A computer model of cellular self-reproduction. Journal of Theoretical Biology, 14(2):187–205, February 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367901130.

Stern:1964:QPB

[Ste64] Alexander W. Stern. Quantum physics and biological systems. Journal of Theoretical Biology, 7(2):318–328, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936490075X.

Stewart:1968:MAE

[Ste68] Robert M. Stewart. Model analysis of electrical conduction through morphologically complex cells and cell masses. Journal of Theoretical Biology, 19(3):311–318, IN1, 319, June 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901422.

Stibitz:1968:MCP

[Sti68] George R. Stibitz. A model and computer program for biological networks. Journal of Theoretical Biology, 19(1):116– 132, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936890009X.

Stonehill:1965:MDR

[Sto65] E. H. Stonehill. On the mechanism of DNA replication. Journal of Theoretical Biology, 9(2):323–328, IN1, 329–331, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365901177.

Strauss:1962:SAM

[Str62] Bernard S. Strauss. Some aspects of mutation. Journal of Theoretical Biology, 2(3):266–278, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519362900309.

Straile:1969:VCO

[Str69] William E. Straile. Vertical cutaneous organization. Journal of Theoretical Biology, 24(2):203–204, IN1–IN3, 205–215, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800469.

Sheppard:1969:SMT

[SU69] C. W. Sheppard and M. B. Uffer. Stochastic models for tracer experiments in the circulation: II. Serial random walks. Journal of Theoretical Biology, 22(1):188–207, January 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519369900873.

Subirana:1968:RSM

[Sub68] Juan A. Subirana. Role of spindle microtabules in mitosis. Journal of Theoretical Biology, 20(1):117–123, July 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368900970.

Sugita:1963:FACa

[Sug63] Motoyosi Sugita. Functional analysis of chemical systems in vivo using a logical circuit equivalent. II. The idea of a molecular automaton. Journal of Theoretical Biology, 4(2):179–184, IN2, 185– 192, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900274.

Sugita:1966:FAC

[Sug66] Motoyosi Sugita. Functional analysis of chemical systems in vivo using a logical circuit equivalent IV. simulation of cellular control systems using a hybrid computing system. Journal of Theoretical Biology, 13(??):330–356, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900269.

Smith:1966:CBQ

[SW66] H. J. Smith and H. Williams. A comment on the binding of quaternary ammonium inhibitors at the anionic site of acety1cholinesterase. Journal of Theoretical Biology, 11(1):177– 179, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900488.

Smith:1967:PTA

[SW67] H. J. Smith and H. Williams. A partition theory for the action of quaternary ammonium compounds at the muscarinic receptor site. Journal of Theoretical Biology, 14(2):218–224, February 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901154.

Swade:1969:CRF

[Swa69] Richard H. Swade. Circadian rhythms in fluctuating light cycles: Toward a new model of entrainment. Journal of Theoretical Biology, 24(2):227–239, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800482.

Sykes:1966:MRR

[Syk66] J. Sykes. A mechanism for the regulation of ribosomal ribonucleic acid synthesis in bacteria. Journal of Theoretical Biology, 12(3):373–384, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901500.

Simha:1962:SDN

[SZ62] R. Simha and J. M. Zimmerman. Sequence distribution and neighbor effects of various orders in polynucleotides. Journal of Theoretical Biology, 2(2):87–104, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519362900383.

Simha:1965:CCD

[SZ65] R. Simha and J. M. Zimmerman. Calculation of chainlength distribution between “markers” in mixed polynucleotides. Journal of Theoretical Biology, 8(1):81–86, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900949.

Taborsky:1967:APT

[Tab67] Robert G. Taborsky. Atherosclerosis: a physical-thermodynamic hypothesis for its genesis. Journal of Theoretical Biology, 16(1): 147–150, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900586.

Takahashi:1966:TBC

[Tak66] Manabu Takahashi. Theoretical basis for cell cycle analysis I. Labelled mitosis wave method. Journal of Theoretical Biology, 13(??):202–211, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900178.

Takahashi:1968:TBC

[Tak68] Manabu Takahashi. Theoretical basis for cell cycle analysis: II. Further studies on labelled mitosis wave method. Journal of Theoretical Biology, 18(2):195–209, February 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368901616.

Trams:1966:MLS

[TB66] Eberhard G. Trams and Elise Ann Brown. Models for lipoprotein synthesis. Journal of Theoretical Biology, 12(3):311–326, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901457.

Trucco:1968:PLM

[TB68] E. Trucco and P. J. Brockwell. Percentage labeled mitoses curves in exponentially growing cell populations. Journal of Theoretical Biology, 20(3):321–337, September 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936890132X.

Terzi:1965:RGC

[Ter65] Mario Terzi. Radiosensitivity and genetic complexity. Journal of Theoretical Biology, 8(2):233–243, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900743.

Thoma:1968:PMA

[Tho68] John A. Thoma. A possible mechanism for amylase catalysis. Journal of Theoretical Biology, 19(3):297–310, June 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368901410.

Takahashi:1967:CCR

[TI67] Manabu Takahashi and Kazuo Inouye. Characteristics of cube root growth of transplantable tumours. Journal of Theoretical Biology, 14(3):275–283, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251936790080X.

Tien:1967:BLM

[Tie67] H. T. Tien. Black lipid membranes: Thickness determination and molecular organization by optical methods. Journal of Theoretical Biology, 16(1):97–110, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900550.

Tallis:1966:SAS

[TL66] G. M. Tallis and Morley Leyton. A stochastic approach to the study of parasite populations. Journal of Theoretical Biology, 13(??):251– 260, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366900208.

Tenforde:1968:MDF

[TM68] Tom Tenforde and Robert Macey. Membrane diffusion with feedback regulation of cell volume; estimation of membrane tortuosity and diffusion coefficients. Journal of Theoretical Biology, 19(2): 232–240, May 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901161.

Tollin:1962:PPP

[Tol62] Gordon Tollin. The primary process of photosynthesis: an interpretation of the Emerson effect and of the light-induced spectral changes at 7,000 å. Journal of Theoretical Biology, 2(2):105– 116, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519362900395.

Tolansky:1963:IBM

[Tol63] S. Tolansky. Interferometry and biological microtopography. Journal of Theoretical Biology, 5(2):306–312, IN1–IN14, 313–320, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900663.

Tomita:1966:MCR

[Tom66a] Tadao Tomita. Membrane capacity and resistance of mammalian smooth muscle. Journal of Theoretical Biology, 12(2):216–227, November 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901147.

Tomlinson:1966:AHP

[Tom66b] J. Tomlinson. The advantages of hermaphroditism and parthenogenesis. Journal of Theoretical Biology, 11(1):54–58, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366900385.

Talmage:1963:ARM

[TP63] David W. Talmage and David S. Pearlman. The antibody response: a model based on antagonistic actions of antigen. Journal of Theoretical Biology, 5(2):321–339, September 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900675.

Thiebaux:1967:SSP

[TP67] H. Jean Thiebaux and H. H. Pattee. Statistical studies of protein sequences: Pair and triplet analysis in helical conformations. Journal of Theoretical Biology, 17(1):121–135, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900240.

TenBosch:1963:RET

[TR63] J. J. Ten Bosch and Th. W. Ruijgrok. A remark on the energy transfer in biological systems. Journal of Theoretical Biology, 4(2): 225–227, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900316.

TenHoopen:1965:FCP

[TR65] M. Ten Hoopen and H. A. Reuver. An n-fold coincidence problem in physiology. Journal of Theoretical Biology, 9(1):117– 123, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900603.

Triggle:1964:RMS

[Tri64] D. J. Triggle. The reaction mechanisms, selectivity and specificity of action of 2-halogenoethylamines — an analysis. Journal of Theoretical Biology, 7(2):241–275, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900724.

Triggle:1969:CGF

[Tri69] D. J. Triggle. A comment on “a generalized functional role for phospholipids”. Journal of Theoretical Biology, 25(3):499–501, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800378.

Trucco:1967:CFN

[Tru67] E. Trucco. Collection functions for non-equivivant cell populations. Journal of Theoretical Biology, 15(2):180–189, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901993.

Tsanev:1966:MRM

[TS66] R. Tsanev and Bl. Sendov. A model of the regulatory mechanism of cellular multiplication. Journal of Theoretical Biology, 12(3):327– 341, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901469.

Tsanev:1969:MCS

[TS69] R. Tsanev and Bl. Sendov. A model of cancer studied by a computer. Journal of Theoretical Biology, 23(1):124–134, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936990071X.

Tyler:1966:MIP

[Tyl66] Albert Tyler. A model illustrating possible instability of cellular structure under conditions of weightlessness. Journal of Theoretical Biology, 11(1):59–62, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900397.

Urry:1965:BETa

[UE65a] Dan W. Urry and Henry Eyring. Biological electron transport: I. Imidazole pump model of electron transport: a group transfer model. Journal of Theoretical Biology, 8(1):198–213, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365901037.

Urry:1965:BETb

[UE65b] Dan W. Urry and Henry Eyring. Biological electron transport: II. A variation of the imidazole pump model to include coupling. Journal of Theoretical Biology, 8(1):214–220, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365901049.

Ullrick:1967:TCS

[Ull67] William C. Ullrick. A theory of contraction for striated muscle. Journal of Theoretical Biology, 15(1):53–69, April 1967. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519367900434.

Urbakh:1968:RDR

[Urb68] V. Yu Urbakh. Reversible denaturation of RNA and memory. Journal of Theoretical Biology, 18(2):259–261, February 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901641.

Vaidhyanathan:1964:PET

[Vai64] V. S. Vaidhyanathan. On the permeability of electrolytes through biological membranes. Journal of Theoretical Biology, 7(2):334– 338, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519364900773.

Vaidhyanathan:1965:CAT

[Vai65a] V. S. Vaidhyanathan. On the criteria of active transport. Journal of Theoretical Biology, 9(3):489–501, November 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365900469.

Vaidhyanathan:1965:SMT

[Vai65b] V. S. Vaidhyanathan. Statistical mechanical theory of electrolyte diffusion inside a charged membrane. Journal of Theoretical Biology, 8(2):344–356, March 1965. CODEN JTBIAP.

ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900822.

Vaidhyanathan:1965:STT

[Vai65c] V. S. Vaidhyanathan. Statistical theory of transport of electrolytes across a charged membrane. Journal of Theoretical Biology, 9(3):478–488, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900457.

Vaidhyanathan:1966:SCS

[Vai66a] V. S. Vaidhyanathan. Some comments on steady state transport of ions across a charged biological membrane. Journal of Theoretical Biology, 10(2):159–176, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901202.

Vaidhyanathan:1966:TME

[Vai66b] V. S. Vaidhyanathan. Theory of mechano-electric membrane transducers. Journal of Theoretical Biology, 13(??):18–31, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936690004X.

Vaidhyanathan:1969:CFE

[Vai69] V. S. Vaidhyanathan. Conductance function of electrolyte transport across lipid films. Journal of Theoretical Biology, 23(2): 232–250, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900393.

VanLiew:1967:GAA

[Van67] Hugh D. Van Liew. Graphic analysis of aggregates of linear and exponential processes. Journal of Theoretical Biology, 16(1): 43–53, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900525.

Vayo:1965:DEP

[Vay65] H. Westcott Vayo. Determination of the electrical parameters of vertebrate visceral smooth muscle. Journal of Theoretical Biology, 9(2):263–277, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365901128.

vanderSteen:1969:CVP

[vdS69] W. J. van der Steen. A comment on vitalistic principles. Journal of Theoretical Biology, 24(1):123–125, July 1969. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519369800111.

Vaidhyanathan:1968:SLF

[VG68] V. S. Vaidhyanathan and Narendra S. Goel. Stability of lipid films in aqueous electrolyte media: Electrostatic interactions. Journal of Theoretical Biology, 21(3):331–347, December 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901197.

Vidaver:1966:IPF

[Vid66] George A. Vidaver. Inhibition of parallel flux and augmentation of counter flux shown by transport models not involving a mobile carrier. Journal of Theoretical Biology, 10(2):301– 306, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519366901287.

Vogel:1969:SDS

[VNM69] H. Vogel, H. Niewisch, and G. Matioli. Stochastic development of stem cells. Journal of Theoretical Biology, 22(2):249– 270, February 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900046.

vanOss:1963:AAP

[vO63] C. J. van Oss. Antigen-antibody precipitates considered as specifically impermeable barriers. Journal of Theoretical Biology, 4(2): 223–224, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900304.

Vaidhyanathan:1964:PNE

[VP64] V. S. Vaidhyanathan and W. H. Perkins. On the permeability of non-electrolytes through biological membranes. Journal of Theoretical Biology, 7(2):329–333, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900761.

Veillard:1965:ASE

[VP65a] Alain Veillard and Bernard Pullman. Aspects de la structure électronique de la vitamine B_{12} et de ses analogues. Journal of Theoretical Biology, 8(2):307–316, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900792.

Veillard:1965:CSE

[VP65b] Alain Veillard and Bernard Pullman. Complements sur la structure electronique des ferroporphyrines biologiques. Journal of Theoretical Biology, 8(2):317–326, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900809.

Vaidhyanathan:1966:MII

[VP66a] V. S. Vaidhyanathan and Hannah M. Phillips. Mobile ion interactions in the membrane phase. Journal of Theoretical Biology, 13(??):32–47, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900051.

Vaidhyanathan:1966:MTN

[VP66b] V. S. Vaidhyanathan and Hannah M. Phillips. Molecular theory of nerve potentials. Journal of Theoretical Biology, 10(3): 460–489, April 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936690138X.

Vaidhyanathan:1964:TTA

[VPT64] V. S. Vaidhyanathan, W. H. Perkins, and E. J. Towbin. Theory of transport across a membrane model. Journal of Theoretical Biology, 7(2):339–351, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900785.

Vander:1961:TMA

[VWM61] Arthur J. Vander, Walter S. Wilde, and Richard L. Malvin. A theoretical mode of action of aldosterone. Journal of Theoretical Biology, 1(2):236–243, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900479.

Wong:1965:CHA

[WA65] James Wong and Frederick Aladjem. Competitive hapten-antibody binding: Solution of the binding equation. Journal of Theoretical Biology, 8(1):41–48, January 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365900901.

Waddington:1965:ACP

[Wad65] C. H. Waddington. Autogenous cellular periodicities as (a) temporal templates and (b) the basis of “morphogenetic fields”. Journal of Theoretical Biology, 8(2):367–369, March 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365900846.

Wagner:1968:KPR

[Wag68] J. G. Wagner. Kinetics of pharmacologic response I. Proposed relationships between response and drug concentration in the intact animal and man. Journal of Theoretical Biology, 20(2):173– 201, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901884.

Wallace:1965:RHI

[Wal65] Margaret E. Wallace. The relative homozygosity of inbred lines and closed colonies. Journal of Theoretical Biology, 9(1):93– 116, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900597.

Walser:1966:MARb

[Wal66a] Mackenzie Walser. Mathematical aspects of renal function: Reabsorption of individual solutes as interdependent processes. Journal of Theoretical Biology, 10(2):327–335, February 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366901305.

Walser:1966:MARa

[Wal66b] Mackenzie Walser. Mathematical aspects of renal function: The dependence of solute reabsorption on water reabsorption, and the mechanism of osmotic natriuresis. Journal of Theoretical Biology, 10(2):307–326, February 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366901299.

Walter:1966:QSS

[Wal66c] Charles Walter. Quasi-steady state in a general enzyme system. Journal of Theoretical Biology, 11(2):181–206, July 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366901603.

Walter:1967:PUS

[Wal67] Charles Walter. The practicality of the use of the steadystate assumption and the inflection point method in enzyme kinetics. Journal of Theoretical Biology, 15(1):1–33, April 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900409.

Walter:1969:ASC

[Wal69a] Charles Walter. The absolute stability of certain types of controlled biological systems. Journal of Theoretical Biology, 23(1): 39–52, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900666.

Walter:1969:SCB

[Wal69b] Charles Walter. Stability of controlled biological systems. Journal of Theoretical Biology, 23(1):23–38, April 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900654.

Wang:1963:CRA

[Wan63] Jui H. Wang. Chemical reaction accelerated transport of molecules through membranes. Journal of Theoretical Biology, 4(2):175– 178, March 1963. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519363900262.

Warner:1964:MMI

[War64] Donald T. Warner. Molecular models: IV. a suggested conformation for the protein subunit of tobacco mosaic virus. Journal of Theoretical Biology, 6(1):118–136, January 1964. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519364900682.

Warburton:1967:MNS

[War67a] Frederick E. Warburton. A model of natural selection based on a theory of guessing games. Journal of Theoretical Biology, 16 (1):78–96, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900549.

Ware:1967:SBC

[War67b] G. C. Ware. The simulation of bacterial cultures on a digital computer. Journal of Theoretical Biology, 17(1):91–107, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900227.

Watkins:1965:PRG

[Wat65] J. C. Watkins. Pharmacological receptors and general permeability phenomena of cell membranes. Journal of Theoretical Biology, 9(1):37–50, July 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936590055X.

Waddington:1969:CSM

[WC69] C. H. Waddington and Russell J. Cowe. Computer simulation of a mulluscan pigmentation pattern. Journal of Theoretical Biology, 25(2):219–220, IN3–IN4, 221–225, November 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800603.

Weiss:1962:CMC

[Wei62] Leonard Weiss. Cell movement and cell surfaces: a working hypothesis. Journal of Theoretical Biology, 2(3):236–250, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519362900267.

Weiss:1963:CU

[Wei63] Paul Weiss. The cell as unit. Journal of Theoretical Biology, 5(3):389–397, November 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900857.

Weiss:1964:CLP

[Wei64] Leonard Weiss. Cellular locomotive pressure in relation to initial cell contacts. Journal of Theoretical Biology, 6(2):275–281, March 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900335.

Weiss:1968:SCD

[Wei68] Leonard Weiss. Studies on cell deformability: V. Some effects of ribonuclease. Journal of Theoretical Biology, 18(1):9– 18, January 1968. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519368901677.

Wenner:1961:DLH

[Wen61] Adrian M. Wenner. Division of labor in a honey bee colony — a Markov process? Journal of Theoretical Biology, 1(3): 324–327, July 1961. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519361900340.

Werman:1965:SMP

[Wer65] R. Werman. The specificity of molecular processes involved in neural transmission. Journal of Theoretical Biology, 9(3):471– 477, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900445.

Wever:1966:DRE

[Wev66] R. Wever. The duration of re-entrainment of circadian rhythms after phase shifts of the zeitgeber a theoretical investigation. Journal of Theoretical Biology, 13(??):187–201, December 1966. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519366900166.

Woolf:1966:LRT

[WGP66] L. I. Woolf, B. L. Goodwin, and C. E. Phelps. T_{m}-limited renal tubular reabsorption and the genetics of renal glucosuria. Journal of Theoretical Biology, 11(1):10–21, May 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519366900348.

Whitehead:1965:TQS

[Whi65] E. P. Whitehead. A theory of the quaternary structure of dehydrogenases, dehydrogenating complexes, and other proteins. Journal of Theoretical Biology, 8(2):276–306, March 1965. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519365900780.

Whyte:1967:DAO

[Why67] Lancelot Law Whyte. Directive agencies in organic evolution. Journal of Theoretical Biology, 17(2):312–314, November 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367901749.

Williams:1961:PFC

[Wil61a] R. J. P. Williams. Possible functions of chains of catalysts. Journal of Theoretical Biology, 1(1):1–17, January 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900236.

Willmer:1961:HCV

[Wil61b] E. N. Willmer. Human colour vision and the perception of blue. Journal of Theoretical Biology, 1(2):141–179, April 1961. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519361900431.

Williams:1962:PFC

[Wil62] R. J. P. Williams. Possible functions of chains of catalysts II. Journal of Theoretical Biology, 3(2):209–229, September 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519362800150.

Wilkins:1964:TPM

[Wil64] B. R. Wilkins. A theory of position memory. Journal of Theoretical Biology, 7(2):374–387, September 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900815.

Wilson:1966:CSM

[Wil66] Theodore A. Wilson. Capillary spacing for minimum entropy production. Journal of Theoretical Biology, 11(3):436–445, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366901032.

Williams:1967:MCG

[Wil67] F. M. Williams. A model of cell growth dynamics. Journal of Theoretical Biology, 15(2):190–207, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367902007.

Williams:1969:TME

[Wil69] Richard B. Williams. A table of mean effective temperatures for the metabolism of biological systems subjected to sinusoidal cycles in temperature. Journal of Theoretical Biology, 24(2):240– 245, August 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800494.

Winkler:1963:SSA

[Win63] Marvin H. Winkler. On the structure of a specific antibody site. Journal of Theoretical Biology, 4(3):237–241, May 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251936390002X.

Winfree:1967:BRB

[Win67] Arthur T. Winfree. Biological rhythms and the behavior of populations of coupled oscillators. Journal of Theoretical Biology, 16 (1):15–42, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900513.

Winne:1967:FKR

[WO67] D. Winne and H. Ochsenfahrt. Die formale kinetik der resorption unter berücksichtigung der darmdurchblutung. (German) [Formal kinetics of absorption taking into account digestive circulation]. Journal of Theoretical Biology, 14(3):293–315, March 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519367900823.

Wilkins:1962:FSLa

[WOB62a] D. J. Wilkins, R. H. Ottewill, and A. D. Bangham. On the flocculation of sheep leucocytes: I. Electrophoretic studies. Journal of Theoretical Biology, 2(2):165–175, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519362900450.

Wilkins:1962:FSLb

[WOB62b] D. J. Wilkins, R. H. Ottewill, and A. D. Bangham. On the flocculation of sheep leucocytes: II. Stability studies. Journal of Theoretical Biology, 2(2):176–191, March 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519362900462.

Wobschall:1968:EMN

[Wob68] D. Wobschall. An electret model of the nerve membrane. Journal of Theoretical Biology, 21(3):439–448, December 1968. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519368901264.

Wolff:1962:KTB

[Wol62a] Sheldon Wolff. The kinetics for two-break chromosome exchanges. Journal of Theoretical Biology, 3(2):304–314, September 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519362800198.

Wolken:1962:PST

[Wol62b] Jerome J. Wolken. Photoreceptor structures: + their molecular organization for energy transfer. Journal of Theoretical Biology, 3(2):192–194, IN2, 195–204, IN4, 205–208, September 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519362800149.

Wolpert:1969:PIS

[Wol69] L. Wolpert. Positional information and the spatial pattern of cellular differentiation. Journal of Theoretical Biology, 25(1):1– 47, October 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519369800160.

Willard:1963:EPM

[WP63] A. Willard and B. Pullman. Etude par la méthode du champ moleculaire self-consistant de la structure electronique des bases puriques et pyrimidiques d’intérêt biochimique. Journal of Theoretical Biology, 4(1):37–51, January 1963. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519363900997.

Walter:1967:MBL

[WPY67] Charles Walter, Rodger Parker, and Martynas Ycas. A model for binary logic in biochemical systems. Journal of Theoretical Biology, 15(2):208–217, May 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367902019.

Wu:1964:NGM

[Wu64] Tai Te Wu. A note on genetic mapping. Journal of Theoretical Biology, 7(3):430–434, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519364900153.

Wu:1967:RFP

[Wu67] T. T. Wu. Recombination frequencies of proximal markers in bacterial conjugation. Journal of Theoretical Biology, 17(1):40– 46, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900185.

Wuenscher:1969:NSC

[Wue69] James E. Wuenscher. Niche specification and competition modeling. Journal of Theoretical Biology, 25(3):436–443, December 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519369800317.

Wacker:1968:MCB

[WW68a] W. E. C. Wacker and R. J. P. Williams. Magnesium/calcium balances and steady states of biological systems. Journal of Theoretical Biology, 20(1):65–78, July 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900921.

Watts:1968:IME

[WW68b] R. L. Watts and D. C. Watts. The implications for molecular evolution of possible mechanisms of primary gene duplication. Journal of Theoretical Biology, 20(2):227–244, August 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368901926.

Wyatt:1964:CEC

[Wya64] H. V. Wyatt. Cations, enzymes and control of cell metabolism. Journal of Theoretical Biology, 6(3):441–470, May 1964. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251936490058X.

Weiss:1969:KRB

[WZ69] George H. Weiss and Gershom Zajicek. Kinetics of red blood cells following hemolysis. Journal of Theoretical Biology, 23(3): 475–491, June 1969. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519369900332.

Yamagata:1966:HAA

[Yam66] Yukio Yamagata. A hypothesis for the asymmetric appearance of biomolecules on earth. Journal of Theoretical Biology, 11(3):495– 498, August 1966. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251936690110X.

Ycas:1961:RAA

[Yca61] Martynas Ycas. Replacement of amino acids in proteins. Journal of Theoretical Biology, 1(2):244–257, April 1961. CODEN JTBIAP.

ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519361900480.

Yeisley:1964:ACS

[YP64] Warren G. Yeisley and Ernest C. Pollard. An analog computer study of differential equations concerned with bacterial cell synthesis. Journal of Theoretical Biology, 7(3):485–503, November 1964. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519364900189.

Yasuda:1967:TSF

[YS67] H. Yasuda and William Stone. Theoretical study of the fluid transport and the hydration of the cornea. Journal of Theoretical Biology, 16(1):111–134, July 1967. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519367900562.

Yas:1965:MCS

[YSB65] M. Yăs, M. Sugita, and Arlene Bensam. A model of cell size regulation. Journal of Theoretical Biology, 9(3):444–470, November 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519365900433.

Zajicek:1968:CMS

[Zaj68] Gershom Zajicek. A computer model simulating the behavior of adult red blood cells. Red cell model. Journal of Theoretical Biology, 19(1):51–66, April 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519368900040.

Zarybnicky:1969:MED

[Zár69] V. Zárybnický. Mechanism of t-even DNA ejection. Journal of Theoretical Biology, 22(1):33–40, IN1–IN2, 41–42, January 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519369900782.

Zelman:1968:ACF

[Zel68] D. Allen Zelman. An analysis of the constant field equation. Journal of Theoretical Biology, 18(3):396–398, March 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900878.

Zimmerman:1968:CAA

[ZES68] J. M. Zimmerman, Naomi Eliezer, and R. Simha. The characterization of amino acid sequences in proteins by statistical methods. Journal of Theoretical Biology, 21(2):170–201, November 1968. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519368900696.

Zuckerkandl:1965:MDE

[ZP65] Emile Zuckerkandl and Linus Pauling. Molecules as documents of evolutionary history. Journal of Theoretical Biology, 8(2):357– 366, March 1965. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519365900834.

Zimmerman:1965:KMM

[ZS65] J. M. Zimmerman and Robert Simha. The kinetics of multicenter macromolecule growth along a template. Journal of Theoretical Biology, 9(2):156–185, September 1965. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519365901062.

Zimmerman:1966:KCU

[ZS66] J. M. Zimmerman and Robert Simha. The kinetics of cooperative unwinding and template replication of biological macromolecales. Journal of Theoretical Biology, 13(??):106–130, December 1966. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519366900129.

Zotin:1967:TAD

[ZZ67] A. I. Zotin and R. S. Zotina. Thermodynamic aspects of developmental biology. Journal of Theoretical Biology, 17(1):57– 75, October 1967. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519367900203.

Zotin:1969:CAP

[ZZ69] A. I. Zotin and R. S. Zotina. Comments by the authors on the paper “thermodynamic aspect of developmental biology”. Journal of Theoretical Biology, 23(2):342–343, May 1969. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519369900459.